### v1

# Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment

Abstract
While large-scale unsupervised language models (LMs) capture broad world knowledge and reasoning capabilities, steering their behavior toward desired objectives remains challenging due to the lack of explicit supervision. Existing alignment techniques, such as reinforcement learning from human feedback (RLHF), rely on training a reward model and performing reinforcement learning to align with human preferences. However, RLHF is often computationally intensive, unstable, and sensitive to hyperparameters.
To address these limitations, Direct Preference Optimization (DPO) was introduced as a lightweight and stable alternative, enabling direct alignment of language models with pairwise preference data via classification loss. However, DPO and its extensions generally assume a single static preference distribution, limiting flexibility in multi-objective or dynamic alignment settings.
In this paper, we propose a novel framework: Multi-Preference Lambda-weighted Listwise DPO, which extends DPO to incorporate multiple human preference dimensions (e.g., helpfulness, harmlessness, informativeness) and enables dynamic interpolation through a controllable simplex-weighted formulation. Our method supports both listwise preference feedback and flexible alignment across varying user intents without re-training. Empirical and theoretical analysis demonstrates that our method is as effective as traditional DPO on static objectives while offering greater generality and adaptability for real-world deployment.
Introduction
Large language models (LLMs) trained on large-scale text corpora using next-token prediction objectives have demonstrated strong capabilities across a range of natural language understanding and generation tasks. However, these models often exhibit misalignment with human values and intentions, producing responses that may be factually incorrect, verbose, offensive, or otherwise undesirable. This gap arises in part because the standard unsupervised training objective does not explicitly encode human preferences or task-specific constraints.
To improve alignment with human intent, recent research has introduced fine-tuning techniques that incorporate human feedback. Reinforcement learning from human feedback (RLHF) is one prominent approach: it first trains a reward model from pairwise human comparisons, and then optimizes the language model via reinforcement learning to maximize this reward while maintaining similarity to the original policy. While effective, RLHF is known to be unstable, expensive to train, and sensitive to hyperparameter choices. It also requires online sampling and reward estimation, making it cumbersome to deploy at scale.
Direct Preference Optimization (DPO) was proposed as a simpler alternative to RLHF. Rather than relying on reinforcement learning, DPO directly optimizes the model’s output distribution to reflect human preferences using a pairwise classification objective. DPO has been shown to produce high-quality generations with significantly lower training complexity and resource demands. Despite its advantages, the original DPO formulation is limited in its ability to handle diverse or conflicting user preferences, as it is designed for a single fixed reward distribution.
In practical applications, however, alignment often involves multiple criteria—such as helpfulness, harmlessness, factuality, and succinctness—that may vary in importance depending on the use case or user. This motivates a more flexible framework capable of learning from multiple preference dimensions and generalizing across different preference trade-offs.
In this work, we introduce Multi-Preference Lambda-weighted Listwise DPO, a generalization of the DPO framework that incorporates multiple dimensions of preference feedback and enables interpolation over them via a weight vector $\lambda$ drawn from the probability simplex. Our method extends DPO from pairwise to listwise settings and supports alignment under dynamically changing preference configurations. At inference time, users or downstream systems can modulate the alignment behavior by selecting different $\lambda$ vectors without re-training the model.
We present a detailed theoretical derivation of our method starting from the KL-regularized RL objective, and show how it leads to a clean listwise loss that aggregates multiple reward dimensions. We demonstrate the effectiveness of our approach on multi-objective alignment benchmarks, showing that it achieves competitive or superior results compared to conventional DPO, while offering greater flexibility and control.
Related Work
Reinforcement Learning from Human Feedback (RLHF).
RLHF [19, 1] has become the de facto paradigm for aligning large language models (LLMs) with human preferences. This framework typically proceeds in three steps: (1) collect human comparisons over model outputs, (2) train a reward model to mimic these preferences, and (3) fine-tune the base model via reinforcement learning, most commonly using Proximal Policy Optimization (PPO) [22]. While effective in improving response helpfulness and safety, RLHF is complex to implement and computationally intensive, requiring online rollouts, heavy sampling, and sensitive hyperparameter tuning [31]. Moreover, it often suffers from reward hacking and policy collapse [4], making it difficult to scale or adapt in practical applications.
Direct Preference Optimization (DPO).
To address the limitations of RLHF, Direct Preference Optimization (DPO) [20] reparameterizes the reward model such that the optimal policy is analytically defined as a function of preference probabilities. This reformulation allows replacing RL with a simple binary classification objective on preference-labeled data. DPO avoids on-policy rollouts, maintains training stability, and significantly reduces computational cost. Several extensions have since been proposed, such as Quark [27] and LLaMA-DPO [16], which demonstrate competitive or superior performance to PPO-based RLHF in tasks like summarization, question answering, and dialogue generation. However, these approaches are limited to pairwise preferences and assume a single fixed alignment objective.
Listwise Preference Learning.
Traditional learning-to-rank methods distinguish between pairwise [3], pointwise [17], and listwise [25] approaches. Listwise methods are known to better capture global ranking structure, and have recently been adopted for preference modeling in LLMs. UltraFeedback [8] introduces a benchmark dataset where human raters assign rankings to multiple outputs per prompt. Listwise approaches in alignment, such as GLOW [26], use these rankings to learn better reward functions. Our work builds upon these insights but optimizes a listwise loss directly in policy space rather than via supervised reward modeling, enabling end-to-end fine-tuning.
Multi-Objective Alignment.
In many alignment scenarios, objectives such as helpfulness, harmlessness, factuality, and informativeness can conflict. Prior work has addressed this challenge via multi-reward reinforcement learning [12], reward decomposition [13], or manually interpolating between multiple reward functions [24]. However, these methods often require training a separate model for each configuration of preference weights. By contrast, our method incorporates preference weight vectors $\lambda$ during training and samples over the probability simplex, enabling the model to generalize to arbitrary trade-offs without retraining. This mechanism draws inspiration from meta-learning and universal value function approximators [21], but is specifically designed for alignment with human judgments.
Controllable and Steerable LMs.
Beyond RLHF, several techniques aim to control model behavior explicitly. These include prompt-based methods such as zero-shot prompting [2], few-shot prompting [30], and chain-of-thought [23], as well as supervised fine-tuning with control tokens [11] or conditioning variables [29]. Constrained decoding [18] and plug-and-play language models [7] provide post-hoc control. However, these methods typically address narrow or static attributes, lacking the flexibility to express complex human preferences. Our method introduces a theoretically grounded and fine-grained approach for multi-dimensional preference steerability without modifying prompts or decoding strategies.
Preliminaries
We begin by reviewing the foundation of Direct Preference Optimization (DPO) and its connection to KL-regularized reinforcement learning from human feedback (RLHF). Our setting assumes the existence of a base language model $\pi_{\text{ref}}$, a supervised fine-tuned (SFT) model, and a collection of ranked or pairwise human preferences. We aim to train a target policy $\pi_{\theta}$ that better aligns with human preferences while remaining close to $\pi_{\text{ref}}$.
Pairwise Preference Modeling.
Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one). These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$. A commonly adopted model is the Bradley-Terry (BT) model [bradley1952rank]:
$p^{*}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$ (1)
Given a dataset of comparisons $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, we can learn the reward function $r(x,y)$ by minimizing the negative log-likelihood under the BT model:
$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right],$ (2)
where $\sigma(z)=1/(1+\exp(-z))$ is the logistic sigmoid.
KL-Regularized RL Objective.
Once a reward model is trained, the standard RLHF approach fine-tunes the policy $\pi_{\theta}$ to maximize expected rewards while constraining deviation from the reference policy $\pi_{\text{ref}}$ via KL regularization:
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right],$ (3)
where $\beta>0$ controls the strength of the KL constraint.
This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$. However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO [22] to optimize.
Reparameterization and DPO.
DPO sidesteps explicit reward modeling and reinforcement learning by recognizing that the optimal policy $\pi_{r}$ under Equation 3 can be written as:
$\pi_{r}(y|x)\propto\pi_{\text{ref}}(y|x)\cdot\exp\left(\frac{1}{\beta}r(x,y)\right).$ (4)
Using Equation 4, we reparameterize the reward difference into the log-ratio of policies:
$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}.$ (5)
Replacing $\pi_{r}$ with $\pi_{\theta}$ yields the DPO loss:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$ (6)
This loss directly optimizes the policy parameters $\theta$ using preference comparisons, without sampling or reinforcement learning. It forms the foundation upon which we build our multi-preference and listwise extensions in later sections.
Multi-Preference Lambda-weighted Listwise DPO
Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we introduce a generalized framework that supports listwise preference supervision across multiple human objectives. In real-world settings, human judgments often reflect diverse criteria—such as helpfulness, harmlessness, and conciseness—and capturing these via a unified formulation enables controllable alignment across axes of interest.
Our key insight is that we can reparameterize multi-dimensional listwise preferences using a convex combination of multiple reward signals. Specifically, we define a simplex-weighted listwise loss that extends the DPO framework to dynamic, multi-objective optimization.
Listwise Preference Modeling.
Let $\{y_{1},y_{2},\dots,y_{N}\}$ be a set of candidate completions for a prompt $x$, and let $p^{*}(y_{i}|x)$ be a listwise human preference distribution over these candidates. Analogous to the pairwise DPO reparameterization, we define the listwise policy distribution under $\pi_{\theta}$ as:
$P_{\theta}(y_{i}|x)=\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}},$ (7)
which defines a softmax over reweighted log-ratios, consistent with KL-derived optimal policies.
We then define the listwise DPO objective as a cross-entropy loss between human preferences $p^{*}(y_{i}|x)$ and model predictions:
$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (8)
Lambda-weighted Preference Composition.
To enable multi-objective control, we assume access to $m$ distinct preference dimensions, each defined by a distribution $p^{*(k)}(y_{i}|x)$ for objective $k\in\{1,\dots,m\}$. For each prompt, we sample or specify a weight vector $\lambda\in\Delta^{m}$ (the probability simplex), and define the aggregated preference distribution:
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}|x).$ (9)
This allows us to interpolate among various alignment goals without retraining.
The full Lambda-weighted Listwise DPO objective becomes:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (10)
This expectation can be approximated by sampling $\lambda\sim\text{Unif}(\Delta^{m})$ or conditioning on user-specified preferences.
What does $\lambda$-DPO update do?
To understand the optimization dynamics, consider the gradient of the loss:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right].$ (11)
This gradient updates the model to increase likelihood of responses aligned with weighted human preferences, while decreasing likelihood of responses overrepresented in the reference policy. Importantly, when $\lambda$ is fixed (e.g., $\lambda=(1,0,0)$), the objective reduces to a single-objective listwise DPO loss.
Lambda-DPO Pipeline.
The training process proceeds as follows:
  1. 1.
    For each prompt $x$, obtain $N$ candidate completions $\{y_{1},\dots,y_{N}\}\sim\pi_{\text{ref}}(\cdot|x)$.
  2. 2.
    For each alignment dimension $k$, elicit a human listwise preference distribution $p^{*(k)}(y_{i}|x)$.
  3. 3.
    Sample or define $\lambda\in\Delta^{m}$, and compute the mixture preference $p^{\lambda}$.
  4. 4.
    Optimize $\pi_{\theta}$ using the listwise cross-entropy loss between $p^{\lambda}$ and $P_{\theta}$.
In practice, we can reuse datasets like UltraFeedback [8], where multiple preference criteria are annotated. During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
Theoretical Analysis of Multi-Preference Lambda-weighted Listwise DPO
While Direct Preference Optimization (DPO) offers a robust framework for aligning language models with human preferences via pairwise supervision, it inherently simplifies the preference space by assuming binary feedback between two completions. In this section, we analyze the theoretical implications of this simplification and motivate the use of a listwise and multi-preference extension.
Pairwise DPO and Its Limitations.
Given a prompt $x$ and two outputs $y_{w}$ (winner) and $y_{l}$ (loser), DPO defines a binary preference distribution:
$P^{*}(y_{i}|x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
This implicitly constrains the learning objective to satisfy $P^{*}(y_{w}|x)+P^{*}(y_{l}|x)=1$, ignoring the quality of other plausible candidates. Consequently, the optimized policy $\pi_{\theta}$ may fail to capture the full spectrum of user preferences, introducing estimation error.
Listwise Distribution via Softmax.
To mitigate this, we define a listwise target distribution $\pi^{*}$ over all completions $\{y_{1},y_{2},\dots,y_{N}\}$ for prompt $x$, derived from scalar reward scores $r(x,y_{i})$:
$\pi^{*}(y_{i}\mid x)=\frac{\exp(r(x,y_{i}))}{\sum_{j=1}^{N}\exp(r(x,y_{j}))}$ (13)
This distribution captures relative preference over all candidates, allowing for more representative updates and lower variance.
Multi-Preference Supervision.
Real-world alignment involves multiple competing objectives, such as helpfulness, safety, and factuality. Let $p^{*(k)}(y_{i}|x)$ denote the listwise preference under objective $k\in\{1,\dots,m\}$. We define a convex combination of these objectives via a preference vector $\lambda\in\Delta^{m}$:
$p^{\lambda}(y_{i}\mid x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}\mid x)$ (14)
The training objective minimizes the cross-entropy between the aggregated distribution $p^{\lambda}$ and the policy-induced distribution $P_{\theta}$:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (15)
Gradient Interpretation.
The gradient of this objective is:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
This update rule drives $\pi_{\theta}$ toward the weighted human preference distribution $p^{\lambda}$, allowing alignment to reflect fine-grained and multi-dimensional values.
Variance Reduction via Listwise Supervision.
Using a normalized softmax distribution $\pi^{*}$ derived from reward scores avoids the instability and high variance that arise when optimizing based on hard binary labels. Empirically, listwise training improves sample efficiency and convergence by utilizing the full output ranking structure.
Summary.
In summary, Multi-Preference Lambda-weighted Listwise DPO:
  • Recovers a more informative supervision signal via $\pi^{*}$.
  • Reduces estimation error and variance in gradient updates.
  • Enables alignment over a continuous space of user-defined objectives via $\lambda$.
Experiment Setup
Experimental Objectives and Comparison Targets
The primary goal of our experiments is to evaluate the performance of the proposed $\lambda$-DPO algorithm compared to the standard DPO method. Specifically, we assess their effectiveness across different model architectures and training settings using public preference datasets and standard evaluation benchmarks.
We consider two levels of comparison:
Model comparison: We experiment with two strong open-source language models—LLaMA3-8B and Qwen. These models differ in architecture, pretraining corpora, and tokenizer, providing a robust testbed for evaluating generalizability.
Method comparison: We compare DPO and $\lambda$-DPO. For each method and model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure that comparisons are fair and reflect the best achievable performance under comparable resource constraints.
Training Settings
Inspired by the SimPO baseline setup, we adopt two training pipelines for both models:
  • Base: The model is directly trained on the UltraFeedback dataset using DPO or $\lambda$-DPO. No supervised fine-tuning is applied before preference optimization. This setting isolates the influence of the preference learning method alone.
  • Finetuned-Base: The model is first supervised fine-tuned (SFT) on the UltraChat-200k dataset, which contains approximately 200K high-quality dialogue examples. The resulting SFT model is then optimized using UltraFeedback preferences via DPO or $\lambda$-DPO. This pipeline reflects realistic alignment procedures.
All training is conducted on A100 GPUs, with consistent batch sizes and sequence lengths. For each setting, hyperparameters are tuned via validation performance on reward preference accuracy and KL stability.
Evaluation Benchmarks and Metrics
To comprehensively evaluate model behavior, we use six established benchmarks:
  • MMLU [10]: 57 subject areas with around 15K test questions designed to test academic and professional knowledge.
  • ARC (Challenge) [5]: Grade-school level science questions ($\sim$1.1K) focusing on multi-step reasoning.
  • HellaSwag [28]: 10K commonsense sentence completion tasks with subtle distractors.
  • TruthfulQA [15]: 817 questions designed to expose model hallucination and fact inconsistency.
  • Winograd [14]: 273 examples targeting pronoun coreference and syntactic disambiguation.
  • GSM8K [6]: 8.5K arithmetic reasoning problems requiring multi-step solutions.
We evaluate models along several axes:
  • Win Rate (WR): The proportion of pairwise comparisons where the model’s output is preferred by the reward model.
  • lm-eval-harness Score [9]: Average score given by GPT-4 on multi-turn QA prompts, covering helpfulness, correctness, and language quality.
  • Reward Accuracy (%): The percentage of outputs that match the preferred choice of the reward model.
  • KL Divergence: The average KL divergence between the trained policy and the SFT policy, indicating policy shift magnitude.
  • Run Time: Total training duration under fixed hardware settings.
  • Peak GPU Memory: Maximum memory usage during training.
  • Spearman Correlation: Correlation between output log-probability and output length, used to measure length bias.
Ablation Experiment
To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct two complementary sets of ablation experiments, each targeting a different aspect of the framework’s flexibility and effectiveness.
The first set of experiments aims to isolate the effect of explicit preference vector interpolation by comparing $\lambda$-DPO models trained with and without multi-dimensional preference aggregation. In this setting, the random $\lambda$ and Gaussian $\lambda$ variants are used to provide dynamic and smoothly varying interpolation weights over multiple preference dimensions, while the unimodal $\lambda$ variant—represented by fixed one-hot vectors such as $[1,0,0,0]$—serves as the baseline where no interpolation is applied, i.e., each training instance aligns with a single fixed objective. This comparison helps assess whether the simplex-weighted combination of preferences brings measurable benefits in robustness, generalization, or alignment quality.
The second set of experiments directly compares $\lambda$-DPO with the original DPO framework. To ensure a fair comparison that isolates the algorithmic difference—rather than the benefit from multiple objectives—this evaluation uses only the unimodal $\lambda$ setting in $\lambda$-DPO, effectively aligning it with single-objective training, just like standard DPO. This design ensures that both models optimize against one fixed preference dimension at a time, allowing us to accurately assess the impact of listwise supervision and policy reparameterization introduced by $\lambda$-DPO.
For the $\lambda$ schedule in total, we therefore investigate three variants:
The random $\lambda$ variant samples a different vector from the simplex for each instance, introducing stochastic diversity across training and encouraging robustness.
The Gaussian $\lambda$ variant samples values from a standard normal distribution and normalizes them to lie on the simplex, enabling smooth but non-uniform emphasis over dimensions.
The unimodal $\lambda$ variant uses fixed one-hot vectors to eliminate interpolation and isolate individual preference components.
Together, these ablation studies provide insights into the roles of interpolation, multi-dimensional preference modeling, and the structural benefits of our listwise optimization strategy.
Acknowledgements
We would like to thank the contributors of the original DPO framework for their foundational work, which inspired this extension. Special thanks to the Hugging Face team and the LLaMA-Factory project for providing robust tools and open-source infrastructure for preference-based fine-tuning. We also acknowledge valuable discussions with colleagues and reviewers who provided feedback during the early stages of this work. Their insights have significantly improved the clarity and rigor of our theoretical analysis.
This work was supported in part by academic research funding and compute resources generously provided by [Your Institution Name Here].
Appendix A A Mathematical Derivations
A.1 Deriving the Final Training Objective
We begin by defining the base reward model used in DPO. Let $\pi_{\theta}(y|x)$ denote the current policy and $\pi_{\text{ref}}(y|x)$ a fixed reference policy. Given input $x$ and candidate outputs $\{y_{1},\ldots,y_{N}\}$, the reward is defined via KL divergence as:
$r_{\theta}(x,y)=\beta\log\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)+\beta\log Z(x)$ (17)
where $Z(x)$ is a normalizing constant ensuring scale-invariance:
$Z(x)=\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)$ (18)
The predicted distribution over completions is then:
$P_{\theta}(y_{i}|x)=\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}$ (19)
Given a human-provided listwise preference distribution $p^{*}_{i}$, we minimize the cross-entropy between $p^{*}_{i}$ and $P_{\theta}$:
$\mathcal{L}^{\text{DPO}}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}\right)\right]$ (20)
Substituting the reward expression and simplifying:
$\mathcal{L}^{\text{DPO}}(\theta)&=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}Z(x)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}Z(x)^{\beta}}\right)\right]\\ &=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]\end{aligned}$
This becomes the canonical DPO loss. To extend this to multiple preference dimensions, we assume there are $m$ reward components indexed by $k$, each with its own preference distribution $p_{i}^{*(k)}$:
$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]$ (21)
We then define a user-controllable or sampled weight vector $\lambda\in\Delta^{m}$ (the $m$-dimensional probability simplex: $\lambda_{k}\geq 0$, $\sum_{k}\lambda_{k}=1$), and define the aggregated loss:
$\mathcal{L}^{\lambda}(\theta)=\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)$ (22)
To enable generalization over arbitrary preferences, we optimize the expected loss over $\lambda\sim\text{Unif}(\Delta^{m})$:
$\mathcal{L}_{\text{final}}(\theta)=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\mathcal{L}^{\lambda}(\theta)\right]=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)\right]$ (23)
A.2 From Final Objective to Lambda-DPO Form
To convert Eq. (25) to a form that operates on distributions directly, we change the order of summation. Recall:
$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (24)
So the aggregated expected loss becomes:
$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathbb{E}_{(x,\{y_{i}\})}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]\right]$ (25)
$=\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[-\sum_{i=1}^{N}\sum_{k=1}^{m}\lambda_{k}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]$ (26)
We now define the aggregated distribution over outputs:
$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}p^{*(k)}(y_{i}|x)$ (27)
So the objective becomes:
$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (28)
Thus, by switching the summation order (i.e., aggregating over preference weights before computing loss), we arrive at an equivalent but more compact objective that is easier to optimize and sample in practice.
A.3 Gradient of Lambda-DPO Loss
We compute the gradient of $\mathcal{L}_{\lambda\text{-DPO}}$:
$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}$ $=-\mathbb{E}_{x,\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\nabla_{\theta}\log P_{\theta}(y_{i}\mid x)\right]$ (29)
$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (30)
where we use:
$\nabla_{\theta}P_{\theta}(y_{i}|x)=P_{\theta}(y_{i}|x)\cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)$ (31)
This completes the derivation of the Lambda-weighted DPO objective and its gradient form.
Appendix B B Implementation of _compute_lambda_dpo_loss
This function sequentially performs the following steps:
B.1 Input Tensors and Log‑Probabilities
  1. 1.
    Load batch tensors:
[…]
    input_ids = inputs["input_ids"]
    attention_mask = inputs["attention_mask"]
    labels = inputs["labels"]
[…]
  2. 2.
    Compute token-level and sequence-level log-probabilities:
    $\log\pi_{\theta}(y|x)$ $=\log\text{softmax}(\text{model}(x)[:,:-1])$
    mask $=\mathbb{1}_{\{y_{t}\neq\text{PAD}\}}$
    $\log\pi_{\theta}(y|x)$ $=\sum_{t}\texttt{mask}_{t}\cdot\log\pi_{\theta}(y_{t}|x_{t})$
    logits = model(input_ids=input_ids, attention_mask=attention_mask).logits
    log_probs = F.log_softmax(logits[:, :-1], dim=-1)
    labels_shifted = labels[:, 1:].clone()
    mask = labels_shifted != self.label_pad_token_id
    labels_shifted[labels_shifted == self.label_pad_token_id] = 0
    label_log_probs = torch.gather(log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
    seq_log_probs = (label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
B.2 Reference Model and Relative Reward
  1. 3.
    Compute reference log-probabilities if available:
    $\log\pi_{\text{ref}}(y|x)=\begin{cases}\log\text{softmax}(\text{ref\_model}(x)[:,:-1]),&\text{if reference model exists}\\ 0,&\text{otherwise}\end{cases}$
    with torch.no_grad():
    if self.finetuning_args.use_ref_model and self.ref_model is not None:
    ref_logits = self.ref_model(input_ids=input_ids, attention_mask=attention_mask).logits
    ref_log_probs = F.log_softmax(ref_logits[:, :-1], dim=-1)
    ref_label_log_probs = torch.gather(ref_log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
    ref_seq_log_probs = (ref_label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
    else:
    ref_seq_log_probs = torch.zeros_like(seq_log_probs)
  2. 4.
    Compute relative reward and reshape:
    $r_{\theta}(x,y)$ $=\beta(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x))$
    $r_{\theta}$ $\in\mathbb{R}^{B\times 4\times 4},\quad\pi^{*}\in\mathbb{R}^{B\times 4\times 4}$
    r_theta = self.beta * (seq_log_probs - ref_seq_log_probs)
    B = input_ids.size(0) // 16
    r_theta = r_theta.view(B, 4, 4)
    pi_target = pi_target.view(B, 4, 4)
B.3 Listwise Loss by Preference Dimension
  1. 5.
    For each dimension $i=1,2,3,4$:
    $\ell^{(i)}$ $=-\sum_{j}p^{*(i)}_{j}\cdot\log\text{softmax}(r^{(i)})_{j}$
    listwise_losses = []
    for i in range(4):
    r = r_theta[:, i, :]
    p = pi_target[:, i, :] + 1e-10
    log_softmaxed = r - torch.logsumexp(r, dim=1, keepdim=True)
    loss = -(p * log_softmaxed).sum(dim=1).mean()
    listwise_losses.append(loss)
B.4 Lambda Weights and Final Loss
  1. 6.
    Sample weights $\boldsymbol{\lambda}\sim\Delta^{4}$, and compute:
    $\mathcal{L}(\theta)$ $=\sum_{i=1}^{4}\lambda_{i}\ell^{(i)}$
    weights = torch.rand(4, device=input_ids.device)
    weights /= weights.sum()
    final_loss = sum(w * l for w, l in zip(weights, listwise_losses))
    if return_outputs:
    return final_loss, {"loss_components": listwise_losses}
    return final_loss
This function performs sequence probability calculation, reference model comparison, listwise cross-entropy over preference dimensions, and $\lambda$-weighted aggregation. It fully realizes the lambda-DPO loss objective described in the paper’s methodology section.

### v2

@@
 Together, these ablation studies provide insights into the roles of interpolation, multi-dimensional preference modeling, and the structural benefits of our listwise optimization strategy.
+Conclusion
+We presented Multi-Preference Lambda-weighted Listwise DPO, a generalization of Direct Preference Optimization (DPO) that enables alignment of language models with diverse and dynamic human preferences.
+By extending DPO from a pairwise to a listwise formulation and incorporating multiple preference dimensions via a lambda-weighted aggregation, our approach captures richer supervision signals and allows for controllable alignment without retraining.
+Unlike traditional DPO, which assumes a single binary comparison per training instance, Lambda-DPO models the full distribution over ranked outputs, supporting scenarios where user intent is nuanced, multi-objective, or even conflicting.
+This enables practical applications such as personalized assistants that can adapt to individual user trade-offs (e.g., prioritizing helpfulness over conciseness), regulatory-sensitive deployments that balance safety and informativeness, and modular alignment pipelines where preferences may evolve over time.
+Theoretically, our method is grounded in a KL-regularized RL objective, and empirically, we demonstrate competitive or superior performance to standard DPO across alignment benchmarks.
+By reducing estimation variance and providing interpretable preference control via the lambda vector, Lambda-DPO offers a scalable and expressive solution for real-world preference alignment.
+In future work, we plan to explore more structured lambda scheduling, integration with curriculum learning, and extensions to conversational or multimodal alignment tasks.
 Acknowledgements
@@
 This function sequentially performs the following steps:
-B.1 Input Tensors and Log‑Probabilities
+B.1 Input Tensors and Chunked Log-Probabilities
   1. 1.
-    Load batch tensors:
+    Load the batch tensors:
 […]
@@
   2. 2.
-    Compute token-level and sequence-level log-probabilities:
-    $\log\pi_{\theta}(y|x)$ $=\log\text{softmax}(\text{model}(x)[:,:-1])$
-    mask $=\mathbb{1}_{\{y_{t}\neq\text{PAD}\}}$
-    $\log\pi_{\theta}(y|x)$ $=\sum_{t}\texttt{mask}_{t}\cdot\log\pi_{\theta}(y_{t}|x_{t})$
-    logits = model(input_ids=input_ids, attention_mask=attention_mask).logits
+    Set the chunk size and compute token/sequence log-probabilities:
+    To reduce memory usage, we split the batch into chunks of size $C$, denoted by:
+[…]
+[…]
+    chunk_size = self.lambda_dpo_chunk_size or input_ids.size(0)
+    seq_log_probs_list = []
+    ref_seq_log_probs_list = []
+    for start in range(0, input_ids.size(0), chunk_size):
+    end = start + chunk_size
+    chunk_ids = input_ids[start:end]
+    chunk_mask = attention_mask[start:end]
+    chunk_labels = labels[start:end]
+    logits = model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
     log_probs = F.log_softmax(logits[:, :-1], dim=-1)
-    labels_shifted = labels[:, 1:].clone()
+    labels_shifted = chunk_labels[:, 1:].clone()
     mask = labels_shifted != self.label_pad_token_id
@@
     label_log_probs = torch.gather(log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
-    seq_log_probs = (label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
+    seq_lp = (label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
+    seq_log_probs_list.append(seq_lp)
+  3. 3.
+    Compute reference log-probabilities (if applicable):
+    $\log\pi_{\text{ref}}(y|x)=\begin{cases}\log\text{softmax}(\text{ref\_model}(x)),&\text{if ref model is used}\\ 0,&\text{otherwise}\end{cases}$
+    with torch.no_grad():
+    if self.finetuning_args.use_ref_model and self.ref_model is not None:
+    ref_logits = self.ref_model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
+    ref_log_probs = F.log_softmax(ref_logits[:, :-1], dim=-1)
+    ref_label_log_probs = torch.gather(ref_log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
+    ref_lp = (ref_label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
+    else:
+    ref_lp = torch.zeros_like(seq_lp)
+    ref_seq_log_probs_list.append(ref_lp)
+    seq_log_probs = torch.cat(seq_log_probs_list, dim=0)
+    ref_seq_log_probs = torch.cat(ref_seq_log_probs_list, dim=0)
 B.2 Reference Model and Relative Reward
   1. 3.
-    Compute reference log-probabilities if available:
-    $\log\pi_{\text{ref}}(y|x)=\begin{cases}\log\text{softmax}(\text{ref\_model}(x)[:,:-1]),&\text{if reference model exists}\\ 0,&\text{otherwise}\end{cases}$
-    with torch.no_grad():
-    if self.finetuning_args.use_ref_model and self.ref_model is not None:
-    ref_logits = self.ref_model(input_ids=input_ids, attention_mask=attention_mask).logits
-    ref_log_probs = F.log_softmax(ref_logits[:, :-1], dim=-1)
-    ref_label_log_probs = torch.gather(ref_log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
-    ref_seq_log_probs = (ref_label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
-    else:
-    ref_seq_log_probs = torch.zeros_like(seq_log_probs)
-  2. 4.
-    Compute relative reward and reshape:
-    $r_{\theta}(x,y)$ $=\beta(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x))$
+    Compute the relative reward:
+    $r_{\theta}(x,y)=\beta\left(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x)\right)$
+    Then reshape:
     $r_{\theta}\in\mathbb{R}^{B\times 4\times 4},\quad\pi^{*}\in\mathbb{R}^{B\times 4\times 4}$
@@
   1. 5.
-    For each dimension $i=1,2,3,4$:
-    $\ell^{(i)}$ $=-\sum_{j}p^{*(i)}_{j}\cdot\log\text{softmax}(r^{(i)})_{j}$
+    For each preference axis $i\in\{1,2,3,4\}$, compute the listwise loss:
+    $\ell^{(i)}=-\sum_{j}p^{*(i)}_{j}\cdot\log\left(\frac{\exp(r^{(i)}_{j})}{\sum_{j^{\prime}}\exp(r^{(i)}_{j^{\prime}})}\right)$
     listwise_losses = []
@@
   1. 6.
-    Sample weights $\boldsymbol{\lambda}\sim\Delta^{4}$, and compute:
+    Sample a random weight vector $\boldsymbol{\lambda}\sim\Delta^{4}$, and compute:
     $\mathcal{L}(\theta)=\sum_{i=1}^{4}\lambda_{i}\ell^{(i)}$
@@
     return final_loss
+Appendix C: Humorous Overview of the Added Classes
+C.1 UltrafeedbackDatasetConverter
+Input
+A single UltraFeedback record. It contains:
+  • a user prompt (instruction)
+  • multiple completions, each with four ratings: helpfulness, honesty, instruction_following, and truthfulness.
+Output
+A dictionary with the prompt and, for every dimension, the raw scores plus a preference vector.
+Example
+Input:
+instruction = "Explain photosynthesis."
+completions = [
+{"response": "...", "annotations": {"helpfulness": {"Rating": 4}}},
+{"response": "...", "annotations": {"helpfulness": {"Rating": 2}}},
+{"response": "...", "annotations": {"helpfulness": {"Rating": 5}}},
+{"response": "...", "annotations": {"helpfulness": {"Rating": 3}}}
+Output (simplified):
+"_prompt": [{"role": "user", "content": "Explain photosynthesis."}],
+"helpfulness": [
+{"model": "?", "content": "...", "score": 4},
+"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22]}
+C.2 ListwiseDatasetProcessor
+Input
+A dictionary produced by the converter that contains _prompt, the lists of responses for each dimension, and _pi_target.
+Output
+A flattened structure ready for training: for every sample there are 16 sequences of token IDs (one per response per dimension), matching labels and attention masks, and the list of probabilities.
+Example
+Input:
+"_prompt": ...,
+"helpfulness": [
+{"content": "A"}, {"content": "B"}, {"content": "C"}, {"content": "D"}
+"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22], ...}
+Output (illustrative shapes):
+"input_ids": [seq1, seq2, ..., seq16], # each seq corresponds to a prompt+response
+"labels": [...], # same layout as input_ids
+"attention_mask": [...], # masks per sequence
+"pi_target": [0.24, 0.09, 0.45, 0.22, ...] # flattened list of probabilities
+C.3 ListwiseDataCollatorWithPadding
+Input
+A list of processed samples where each sample holds 16 sequences of token IDs and the flattened probability list.
+Output
+A padded batch containing tensors input_ids, attention_mask, labels, and an accompanying tensor of preference values aligned with the responses.
+Example
+Input:
+# List of samples where each sample holds 16 sequences and
+# a corresponding 16-element probability list
+[example1, example2, ..., example_B] # B = batch size
+Output:
+"input_ids": Tensor of shape [B * 16, L],
+"attention_mask": Tensor of shape [B * 16, L],
+"labels": Tensor of shape [B * 16, L],
+"pi_target": Tensor of shape [B * 16]
+These three classes transform UltraFeedback data into tensors suitable for Lambda-weighted DPO training, while preserving the four-response structure and converting ratings into softmax preference vectors.
 This function performs sequence probability calculation, reference model comparison, listwise cross-entropy over preference dimensions, and $\lambda$-weighted aggregation.

### v3

@@
 This design ensures that both models optimize against one fixed preference dimension at a time, allowing us to accurately assess the impact of listwise supervision and policy reparameterization introduced by $\lambda$-DPO.
-For the $\lambda$ schedule in total, we therefore investigate three variants:
-The random $\lambda$ variant samples a different vector from the simplex for each instance, introducing stochastic diversity across training and encouraging robustness.
-The Gaussian $\lambda$ variant samples values from a standard normal distribution and normalizes them to lie on the simplex, enabling smooth but non-uniform emphasis over dimensions.
-The unimodal $\lambda$ variant uses fixed one-hot vectors to eliminate interpolation and isolate individual preference components.
-Together, these ablation studies provide insights into the roles of interpolation, multi-dimensional preference modeling, and the structural benefits of our listwise optimization strategy.
+For the $\lambda$ schedule, we investigate three variants:
+[…]
+  • Gaussian $\lambda$: we sample each component from a standard normal distribution and then normalize the vector to lie on the simplex, thereby enabling smoothly varying but non-uniform emphasis across dimensions.
+  • Unimodal $\lambda$: we use fixed one-hot vectors (e.g., $(1,0,0,\dots)$), effectively disabling interpolation and isolating optimization to a single preference dimension per instance.
+Together, these ablations help us understand how balanced weighting, smooth interpolation, and single-objective optimization each contribute to the effectiveness of our listwise $\lambda$-DPO framework.
+Comparative Experiment
+We conduct a comparative experiment to evaluate the performance of standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups.
+Specifically, we focus on two representative models—LLaMA 3.2B Instruct and Qwen 1.8B Chat—and assess generalization across five commonly used benchmarks using the lm-eval-harness framework.
+Both methods are trained from the same base model (i.e., without supervised fine-tuning), using UltraFeedback preference data.
+Main Results
+Table 1 reports the accuracy of DPO and $\lambda$-DPO on five representative benchmarks for both LLaMA 3.2 1B Instruct and Qwen 1.8B Chat.
+On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +1.8% on GSM8K, and +3.6% on TruthfulQA.
+Qwen 1.8B Chat also sees improvements on most benchmarks, particularly GSM8K (+14.7%) and MMLU (+3.9%).
+DPO only slightly outperforms $\lambda$-DPO on TruthfulQA in the Qwen setting.
+Model Method MMLU ARC HellaSwag TruthfulQA GSM8K
+LLaMA 3.2 1B Instruct DPO 0.458 0.364 0.456 0.378 0.696
+$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714
+Qwen 1.8B Chat DPO 0.379 0.271 0.418 0.366 0.008
+$\lambda$-DPO 0.418 0.284 0.428 0.312 0.155
+Table 1: Accuracy comparison between DPO and $\lambda$-DPO on five benchmarks using LLaMA 3.2 1B Instruct and Qwen 1.8B Chat. Bold numbers indicate better performance.
+Visualization
+To further illustrate these results, we provide bar charts comparing DPO and $\lambda$-DPO performance on the two models.
+Figure 1: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
+Figure 2: Accuracy comparison between DPO and $\lambda$-DPO on Qwen 1.8B Chat.
+Performance Analysis and Discussion
+As shown in both Table 1 and Figures 14, $\lambda$-DPO consistently improves performance across most tasks.
+The method yields notable improvements on ARC, GSM8K, and MMLU, while maintaining parity on Winogrande.
+TruthfulQA is the only benchmark where DPO slightly outperforms $\lambda$-DPO (in the Qwen model).
+These results suggest that $\lambda$-DPO effectively captures human preference signals across diverse task types.
+Its listwise preference modeling contributes to better alignment with human-evaluated quality metrics such as factuality, helpfulness, and fluency.
+Nonetheless, for certain adversarial or logic-heavy tasks like TruthfulQA, the benefits of listwise supervision may be less reliable.
+Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting [23]—could potentially mitigate this gap.
+Conclusion of Comparative Analysis
+In summary, $\lambda$-DPO demonstrates consistent improvements over DPO in most evaluation scenarios, confirming the effectiveness of listwise preference supervision.
+Its generalization advantage makes it a strong candidate for broader alignment tasks.
+Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning methods to further enhance performance on complex benchmarks.
 Conclusion
@@
 Appendix B B Implementation of _compute_lambda_dpo_loss
-This function sequentially performs the following steps:
-B.1 Input Tensors and Chunked Log-Probabilities
+This function executes the following procedure:
+B.1 Input Tensors and Chunked Computation
   1. 1.
     Load the batch tensors:
-[…]
     input_ids = inputs["input_ids"]
@@
     labels = inputs["labels"]
-[…]
+    pi_target = inputs["pi_target"]
   2. 2.
-    Set the chunk size and compute token/sequence log-probabilities:
-    To reduce memory usage, we split the batch into chunks of size $C$, denoted by:
-[…]
-[…]
-    chunk_size = self.lambda_dpo_chunk_size or input_ids.size(0)
+    Assume the batch contains 4 preference dimensions and 4 candidate responses (16 sequences in total). Reshape the tensors so that each preference dimension shares the same 4 unique responses:
+    num_pref, num_resp = 4, 4
+    B = input_ids.size(0) // (num_pref * num_resp)
+    seq_len = input_ids.size(1)
+    ids_view = input_ids.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
+    mask_view = attention_mask.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
+    labels_view = labels.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
+  3. 3.
+    Choose a chunk size (by default equal to ids_view.size(0)) to compute log-probabilities in smaller batches. Optionally compute reference model probabilities:
+    chunk_size = self.lambda_dpo_chunk_size or ids_view.size(0)
     seq_log_probs_list = []
     ref_seq_log_probs_list = []
-    for start in range(0, input_ids.size(0), chunk_size):
+    for start in range(0, ids_view.size(0), chunk_size):
     end = start + chunk_size
-    chunk_ids = input_ids[start:end]
-    chunk_mask = attention_mask[start:end]
-    chunk_labels = labels[start:end]
+    chunk_ids = ids_view[start:end]
+    chunk_mask = mask_view[start:end]
+    chunk_labels = labels_view[start:end]
     logits = model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
-    log_probs = F.log_softmax(logits[:, :-1], dim=-1)
-    labels_shifted = chunk_labels[:, 1:].clone()
-    mask = labels_shifted != self.label_pad_token_id
-    labels_shifted[labels_shifted == self.label_pad_token_id] = 0
-    label_log_probs = torch.gather(log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
-    seq_lp = (label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
-    seq_log_probs_list.append(seq_lp)
-  3. 3.
-    Compute reference log-probabilities (if applicable):
-    $\log\pi_{\text{ref}}(y|x)=\begin{cases}\log\text{softmax}(\text{ref\_model}(x)),&\text{if ref model is used}\\ 0,&\text{otherwise}\end{cases}$
+    seq_lp, valid_len = get_batch_logps(logits, chunk_labels, self.label_pad_token_id)
+    seq_log_probs_list.append(seq_lp / valid_len)
     with torch.no_grad():
@@
     ref_logits = self.ref_model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
-    ref_log_probs = F.log_softmax(ref_logits[:, :-1], dim=-1)
-    ref_label_log_probs = torch.gather(ref_log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
-    ref_lp = (ref_label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
+    ref_lp, ref_len = get_batch_logps(ref_logits, chunk_labels, self.label_pad_token_id)
+    ref_lp = ref_lp / ref_len
     else:
@@
     ref_seq_log_probs_list.append(ref_lp)
+  4. 4.
+    Concatenate the chunk results and share each set of four log-probabilities across the preference dimensions:
     seq_log_probs = torch.cat(seq_log_probs_list, dim=0)
     ref_seq_log_probs = torch.cat(ref_seq_log_probs_list, dim=0)
+    seq_log_probs = seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
+    ref_seq_log_probs = ref_seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
 B.2 Reference Model and Relative Reward
-  1. 3.
-    Compute the relative reward:
-    $r_{\theta}(x,y)=\beta\left(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x)\right)$
-    Then reshape:
-    $r_{\theta}\in\mathbb{R}^{B\times 4\times 4},\quad\pi^{*}\in\mathbb{R}^{B\times 4\times 4}$
+  1. 5.
+    Compute the relative reward and reshape back to $(B,4,4)$:
     r_theta = self.beta * (seq_log_probs - ref_seq_log_probs)
-    B = input_ids.size(0) // 16
-    r_theta = r_theta.view(B, 4, 4)
-    pi_target = pi_target.view(B, 4, 4)
+    r_theta = r_theta.view(B, num_pref, num_resp)
+    pi_target = pi_target.view(B, num_pref, num_resp)
 B.3 Listwise Loss by Preference Dimension
-  1. 5.
+  1. 6.
     For each preference axis $i\in\{1,2,3,4\}$, compute the listwise loss:
-    $\ell^{(i)}=-\sum_{j}p^{*(i)}_{j}\cdot\log\left(\frac{\exp(r^{(i)}_{j})}{\sum_{j^{\prime}}\exp(r^{(i)}_{j^{\prime}})}\right)$
     listwise_losses = []
-    for i in range(4):
+    for i in range(num_pref):
     r = r_theta[:, i, :]
@@
 B.4 Lambda Weights and Final Loss
-  1. 6.
-    Sample a random weight vector $\boldsymbol{\lambda}\sim\Delta^{4}$, and compute:
-    $\mathcal{L}(\theta)=\sum_{i=1}^{4}\lambda_{i}\ell^{(i)}$
-    weights = torch.rand(4, device=input_ids.device)
+  1. 7.
+    The current implementation equally weights the preference dimensions (no random sampling of $\lambda$):
+    num_losses = len(listwise_losses)
+    weights = torch.ones(num_losses, device=input_ids.device)
     weights /= weights.sum()
@@
 It fully realizes the lambda-DPO loss objective described in the paper’s methodology section.
+Appendix C D Implementation of _compute_ultra_dpo_loss
+This routine computes the DPO loss when training on the UltraFeedback dataset, where each example contains four candidate responses with human preference scores. It proceeds as follows:
+D.1 Input Aggregation
+  1. 1.
+    Load the tensors input_ids, attention_mask, labels, and pi_target.
+  2. 2.
+[…]
+  3. 3.
+    Apply uniform weights across the four preference dimensions to obtain an aggregated distribution for each example:
+    $\text{aggregated}=\sum_{i=1}^{4}\tfrac{1}{4}\,\pi^{*(i)}.$
+D.2 Selecting Best and Worst Responses
+  1. 4.
+    Determine the best response index via $\arg\max$ over aggregated and the worst index via $\arg\min$.
+  2. 5.
+    Gather the corresponding best and worst sequences from the original batch for further processing.
+D.3 Log-Probability Computation
+  1. 6.
+    Pass the selected sequences through the model, compute token log-probabilities, and average over tokens to obtain sequence log-probabilities.
+  2. 7.
+    If a reference model is provided, compute its log-probabilities in the same way; otherwise use zero tensors as the reference scores.
+D.4 Final DPO Loss
+  1. 8.
+    Split the policy and reference log-probabilities into best and worst halves.
+  2. 9.
+    Invoke compute_preference_loss on these values to obtain per-example DPO losses, then average across the batch:
+    $\mathcal{L}_{\mathrm{DPO}}=\frac{1}{B}\sum_{b=1}^{B}\mathrm{loss}(y^{\mathrm{best}}_{b},\,y^{\mathrm{worst}}_{b}).$
+Appendix D E Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
+We provide benchmark accuracy comparisons between DPO and $\lambda$-DPO across multiple language models on the UltraFeedback dataset.
+Performance is reported across six tasks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
+E.1 Qwen-1.8 (Base)
+Figure 3 shows that $\lambda$-DPO achieves modest improvements over standard DPO on most tasks using the Qwen-1.8 base model.
+Figure 3: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 (base).
+E.2 Qwen-1.8 Chat
+Figure 4 presents results on Qwen-1.8 Chat, where $\lambda$-DPO consistently improves accuracy, especially on reasoning-heavy benchmarks like arc and hellaswag.
+Figure 4: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 Chat.
+E.3 LLaMA 3.2 1B (Base)
+As shown in Figure 5, $\lambda$-DPO outperforms DPO slightly on LLaMA 3.2 1B base across four out of six tasks, demonstrating generalization beyond the Qwen series.
+Figure 5: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B base model.
+E.4 LLaMA 3.2 1B (Instruct)
+Figure 6 illustrates results for the Instruct-tuned variant of LLaMA 3.2 1B. The performance gap between $\lambda$-DPO and DPO is generally preserved, indicating robustness to instruction tuning.
+Figure 6: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
+The function thus reduces the four-way UltraFeedback preferences to a single pairwise comparison per prompt and optimizes the model via DPO.

### v4

@@
 Abstract
-While large-scale unsupervised language models (LMs) capture broad world knowledge and reasoning capabilities, steering their behavior toward desired objectives remains challenging due to the lack of explicit supervision.
-Existing alignment techniques, such as reinforcement learning from human feedback (RLHF), rely on training a reward model and performing reinforcement learning to align with human preferences.
-However, RLHF is often computationally intensive, unstable, and sensitive to hyperparameters.
-To address these limitations, Direct Preference Optimization (DPO) was introduced as a lightweight and stable alternative, enabling direct alignment of language models with pairwise preference data via classification loss.
-However, DPO and its extensions generally assume a single static preference distribution, limiting flexibility in multi-objective or dynamic alignment settings.
-In this paper, we propose a novel framework: Multi-Preference Lambda-weighted Listwise DPO, which extends DPO to incorporate multiple human preference dimensions (e.g., helpfulness, harmlessness, informativeness) and enables dynamic interpolation through a controllable simplex-weighted formulation.
-Our method supports both listwise preference feedback and flexible alignment across varying user intents without re-training.
-Empirical and theoretical analysis demonstrates that our method is as effective as traditional DPO on static objectives while offering greater generality and adaptability for real-world deployment.
+While large language models (LLMs) excel at text generation, aligning them with human preferences remains challenging.
+Reinforcement learning from human feedback (RLHF) improves alignment but is costly and unstable.
+Direct Preference Optimization (DPO) offers a simpler alternative, yet assumes a fixed, single-dimensional preference.
+We propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that supports multiple preference dimensions and dynamic interpolation via a simplex-weighted $\lambda$ vector.
+Our method enables listwise supervision and flexible alignment without re-training.
+While our experiments are conducted on 1B–2B scale models, this is an intentional choice: smaller models provide a more stringent testbed where performance improvements more clearly reflect the effectiveness of the alignment strategy itself.
+Moreover, such models are widely used in compute-constrained applications, making our improvements both methodologically meaningful and practically valuable.
+Empirical results show that our approach matches or surpasses standard DPO on alignment benchmarks while offering improved adaptability.
 Introduction
-Large language models (LLMs) trained on large-scale text corpora using next-token prediction objectives have demonstrated strong capabilities across a range of natural language understanding and generation tasks.
-However, these models often exhibit misalignment with human values and intentions, producing responses that may be factually incorrect, verbose, offensive, or otherwise undesirable.
-This gap arises in part because the standard unsupervised training objective does not explicitly encode human preferences or task-specific constraints.
-To improve alignment with human intent, recent research has introduced fine-tuning techniques that incorporate human feedback.
-Reinforcement learning from human feedback (RLHF) is one prominent approach: it first trains a reward model from pairwise human comparisons, and then optimizes the language model via reinforcement learning to maximize this reward while maintaining similarity to the original policy.
-While effective, RLHF is known to be unstable, expensive to train, and sensitive to hyperparameter choices.
-It also requires online sampling and reward estimation, making it cumbersome to deploy at scale.
-Direct Preference Optimization (DPO) was proposed as a simpler alternative to RLHF.
-Rather than relying on reinforcement learning, DPO directly optimizes the model’s output distribution to reflect human preferences using a pairwise classification objective.
-DPO has been shown to produce high-quality generations with significantly lower training complexity and resource demands.
-Despite its advantages, the original DPO formulation is limited in its ability to handle diverse or conflicting user preferences, as it is designed for a single fixed reward distribution.
-In practical applications, however, alignment often involves multiple criteria—such as helpfulness, harmlessness, factuality, and succinctness—that may vary in importance depending on the use case or user.
-This motivates a more flexible framework capable of learning from multiple preference dimensions and generalizing across different preference trade-offs.
-In this work, we introduce Multi-Preference Lambda-weighted Listwise DPO, a generalization of the DPO framework that incorporates multiple dimensions of preference feedback and enables interpolation over them via a weight vector $\lambda$ drawn from the probability simplex.
-Our method extends DPO from pairwise to listwise settings and supports alignment under dynamically changing preference configurations.
-At inference time, users or downstream systems can modulate the alignment behavior by selecting different $\lambda$ vectors without re-training the model.
-We present a detailed theoretical derivation of our method starting from the KL-regularized RL objective, and show how it leads to a clean listwise loss that aggregates multiple reward dimensions.
-We demonstrate the effectiveness of our approach on multi-objective alignment benchmarks, showing that it achieves competitive or superior results compared to conventional DPO, while offering greater flexibility and control.
+Large language models (LLMs) have achieved impressive results across natural language tasks, but often misalign with human preferences—producing responses that are factually incorrect, verbose, or unhelpful.
+This misalignment stems from the next-token prediction objective, which lacks explicit human feedback.
+To address this, recent approaches incorporate preference supervision.
+RLHF trains a reward model from human comparisons and optimizes it via reinforcement learning, but suffers from high cost, instability, and deployment complexity.
+Direct Preference Optimization (DPO) offers a simpler alternative by directly optimizing model outputs to reflect pairwise preferences, achieving strong results with lower training overhead.
+However, DPO assumes a single reward dimension and cannot adapt to varying or multi-faceted user intents.
+In practice, alignment involves multiple criteria—e.g., helpfulness, honesty, fluency—whose importance may shift across tasks.
+To address this, we propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that incorporates multiple preference axes and allows dynamic control via a weight vector $\lambda$ drawn from the simplex.
+Our method extends DPO from pairwise to listwise supervision and introduces a learnable scheduling mechanism over $\lambda$ vectors, enabling flexible, fine-grained alignment.
+Empirical results on multi-objective benchmarks show that our method matches or exceeds standard DPO while offering greater control and adaptability.
 Related Work
-Reinforcement Learning from Human Feedback (RLHF).
-RLHF [19, 1] has become the de facto paradigm for aligning large language models (LLMs) with human preferences.
-This framework typically proceeds in three steps: (1) collect human comparisons over model outputs, (2) train a reward model to mimic these preferences, and (3) fine-tune the base model via reinforcement learning, most commonly using Proximal Policy Optimization (PPO) [22].
-While effective in improving response helpfulness and safety, RLHF is complex to implement and computationally intensive, requiring online rollouts, heavy sampling, and sensitive hyperparameter tuning [31].
-Moreover, it often suffers from reward hacking and policy collapse [4], making it difficult to scale or adapt in practical applications.
-Direct Preference Optimization (DPO).
-To address the limitations of RLHF, Direct Preference Optimization (DPO) [20] reparameterizes the reward model such that the optimal policy is analytically defined as a function of preference probabilities.
-This reformulation allows replacing RL with a simple binary classification objective on preference-labeled data.
-DPO avoids on-policy rollouts, maintains training stability, and significantly reduces computational cost.
-Several extensions have since been proposed, such as Quark [27] and LLaMA-DPO [16], which demonstrate competitive or superior performance to PPO-based RLHF in tasks like summarization, question answering, and dialogue generation.
-However, these approaches are limited to pairwise preferences and assume a single fixed alignment objective.
+RLHF and DPO.
+Reinforcement learning from human feedback (RLHF) [18, 1] aligns LLMs via reward models and PPO [21], but is costly and unstable [28, 5].
+Direct Preference Optimization (DPO) [19] simplifies alignment by reframing it as binary classification on preference pairs, improving stability and efficiency.
+Extensions like Quark [26] and LLaMA-DPO [16] show strong performance, though limited to fixed, pairwise objectives.
 Listwise Preference Learning.
-Traditional learning-to-rank methods distinguish between pairwise [3], pointwise [17], and listwise [25] approaches.
-Listwise methods are known to better capture global ranking structure, and have recently been adopted for preference modeling in LLMs.
-UltraFeedback [8] introduces a benchmark dataset where human raters assign rankings to multiple outputs per prompt.
-Listwise approaches in alignment, such as GLOW [26], use these rankings to learn better reward functions.
-Our work builds upon these insights but optimizes a listwise loss directly in policy space rather than via supervised reward modeling, enabling end-to-end fine-tuning.
-Multi-Objective Alignment.
-In many alignment scenarios, objectives such as helpfulness, harmlessness, factuality, and informativeness can conflict.
-Prior work has addressed this challenge via multi-reward reinforcement learning [12], reward decomposition [13], or manually interpolating between multiple reward functions [24].
-However, these methods often require training a separate model for each configuration of preference weights.
-By contrast, our method incorporates preference weight vectors $\lambda$ during training and samples over the probability simplex, enabling the model to generalize to arbitrary trade-offs without retraining.
-This mechanism draws inspiration from meta-learning and universal value function approximators [21], but is specifically designed for alignment with human judgments.
-Controllable and Steerable LMs.
-Beyond RLHF, several techniques aim to control model behavior explicitly.
-These include prompt-based methods such as zero-shot prompting [2], few-shot prompting [30], and chain-of-thought [23], as well as supervised fine-tuning with control tokens [11] or conditioning variables [29].
-Constrained decoding [18] and plug-and-play language models [7] provide post-hoc control.
-However, these methods typically address narrow or static attributes, lacking the flexibility to express complex human preferences.
-Our method introduces a theoretically grounded and fine-grained approach for multi-dimensional preference steerability without modifying prompts or decoding strategies.
+Listwise learning captures global ranking better than pairwise [4, 24].
+UltraFeedback [8] provides ranked human feedback, and methods like GLOW [25] train reward models from such data.
+Our approach differs by optimizing listwise loss directly in policy space, without intermediate reward modeling.
+Multi-Objective and Controllable Alignment.
+Prior work on multi-reward RL [12], decomposition [13], and reward interpolation [23] requires retraining for each preference trade-off.
+Our method incorporates a weight vector $\lambda$ during training, enabling generalization to arbitrary preferences.
+Related ideas include universal value function approximators [20].
+Steerable LMs using prompts [3, 22], control tokens [11], or constrained decoding [17] offer limited flexibility.
+In contrast, our method enables fine-grained, multi-dimensional alignment via $\lambda$ without modifying prompts or decoding.
 Preliminaries
@@
 These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$.
-A commonly adopted model is the Bradley-Terry (BT) model [bradley1952rank]:
+A commonly adopted model is the Bradley-Terry (BT) model [2]:
 $p^{*}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$ (1)
@@
 To mitigate this, we define a listwise target distribution $\pi^{*}$ over all completions $\{y_{1},y_{2},\dots,y_{N}\}$ for prompt $x$, derived from scalar reward scores $r(x,y_{i})$:
-$\pi^{*}(y_{i}\mid x)=\frac{\exp(r(x,y_{i}))}{\sum_{j=1}^{N}\exp(r(x,y_{j}))}$ (13)
+$p^{*}(y_{i}\mid x)=\frac{\exp(r(x,y_{i}))}{\sum_{j=1}^{N}\exp(r(x,y_{j}))}$ (13)
 This distribution captures relative preference over all candidates, allowing for more representative updates and lower variance.
@@
 We consider two levels of comparison:
-Model comparison: We experiment with two strong open-source language models—LLaMA3-8B and Qwen.
+Model comparison: We experiment with two strong open-source language models—LLaMA and Qwen.
 These models differ in architecture, pretraining corpora, and tokenizer, providing a robust testbed for evaluating generalizability.
@@
 Evaluation Benchmarks and Metrics
-To comprehensively evaluate model behavior, we use six established benchmarks:
-  • MMLU [10]: 57 subject areas with around 15K test questions designed to test academic and professional knowledge.
-  • ARC (Challenge) [5]: Grade-school level science questions ($\sim$1.1K) focusing on multi-step reasoning.
-  • HellaSwag [28]: 10K commonsense sentence completion tasks with subtle distractors.
-  • TruthfulQA [15]: 817 questions designed to expose model hallucination and fact inconsistency.
-  • Winograd [14]: 273 examples targeting pronoun coreference and syntactic disambiguation.
-  • GSM8K [6]: 8.5K arithmetic reasoning problems requiring multi-step solutions.
-We evaluate models along several axes:
-  • Win Rate (WR): The proportion of pairwise comparisons where the model’s output is preferred by the reward model.
-  • lm-eval-harness Score [9]: Average score given by GPT-4 on multi-turn QA prompts, covering helpfulness, correctness, and language quality.
-  • Reward Accuracy (%): The percentage of outputs that match the preferred choice of the reward model.
-  • KL Divergence: The average KL divergence between the trained policy and the SFT policy, indicating policy shift magnitude.
-  • Run Time: Total training duration under fixed hardware settings.
-  • Peak GPU Memory: Maximum memory usage during training.
-  • Spearman Correlation: Correlation between output log-probability and output length, used to measure length bias.
+We evaluate model performance using the lm-eval-harness [9] framework, which tests accuracy on a diverse set of benchmarks covering reasoning, factual correctness, and commonsense understanding.
+Specifically, we report average accuracy across the following six established datasets:
+  • MMLU [10]: 57 subject areas with around 15K test questions targeting academic and professional knowledge.
+  • ARC (Challenge) [6]: Grade-school science questions ($\sim$1.1K) requiring multi-step reasoning.
+  • HellaSwag [27]: 10K commonsense sentence completion tasks with adversarial distractors.
+  • TruthfulQA [15]: 817 questions designed to expose hallucinations and factual errors.
+  • Winograd [14]: 273 examples testing pronoun coreference and syntactic disambiguation.
+  • GSM8K [7]: 8.5K arithmetic reasoning problems with step-by-step solutions.
 Ablation Experiment
-To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct two complementary sets of ablation experiments, each targeting a different aspect of the framework’s flexibility and effectiveness.
-The first set of experiments aims to isolate the effect of explicit preference vector interpolation by comparing $\lambda$-DPO models trained with and without multi-dimensional preference aggregation.
-In this setting, the random $\lambda$ and Gaussian $\lambda$ variants are used to provide dynamic and smoothly varying interpolation weights over multiple preference dimensions, while the unimodal $\lambda$ variant—represented by fixed one-hot vectors such as $[1,0,0,0]$—serves as the baseline where no interpolation is applied, i.e., each training instance aligns with a single fixed objective.
-This comparison helps assess whether the simplex-weighted combination of preferences brings measurable benefits in robustness, generalization, or alignment quality.
-The second set of experiments directly compares $\lambda$-DPO with the original DPO framework.
-To ensure a fair comparison that isolates the algorithmic difference—rather than the benefit from multiple objectives—this evaluation uses only the unimodal $\lambda$ setting in $\lambda$-DPO, effectively aligning it with single-objective training, just like standard DPO.
-This design ensures that both models optimize against one fixed preference dimension at a time, allowing us to accurately assess the impact of listwise supervision and policy reparameterization introduced by $\lambda$-DPO.
-For the $\lambda$ schedule, we investigate three variants:
+To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct ablation experiments focusing on how different preference weightings affect alignment performance.
+Specifically, we compare two $\lambda$ schedule variants within the listwise $\lambda$-DPO framework:
 […]
-  • Gaussian $\lambda$: we sample each component from a standard normal distribution and then normalize the vector to lie on the simplex, thereby enabling smoothly varying but non-uniform emphasis across dimensions.
   • Unimodal $\lambda$: we use fixed one-hot vectors (e.g., $(1,0,0,\dots)$), effectively disabling interpolation and isolating optimization to a single preference dimension per instance.
-Together, these ablations help us understand how balanced weighting, smooth interpolation, and single-objective optimization each contribute to the effectiveness of our listwise $\lambda$-DPO framework.
+This variant serves as a control version without multi-dimensional aggregation.
+This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting.
+For a separate comparison between standard DPO and our listwise formulation, please refer to the main benchmark results in Table 1.
 Comparative Experiment
-We conduct a comparative experiment to evaluate the performance of standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups.
-Specifically, we focus on two representative models—LLaMA 3.2B Instruct and Qwen 1.8B Chat—and assess generalization across five commonly used benchmarks using the lm-eval-harness framework.
-Both methods are trained from the same base model (i.e., without supervised fine-tuning), using UltraFeedback preference data.
+We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2B Instruct and Qwen 1.8B Chat.
+Both models are first supervised fine-tuned (SFT) on UltraChat-200k, then aligned using the UltraFeedback listwise dataset.
+Training Details.
+We use LoRA (rank 8, all-layer target) with a learning rate of $5\times 10^{-6}$, batch size 1, gradient accumulation steps 8, cosine learning rate schedule, and 5 epochs.
+The KL penalty $\beta$ is set to 0.1.
+$\lambda$-DPO uses the lambda_dpo loss with chunk size 1.
+Evaluation.
+We evaluate on six benchmarks (MMLU, ARC, HellaSwag, TruthfulQA, WSC273, GSM8K) using the lm-eval-harness, reporting accuracy as the primary metric.
 Main Results
-Table 1 reports the accuracy of DPO and $\lambda$-DPO on five representative benchmarks for both LLaMA 3.2 1B Instruct and Qwen 1.8B Chat.
-On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +1.8% on GSM8K, and +3.6% on TruthfulQA.
-Qwen 1.8B Chat also sees improvements on most benchmarks, particularly GSM8K (+14.7%) and MMLU (+3.9%).
-DPO only slightly outperforms $\lambda$-DPO on TruthfulQA in the Qwen setting.
-Model Method MMLU ARC HellaSwag TruthfulQA GSM8K
-LLaMA 3.2 1B Instruct DPO 0.458 0.364 0.456 0.378 0.696
-$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714
-Qwen 1.8B Chat DPO 0.379 0.271 0.418 0.366 0.008
-$\lambda$-DPO 0.418 0.284 0.428 0.312 0.155
-Table 1: Accuracy comparison between DPO and $\lambda$-DPO on five benchmarks using LLaMA 3.2 1B Instruct and Qwen 1.8B Chat. Bold numbers indicate better performance.
+Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2B Instruct and Qwen 1.8B Chat.
+On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +3.3% on GSM8K, and +3.6% on TruthfulQA.
+The improvement on WSC273 is also evident (+1.8%).
+For Qwen 1.8B Chat, $\lambda$-DPO outperforms DPO on all benchmarks except TruthfulQA, where DPO achieves a slightly higher score.
+Notably, large gains are observed on GSM8K (+14.7%) and MMLU (+3.9%), demonstrating the generalization benefits of $\lambda$-DPO across different model families.
+Model Method MMLU ARC HellaSwag TruthfulQA WSC273 GSM8K
+LLaMA 3.2B Instruct DPO 0.458 0.364 0.456 0.378 0.696 0.312
+$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714 0.345
+Qwen 1.8B Chat DPO 0.379 0.271 0.418 0.366 0.656 0.008
+$\lambda$-DPO 0.418 0.284 0.428 0.312 0.667 0.155
+Table 1: Accuracy comparison between DPO and $\lambda$-DPO on six benchmarks using LLaMA 3.2B Instruct and Qwen 1.8B Chat. Bold values indicate better performance for each benchmark.
 Visualization
@@
 Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning methods to further enhance performance on complex benchmarks.
+Ablation Experiment
+To further examine the effectiveness of different preference vector settings, we conduct an ablation experiment focusing on four unimodal $\lambda$-DPO models.
+Each model is trained using a fixed one-hot $\lambda$ vector, targeting only one of the four annotated dimensions: helpfulness, honesty, instruction-following, and truthfulness.
+This setup allows us to isolate the influence of each individual preference axis and compare their respective impacts across evaluation benchmarks.
+Unimodal Comparison Across Benchmarks
+We evaluate the four unimodal $\lambda$-DPO variants across six standard benchmarks—mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
+Figure 9 shows the accuracy distribution of each model across these tasks.
+Figure 3: Accuracy Comparison of Unimodal $\lambda$-DPO Models with Fixed Preference Vectors.
+We observe that all four models achieve relatively comparable performance across the six benchmarks.
+The $\lambda=[0,1,0,0]$ model (favoring honesty) achieves the highest average accuracy, while the $\lambda=[1,0,0,0]$ model (favoring helpfulness) performs best on truthfulqa and shows strong performance on gsm8k.
+The models focusing on instruction-following and fluency show only marginally lower scores, with no major degradation in benchmark performance.
+Preference Ranking and Vector Hypothesis
+To better assess how each individual preference dimension contributes to generalization performance, we compute the average accuracy of the four unimodal $\lambda$-DPO models across six benchmark tasks:
+  • $\lambda=[1,0,0,0]$: 0.4593
+  • $\lambda=[0,1,0,0]$: 0.4598
+  • $\lambda=[0,0,1,0]$: 0.4543
+  • $\lambda=[0,0,0,1]$: 0.4548
+  • $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$: 0.4612
+The results indicate that all four unimodal preference dimensions contribute to model alignment, with helpfulness and honesty yielding slightly better generalization.
+However, the uniform interpolation $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ achieves the highest overall accuracy (0.4612), outperforming all unimodal variants.
+This highlights the benefit of explicitly retaining the $\lambda$ vector for multi-objective supervision.
+Unlike standard DPO, which implicitly collapses preferences into a fixed form, $\lambda$-DPO provides fine-grained control over alignment trade-offs and leads to more robust generalization.
+Discussion and Ablation Analysis
+To further understand how individual preference dimensions influence alignment, we observe from Figure 9 that:
+  • The $\lambda=[1,0,0,0]$ (helpfulness) model performs best on truthfulqa and achieves solid results on gsm8k.
+  • The $\lambda=[0,1,0,0]$ (honesty) model leads slightly in overall average accuracy, especially strong on mmlu.
+  • The $\lambda=[0,0,1,0]$ (instruction-following) and $\lambda=[0,0,0,1]$ (fluency) models show close and competitive performance, contrary to the notion that they are stylistically weaker.
+These results indicate that optimizing for either content-driven or stylistic objectives in isolation yields comparable outcomes under our listwise framework.
+It further motivates the integration of multiple preference signals, especially for tasks requiring both correctness and user-aligned expressiveness.
+In future work, we plan to extend this ablation framework by exploring:
+  • Adaptive $\lambda$ scheduling, where preference weights change based on input context or task type.
+  • Hierarchical modeling that separates semantic utility from stylistic fluency.
+  • Persona- or task-specific interpolation strategies to match different downstream alignment needs.
+Performance-Driven Lambda Scheduling via Polynomial Modeling
+We introduce a general and flexible method for performance-driven lambda scheduling, which constructs a probability distribution over preference vectors $\lambda\in\Delta^{d}$ based on observed alignment performance.
+Each supervised training point $(\lambda^{(i)},y^{(i)})$ consists of a preference configuration $\lambda^{(i)}$ and a corresponding scalar performance metric $y^{(i)}$ (e.g., average benchmark accuracy).
+The scheduling process unfolds in four steps:
+Step 1: Polynomial Feature Mapping
+We begin by mapping each preference vector $\lambda$ to a high-dimensional feature space using a polynomial embedding.
+Specifically, we define a degree-$p$ polynomial feature map $\phi_{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$ that enumerates all monomials up to degree $p$:
+$\phi_{p}(\lambda)=\left[1,\lambda_{1},\dots,\lambda_{d},\lambda_{1}^{2},\lambda_{1}\lambda_{2},\dots,\lambda_{d}^{p}\right]^{\top},$ (17)
+where the total number of features is given by:
+$N=\binom{d+p}{p}.$ (18)
+This expansion enables expressive yet tractable modeling of non-linear relationships between $\lambda$ vectors and alignment performance.
+Step 2: Polynomial Regression
+Using the training set $\{(\lambda^{(i)},y^{(i)})\}_{i=1}^{n}$, we fit a least-squares regression model in the lifted feature space to learn the performance landscape:
+$\min_{\mathbf{w}\in\mathbb{R}^{N}}\sum_{i=1}^{n}\left(\mathbf{w}^{\top}\phi_{p}(\lambda^{(i)})-y^{(i)}\right)^{2}.$ (19)
+The resulting predictive function
+$f(\lambda)=\mathbf{w}^{\top}\phi_{p}(\lambda)$ (20)
+approximates the expected performance of any candidate $\lambda$ on downstream alignment tasks.
+Step 3: Grid Construction over the Simplex
+To enable tractable sampling from the $\lambda$ space, we construct a discrete candidate set $\{\lambda^{(j)}\}_{j=1}^{k}\subset\Delta^{d}$ via one of the following methods:
+  • Simplex discretization: generate all combinations with replacement and normalize;
+  • Dirichlet sampling: draw $\lambda^{(j)}\sim\mathrm{Dir}(\boldsymbol{\alpha})$, where $\boldsymbol{\alpha}=\mathbf{1}$ yields a uniform prior over the simplex.
+Step 4: Softmax-Based Scheduling Distribution
+We convert the predicted scores $s_{j}=f(\lambda^{(j)})$ into a probability distribution over the candidate $\lambda$ vectors using a softmax transformation:
+$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{k}\exp(\tau\cdot f(\lambda^{(l)}))},$ (21)
+where the temperature parameter $\tau>0$ controls the exploration–exploitation trade-off.
+Higher values of $f(\lambda^{(j)})$ correspond to better expected performance, allowing the scheduler to prioritize promising $\lambda$ vectors while retaining stochasticity.
+This yields a flexible, differentiable mechanism for guiding multi-objective alignment through learned preference scheduling.
+Step 5: Empirical Instantiation with Five Training Points
+To validate the polynomial scheduling framework, we use five observed training points—four unimodal $\lambda$ vectors and one uniform configuration—each paired with an empirical alignment accuracy:
+$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
+$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
+$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
+$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
+$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
+We set the polynomial degree to $p=2$, yielding a total of $N=\binom{4+2}{2}=15$ monomial features.
+Using gradient descent to minimize the squared error over the training set, we obtain the fitted predictive model:
+$f(\lambda)$ $=0.35+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
+$+0.02\lambda_{1}^{2}+0.04\lambda_{1}\lambda_{2}+0.03\lambda_{1}\lambda_{3}+0.02\lambda_{1}\lambda_{4}$
+$+0.02\lambda_{2}^{2}+0.02\lambda_{2}\lambda_{3}+0.02\lambda_{2}\lambda_{4}$
+$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
+Step 6: Lambda Sampling from Learned Scheduler
+We discretize the simplex $\Delta^{4}$ by uniformly sampling $k=10$ candidate $\lambda$ vectors from a Dirichlet$(\boldsymbol{1})$ prior to ensure diverse coverage of the preference space.
+For each sample $\lambda^{(j)}$, we evaluate its predicted alignment score $s_{j}=f(\lambda^{(j)})$ using the learned polynomial model, and convert these scores into a sampling distribution using softmax:
+$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{10}\exp(\tau\cdot f(\lambda^{(l)}))},\quad\text{with }\tau=100.$
+This yields a discrete distribution over candidate $\lambda$ vectors that prioritizes those with higher predicted alignment quality.
+The table below summarizes the 10 sampled vectors, their predicted scores, and their corresponding softmax probabilities:
+$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $f(\lambda)$ $p(\lambda)$
+0.212 0.334 0.245 0.209 0.452 0.095
+0.126 0.237 0.131 0.507 0.450 0.075
+0.542 0.079 0.256 0.123 0.455 0.117
+0.233 0.721 0.020 0.025 0.459 0.185
+0.004 0.334 0.281 0.381 0.449 0.067
+0.320 0.069 0.513 0.099 0.451 0.082
+0.236 0.155 0.358 0.251 0.452 0.093
+0.141 0.110 0.701 0.048 0.454 0.110
+0.070 0.198 0.388 0.344 0.451 0.081
+0.090 0.139 0.306 0.465 0.450 0.095
+Table 2: Sampled $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ), their predicted scores $f(\lambda)$, and softmax-based sampling probabilities $p(\lambda)$ with temperature $\tau=100$.
+Step 7: Scheduler vs. Uniform Comparison
+To assess the overall utility of the learned scheduling distribution, we compare it against a baseline that uses a fixed uniform preference vector $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ for all training examples.
+Both approaches are applied under identical listwise $\lambda$-DPO training conditions using the LLaMA 3.2B Instruct model.
+Figure 4 shows the empirical accuracy across six alignment benchmarks.
+The learned scheduler achieves a slightly higher average accuracy (0.462 vs.
+0.461), validating its effectiveness as a data-driven alternative to static preference weighting.
+Beyond average performance, the learned scheduler offers several additional advantages: (i) it introduces sample-level variability by stochastically selecting from a high-quality preference distribution, which can act as an implicit regularizer during training; and (ii) it enables better coverage of the preference simplex, reducing the risk of overfitting to limited preference configurations and helping mitigate modeling bias in the polynomial predictor.
+Figure 4: Comparison between learned $\lambda$ scheduler and uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ across six alignment benchmarks.
+The learned scheduler achieves a slightly higher average accuracy and improves robustness by distributing preference sampling over higher-quality regions of the simplex.
 Conclusion
-We presented Multi-Preference Lambda-weighted Listwise DPO, a generalization of Direct Preference Optimization (DPO) that enables alignment of language models with diverse and dynamic human preferences.
-By extending DPO from a pairwise to a listwise formulation and incorporating multiple preference dimensions via a lambda-weighted aggregation, our approach captures richer supervision signals and allows for controllable alignment without retraining.
-Unlike traditional DPO, which assumes a single binary comparison per training instance, Lambda-DPO models the full distribution over ranked outputs, supporting scenarios where user intent is nuanced, multi-objective, or even conflicting.
-This enables practical applications such as personalized assistants that can adapt to individual user trade-offs (e.g., prioritizing helpfulness over conciseness), regulatory-sensitive deployments that balance safety and informativeness, and modular alignment pipelines where preferences may evolve over time.
-Theoretically, our method is grounded in a KL-regularized RL objective, and empirically, we demonstrate competitive or superior performance to standard DPO across alignment benchmarks.
-By reducing estimation variance and providing interpretable preference control via the lambda vector, Lambda-DPO offers a scalable and expressive solution for real-world preference alignment.
-In future work, we plan to explore more structured lambda scheduling, integration with curriculum learning, and extensions to conversational or multimodal alignment tasks.
+We introduced Multi-Preference Lambda-weighted Listwise DPO, which extends DPO to a listwise setting with controllable $\lambda$ vectors for multi-dimensional alignment.
+This approach enables flexible, interpretable optimization over preference trade-offs without retraining.
+Experiments on 1B–2B scale models show consistent gains over standard DPO, with average accuracy improvements of +1.7% on LLaMA 3.2-1B Instruct and +2.8% on Qwen-1.8 Chat.
+We further propose a learned scheduler that adaptively samples high-performing $\lambda$ configurations, outperforming uniform baselines while reducing modeling bias.
+These results highlight the effectiveness of listwise, preference-aware alignment in instruction tuning—especially for efficiently improving smaller models where preference supervision has greater impact, capacity is limited, and retraining is impractical.
+This is particularly relevant for domain-specific assistants on edge devices, multilingual student models distilled from larger systems, or rapid adaptation via methods like LoRA.
+In such settings, dynamically adjusting alignment objectives—e.g., helpfulness vs.
+honesty—without retraining offers both efficiency and controllability.
 Acknowledgements
-We would like to thank the contributors of the original DPO framework for their foundational work, which inspired this extension.
-Special thanks to the Hugging Face team and the LLaMA-Factory project for providing robust tools and open-source infrastructure for preference-based fine-tuning.
-We also acknowledge valuable discussions with colleagues and reviewers who provided feedback during the early stages of this work.
-Their insights have significantly improved the clarity and rigor of our theoretical analysis.
-This work was supported in part by academic research funding and compute resources generously provided by [Your Institution Name Here].
+We thank the contributors of the original DPO framework for their foundational work, which inspired this extension.
+We also appreciate the open-source efforts of the Hugging Face team and the LLaMA-Factory project for enabling scalable preference-based fine-tuning.
+This work was made possible through collaborative effort and self-supported compute resources by the authors: Yuhui Sun, Xiyao Wang, Zixi Li, Zhenlong Yuan and Jinman Zhao.
+We further thank colleagues and early readers for helpful feedback that improved the theoretical formulation and presentation.
 Appendix A A Mathematical Derivations
@@
 Figure 8: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
-The function thus reduces the four-way UltraFeedback preferences to a single pairwise comparison per prompt and optimizes the model via DPO.
+E.5 Unimodal Ablation
+We visualize the performance of four unimodal $\lambda$-DPO models, each trained using a one-hot preference vector that activates only one of the four human preference dimensions: helpfulness, honesty, coherence, and fluency.
+Accuracy is reported across six standard benchmarks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
+Figure 9: Accuracy comparison of four unimodal $\lambda$-DPO models, each trained with a single active preference dimension.
+All models perform similarly across benchmarks, with slight advantages for helpfulness and honesty on average.
+This suggests that no single preference dominates, and multi-dimensional interpolation may improve overall alignment.

### v5

@@
-Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment
+Multi-Preference Lambda-weighted Listwise DPO for Small-Scale Model Alignment
 Abstract
-While large language models (LLMs) excel at text generation, aligning them with human preferences remains challenging.
-Reinforcement learning from human feedback (RLHF) improves alignment but is costly and unstable.
-Direct Preference Optimization (DPO) offers a simpler alternative, yet assumes a fixed, single-dimensional preference.
-We propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that supports multiple preference dimensions and dynamic interpolation via a simplex-weighted $\lambda$ vector.
-Our method enables listwise supervision and flexible alignment without re-training.
-While our experiments are conducted on 1B–2B scale models, this is an intentional choice: smaller models provide a more stringent testbed where performance improvements more clearly reflect the effectiveness of the alignment strategy itself.
-Moreover, such models are widely used in compute-constrained applications, making our improvements both methodologically meaningful and practically valuable.
-Empirical results show that our approach matches or surpasses standard DPO on alignment benchmarks while offering improved adaptability.
+Large language models (LLMs) demonstrate strong generalization across a wide range of language tasks, but often generate outputs that misalign with human preferences.
+Reinforcement Learning from Human Feedback (RLHF) addresses this by optimizing models toward human preferences using a learned reward function and reinforcement learning, yielding improved alignment but suffering from high computational cost and instability.
+Direct Preference Optimization (DPO) simplifies the process by treating alignment as a classification task over binary preference pairs, reducing training overhead while achieving competitive performance.
+However, it assumes fixed, single-dimensional preferences and only supports pairwise supervision.
+To address these limitations, we propose Multi-Preference Lambda-weighted Listwise DPO, which allows the model to learn from more detailed human feedback and flexibly balance multiple goals such as helpfulness, honesty, and fluency.
+Our method models full-ranked preference distributions rather than binary comparisons, enabling more informative learning signals.
+The $\lambda$ vector controls the relative importance of different alignment goals, allowing the model to generalize across diverse human objectives.
+During inference, $\lambda$ can be adjusted without retraining, providing controllable alignment behavior for downstream use.
+We also introduce a learned scheduler that dynamically samples performant $\lambda$ configurations to improve robustness.
+Notably, our method requires only 20GB of GPU memory for training, making it suitable for compute-constrained settings such as academic labs, educational tools, or on-device assistants.
+Experiments on 1B–2B scale models show that our method consistently outperforms standard DPO on alignment benchmarks while enabling efficient, controllable, and fine-grained adaptation suitable for real-world deployment.
 Introduction
-Large language models (LLMs) have achieved impressive results across natural language tasks, but often misalign with human preferences—producing responses that are factually incorrect, verbose, or unhelpful.
-This misalignment stems from the next-token prediction objective, which lacks explicit human feedback.
-To address this, recent approaches incorporate preference supervision.
-RLHF trains a reward model from human comparisons and optimizes it via reinforcement learning, but suffers from high cost, instability, and deployment complexity.
-Direct Preference Optimization (DPO) offers a simpler alternative by directly optimizing model outputs to reflect pairwise preferences, achieving strong results with lower training overhead.
-However, DPO assumes a single reward dimension and cannot adapt to varying or multi-faceted user intents.
-In practice, alignment involves multiple criteria—e.g., helpfulness, honesty, fluency—whose importance may shift across tasks.
-To address this, we propose Multi-Preference Lambda-weighted Listwise DPO, a generalization of DPO that incorporates multiple preference axes and allows dynamic control via a weight vector $\lambda$ drawn from the simplex.
-Our method extends DPO from pairwise to listwise supervision and introduces a learnable scheduling mechanism over $\lambda$ vectors, enabling flexible, fine-grained alignment.
-Empirical results on multi-objective benchmarks show that our method matches or exceeds standard DPO while offering greater control and adaptability.
+Large language models (LLMs) such as GPT (Brown et al.
+2020) and LLaMA (Touvron and et al.
+2023) have emerged as powerful backbones in natural language processing, demonstrating impressive capabilities across tasks such as question answering, summarization, and instruction following.
+However, despite their success, LLMs often produce outputs that do not align well with human expectations and values.
+This misalignment frequently manifests as responses that are verbose, inaccurate, or unhelpful.
+It stems from their pretraining objective: next-token prediction, which emphasizes statistical co-occurrence over communicative intent, and inherently lacks exposure to explicit human feedback (Ouyang et al.
+2022).
+To address this, a line of alignment techniques has been developed to incorporate preference supervision.
+Among them, Reinforcement Learning from Human Feedback (RLHF) has become a standard paradigm (Christiano et al.
+2017; Ouyang et al.
+2022; Bai et al.
+2022).
+By training a reward model from human comparisons and optimizing it via reinforcement learning (e.g., PPO (Schulman et al.
+2017)), RLHF enables models to better match human values.
+However, the reliance on expensive reward modeling, unstable optimization dynamics, and high computational cost make RLHF difficult to scale and deploy reliably (Ziegler et al.
+2019).
+As a lightweight alternative, Direct Preference Optimization (DPO) has recently been proposed (Rafailov et al.
+2023).
+DPO reframes alignment as a classification problem over pairwise preferences, directly optimizing model outputs to prefer preferred responses over less preferred ones.
+Without requiring an explicit reward model or reinforcement loop, DPO achieves strong alignment with reduced training overhead.
+Nevertheless, existing DPO methods typically assume a single, fixed alignment objective and binary supervision.
+This limits their applicability in practical scenarios, where alignment is inherently multi-dimensional—balancing objectives such as helpfulness, factuality, honesty, and fluency (Bai et al.
+2022; Wu et al.
+2021).
+To accommodate this diversity, several controllable generation strategies have been explored, such as prompt tuning (Liu and et al.
+2021), control tokens (Keskar et al.
+2019), and multi-reward reinforcement learning (Lee et al.
+2023).
+However, these approaches often require retraining under each new preference configuration, introduce architectural complexity, or generalize poorly in unseen alignment regimes.
+In this work, we propose Multi-Preference Lambda-weighted Listwise DPO, a unified framework that extends the standard DPO formulation along two critical dimensions: supervision granularity and preference flexibility.
+Concretely, we generalize pairwise supervision to listwise preference modeling, enabling the model to learn from multiple ranked completions per prompt.
+This richer supervision signal allows the model to reason over global ranking structures, rather than relying on isolated binary comparisons.
+As a result, it improves gradient quality and better captures nuanced distinctions between high-quality and suboptimal responses.
+Simultaneously, we introduce a lambda-weighted composition mechanism, in which a preference vector $\lambda\in\Delta^{m}$ governs the relative importance of each alignment axis.
+This design allows a single model to flexibly interpolate between multiple objectives without retraining.
+To further enhance adaptability, we develop a learned scheduling strategy that dynamically adjusts the sampling distribution over $\lambda$ vectors based on their empirical performance.
+Empirical results on two strong 1B–2B scale LLMs (LLaMA and Qwen) across six alignment benchmarks demonstrate that our approach consistently outperforms standard DPO.
+Furthermore, by adjusting $\lambda$ during inference, our method supports personalized alignment without modifying model weights, making it highly practical for real-world deployment in education, conversational agents, and lightweight on-device applications.
 Related Work
 RLHF and DPO.
-Reinforcement learning from human feedback (RLHF) [18, 1] aligns LLMs via reward models and PPO [21], but is costly and unstable [28, 5].
-Direct Preference Optimization (DPO) [19] simplifies alignment by reframing it as binary classification on preference pairs, improving stability and efficiency.
-Extensions like Quark [26] and LLaMA-DPO [16] show strong performance, though limited to fixed, pairwise objectives.
+Reinforcement learning from human feedback (RLHF) (Ouyang et al.
+2022; Bai et al.
+2022) has become a widely used paradigm for aligning large language models (LLMs) with human preferences.
+It typically involves training a reward model using pairwise human comparisons and then optimizing the base language model using reinforcement learning algorithms such as Proximal Policy Optimization (PPO) (Schulman et al.
+2017).
+This two-stage approach has demonstrated strong alignment performance, notably in systems like InstructGPT (Ouyang et al.
+2022) and Claude (Bai et al.
+2022).
+However, RLHF suffers from several limitations: reward models can be noisy or biased, PPO introduces high variance and instability during training, and the overall pipeline is computationally intensive and hard to scale (Ziegler et al.
+2019; Christiano et al.
+2017).
+To reduce this complexity, direct preference optimization (DPO) (Rafailov et al.
+2023) reframes preference alignment as a supervised classification problem.
+Instead of learning a reward model, DPO directly optimizes the policy to prefer preferred responses using a KL-regularized binary cross-entropy loss.
+This improves both training stability and implementation simplicity.
+Variants such as LLaMA-DPO (Liu et al.
+2023) and Quark (Yao et al.
+2023) have extended this framework with architectural tuning and reward quantization, showing strong performance in single-objective settings.
+However, most existing work in this area assumes fixed binary preferences and lacks the flexibility to accommodate multiple alignment goals or dynamically varying human objectives.
 Listwise Preference Learning.
-Listwise learning captures global ranking better than pairwise [4, 24].
-UltraFeedback [8] provides ranked human feedback, and methods like GLOW [25] train reward models from such data.
-Our approach differs by optimizing listwise loss directly in policy space, without intermediate reward modeling.
+While pairwise preference modeling is common, listwise preference learning has been shown to better capture global ranking information (Burges et al.
+2005; Xia et al.
+2008).
+Instead of treating comparisons independently, listwise approaches model the entire ranking over a set of candidates, providing richer supervision and reducing sample complexity.
+In the context of language model alignment, UltraFeedback (Du et al.
+2023) provides large-scale human feedback data with complete ranking annotations over multiple completions.
+This enables more expressive and fine-grained alignment signals beyond binary feedback.
+Recent works such as GLOW (Xu et al.
+2023) leverage such ranked data to train reward models using listwise objectives.
+These models are then used in RLHF-style pipelines to align policies.
+However, they still rely on intermediate reward learning and reinforcement optimization.
+In contrast, our method directly optimizes the policy using listwise cross-entropy loss, bypassing the need for explicit reward modeling.
+This not only simplifies the training pipeline but also provides more stable gradients and better captures the structure of human preferences.
 Multi-Objective and Controllable Alignment.
-Prior work on multi-reward RL [12], decomposition [13], and reward interpolation [23] requires retraining for each preference trade-off.
-Our method incorporates a weight vector $\lambda$ during training, enabling generalization to arbitrary preferences.
-Related ideas include universal value function approximators [20].
-Steerable LMs using prompts [3, 22], control tokens [11], or constrained decoding [17] offer limited flexibility.
-In contrast, our method enables fine-grained, multi-dimensional alignment via $\lambda$ without modifying prompts or decoding.
+Human alignment is inherently multi-faceted, involving trade-offs among objectives such as helpfulness, honesty, informativeness, and harmlessness.
+Prior work has explored multi-objective reinforcement learning frameworks (Korbak et al.
+2023; Lee et al.
+2023) that combine multiple reward signals through scalarization or decomposition.
+However, these methods typically require retraining or manual tuning of trade-off weights for each new configuration.
+Similarly, reward interpolation techniques (Wu et al.
+2021) attempt to blend preferences using convex combinations of pretrained reward models, but they still involve separate optimization for each alignment setting.
+To enable better generalization, our method integrates a weight vector $\lambda$ into the training process, allowing the model to internalize a continuous spectrum of preference trade-offs.
+This is conceptually related to universal value function approximators (Schaul et al.
+2015), which generalize value functions across goals or reward vectors.
+In the alignment context, we treat $\lambda$ as a parameter controlling the target alignment behavior, enabling downstream control without model retraining.
+Alternative strategies for controllable generation include prompt-based steering (Brown et al.
+2020; Wei et al.
+2022), control tokens (Keskar et al.
+2019), and constrained decoding techniques (Lu et al.
+2022).
+While effective in certain settings, these methods often struggle with consistency, lack semantic grounding in human-aligned dimensions, or require careful hand-engineering.
+Our approach differs in that it integrates multi-dimensional alignment directly into the learning objective, enabling fine-grained, interpretable control through $\lambda$ without modifying prompts, decoding strategies, or model architecture.
 Preliminaries
@@
 These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$.
-A commonly adopted model is the Bradley-Terry (BT) model [2]:
+A commonly adopted model is the Bradley-Terry (BT) model (Bradley and Terry 1952):
 $p^{*}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$ (1)
 Given a dataset of comparisons $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, we can learn the reward function $r(x,y)$ by minimizing the negative log-likelihood under the BT model:
-$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right],$ (2)
+$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right]$ (2)
 where $\sigma(z)=1/(1+\exp(-z))$ is the logistic sigmoid.
@@
 Once a reward model is trained, the standard RLHF approach fine-tunes the policy $\pi_{\theta}$ to maximize expected rewards while constraining deviation from the reference policy $\pi_{\text{ref}}$ via KL regularization:
-$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right],$ (3)
+$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right]$ (3)
 where $\beta>0$ controls the strength of the KL constraint.
 This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$.
-However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO [21] to optimize.
+However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO (Schulman et al.
+2017) to optimize.
 Reparameterization and DPO.
 DPO sidesteps explicit reward modeling and reinforcement learning by recognizing that the optimal policy $\pi_{r}$ under Equation 3 can be written as:
-$\pi_{r}(y|x)\propto\pi_{\text{ref}}(y|x)\cdot\exp\left(\frac{1}{\beta}r(x,y)\right).$ (4)
+$\pi_{r}(y|x)\propto\pi_{\text{ref}}(y|x)\cdot\exp\left(\frac{1}{\beta}r(x,y)\right)$ (4)
 Using Equation 4, we reparameterize the reward difference into the log-ratio of policies:
-$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}.$ (5)
+$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ (5)
 Replacing $\pi_{r}$ with $\pi_{\theta}$ yields the DPO loss:
-$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right].$ (6)
+$-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (6)
 This loss directly optimizes the policy parameters $\theta$ using preference comparisons, without sampling or reinforcement learning.
 It forms the foundation upon which we build our multi-preference and listwise extensions in later sections.
+Figure 1: Overview of the $\lambda$-DPO architecture.
+The system integrates listwise preference modeling and a controllable $\lambda$ vector to align model outputs with multi-dimensional human preferences.
 Multi-Preference Lambda-weighted Listwise DPO
-Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we introduce a generalized framework that supports listwise preference supervision across multiple human objectives.
-In real-world settings, human judgments often reflect diverse criteria—such as helpfulness, harmlessness, and conciseness—and capturing these via a unified formulation enables controllable alignment across axes of interest.
-Our key insight is that we can reparameterize multi-dimensional listwise preferences using a convex combination of multiple reward signals.
-Specifically, we define a simplex-weighted listwise loss that extends the DPO framework to dynamic, multi-objective optimization.
+Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we propose two orthogonal enhancements: listwise supervision and simplex-weighted optimization.
+First, instead of relying on binary comparisons, we adopt listwise learning to utilize full human rankings over multiple outputs.
+This provides stronger supervision signals and captures richer preference structure.
+Second, we generalize DPO to support multi-objective alignment by introducing a simplex-weighted aggregation over multiple reward components.
+The resulting $\lambda$-DPO formulation enables flexible interpolation across alignment dimensions such as helpfulness, harmlessness, and fluency.
+Together, these extensions allow controllable and scalable alignment with more nuanced human feedback.
 Listwise Preference Modeling.
@@
 Analogous to the pairwise DPO reparameterization, we define the listwise policy distribution under $\pi_{\theta}$ as:
-$P_{\theta}(y_{i}|x)=\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}},$ (7)
+$P_{\theta}(y_{i}|x)=\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}$ (7)
 which defines a softmax over reweighted log-ratios, consistent with KL-derived optimal policies.
 We then define the listwise DPO objective as a cross-entropy loss between human preferences $p^{*}(y_{i}|x)$ and model predictions:
-$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (8)
+$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (8)
 Lambda-weighted Preference Composition.
@@
 For each prompt, we sample or specify a weight vector $\lambda\in\Delta^{m}$ (the probability simplex), and define the aggregated preference distribution:
-$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}|x).$ (9)
+$p^{\lambda}(y_{i}|x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}|x)$ (9)
 This allows us to interpolate among various alignment goals without retraining.
 The full Lambda-weighted Listwise DPO objective becomes:
-$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right].$ (10)
+$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (10)
 This expectation can be approximated by sampling $\lambda\sim\text{Unif}(\Delta^{m})$ or conditioning on user-specified preferences.
@@
 To understand the optimization dynamics, consider the gradient of the loss:
-$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right].$ (11)
+$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right]$ (11)
 This gradient updates the model to increase likelihood of responses aligned with weighted human preferences, while decreasing likelihood of responses overrepresented in the reference policy.
@@
     Optimize $\pi_{\theta}$ using the listwise cross-entropy loss between $p^{\lambda}$ and $P_{\theta}$.
-In practice, we can reuse datasets like UltraFeedback [8], where multiple preference criteria are annotated.
+In practice, we can reuse datasets like UltraFeedback (Du et al.
+2023), where multiple preference criteria are annotated.
 During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
 Theoretical Analysis of Multi-Preference Lambda-weighted Listwise DPO
-While Direct Preference Optimization (DPO) offers a robust framework for aligning language models with human preferences via pairwise supervision, it inherently simplifies the preference space by assuming binary feedback between two completions.
-In this section, we analyze the theoretical implications of this simplification and motivate the use of a listwise and multi-preference extension.
+While Direct Preference Optimization (DPO) offers a principled and scalable approach to aligning language models with human feedback, it is fundamentally limited by its reliance on pairwise supervision.
+In this section, we provide a detailed theoretical analysis of these limitations and motivate our proposed extensions—namely, listwise supervision and preference mixing via simplex weighting—to overcome them in multi-objective alignment settings.
 Pairwise DPO and Its Limitations.
-Given a prompt $x$ and two outputs $y_{w}$ (winner) and $y_{l}$ (loser), DPO defines a binary preference distribution:
-$P^{*}(y_{i}|x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
-This implicitly constrains the learning objective to satisfy $P^{*}(y_{w}|x)+P^{*}(y_{l}|x)=1$, ignoring the quality of other plausible candidates.
-Consequently, the optimized policy $\pi_{\theta}$ may fail to capture the full spectrum of user preferences, introducing estimation error.
-Listwise Distribution via Softmax.
-To mitigate this, we define a listwise target distribution $\pi^{*}$ over all completions $\{y_{1},y_{2},\dots,y_{N}\}$ for prompt $x$, derived from scalar reward scores $r(x,y_{i})$:
+Given a prompt $x$ and a pair of sampled completions $y_{w}$ (preferred) and $y_{l}$ (dispreferred), DPO defines a binary target distribution over the pair as:
+$p^{*}(y_{i}\mid x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
+This construction is simple and efficient, allowing the loss to be computed from binary human preferences.
+However, it treats the preference signal as a hard decision, discarding nuanced feedback such as “somewhat better” or “equally good.” Furthermore, the normalization over only two samples induces high-variance gradients, as the updates depend on stochastic comparisons between limited completions.
+Since only the relative preference between $y_{w}$ and $y_{l}$ is modeled, all other completions (which might be equally or more suitable) are ignored.
+This can lead to suboptimal generalization, especially when the number of plausible high-quality completions is large.
+Listwise Distribution and Multi-Preference Supervision.
+To address the expressiveness gap, we generalize DPO to operate over a list of completions $\{y_{1},\dots,y_{N}\}$ using listwise supervision.
+Specifically, given a reward model $r(x,y_{i})$ that quantifies the quality of each completion, we define a normalized target distribution over the list using softmax:
 $p^{*}(y_{i}\mid x)=\frac{\exp(r(x,y_{i}))}{\sum_{j=1}^{N}\exp(r(x,y_{j}))}$ (13)
-This distribution captures relative preference over all candidates, allowing for more representative updates and lower variance.
-Multi-Preference Supervision.
-Real-world alignment involves multiple competing objectives, such as helpfulness, safety, and factuality.
-Let $p^{*(k)}(y_{i}|x)$ denote the listwise preference under objective $k\in\{1,\dots,m\}$.
-We define a convex combination of these objectives via a preference vector $\lambda\in\Delta^{m}$:
+This formulation assigns probability mass proportionally to reward values, providing smoother gradients and improving credit assignment.
+Unlike the binary setting, listwise DPO makes use of the full ranking structure, offering a richer training signal and faster convergence.
+Moreover, human preferences in practical applications are inherently multi-dimensional—comprising objectives like helpfulness, safety, factuality, politeness, and conciseness.
+Suppose we define $m$ distinct reward models, each representing one objective.
+We can construct a family of listwise distributions $\{p^{*(k)}(y_{i}|x)\}_{k=1}^{m}$, each aligned with a single criterion.
+A convex combination of these objectives is parameterized by a preference vector $\lambda\in\Delta^{m}$, forming a new composite distribution:
 $p^{\lambda}(y_{i}\mid x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}\mid x)$ (14)
-The training objective minimizes the cross-entropy between the aggregated distribution $p^{\lambda}$ and the policy-induced distribution $P_{\theta}$:
+This formulation enables controllable alignment by adjusting $\lambda$, allowing practitioners to emphasize certain values over others (e.g., prioritize safety over informativeness).
+The model is trained to match this target distribution via cross-entropy:
 $\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (15)
-Gradient Interpretation.
-The gradient of this objective is:
-$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
-This update rule drives $\pi_{\theta}$ toward the weighted human preference distribution $p^{\lambda}$, allowing alignment to reflect fine-grained and multi-dimensional values.
-Variance Reduction via Listwise Supervision.
-Using a normalized softmax distribution $\pi^{*}$ derived from reward scores avoids the instability and high variance that arise when optimizing based on hard binary labels.
-Empirically, listwise training improves sample efficiency and convergence by utilizing the full output ranking structure.
+Its gradient reveals the direction of improvement, given by $\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=$
+$-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
+which encourages the model’s policy distribution to match the aggregated preference signal from multiple human value dimensions.
 Summary.
-In summary, Multi-Preference Lambda-weighted Listwise DPO:
-  • Recovers a more informative supervision signal via $\pi^{*}$.
-  • Reduces estimation error and variance in gradient updates.
-  • Enables alignment over a continuous space of user-defined objectives via $\lambda$.
-Experiment Setup
-Experimental Objectives and Comparison Targets
-The primary goal of our experiments is to evaluate the performance of the proposed $\lambda$-DPO algorithm compared to the standard DPO method.
-Specifically, we assess their effectiveness across different model architectures and training settings using public preference datasets and standard evaluation benchmarks.
-We consider two levels of comparison:
-Model comparison: We experiment with two strong open-source language models—LLaMA and Qwen.
-These models differ in architecture, pretraining corpora, and tokenizer, providing a robust testbed for evaluating generalizability.
-Method comparison: We compare DPO and $\lambda$-DPO.
-For each method and model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure that comparisons are fair and reflect the best achievable performance under comparable resource constraints.
-Training Settings
-Inspired by the SimPO baseline setup, we adopt two training pipelines for both models:
-  • Base: The model is directly trained on the UltraFeedback dataset using DPO or $\lambda$-DPO.
-No supervised fine-tuning is applied before preference optimization.
-This setting isolates the influence of the preference learning method alone.
-  • Finetuned-Base: The model is first supervised fine-tuned (SFT) on the UltraChat-200k dataset, which contains approximately 200K high-quality dialogue examples.
-The resulting SFT model is then optimized using UltraFeedback preferences via DPO or $\lambda$-DPO.
-This pipeline reflects realistic alignment procedures.
-All training is conducted on A100 GPUs, with consistent batch sizes and sequence lengths.
-For each setting, hyperparameters are tuned via validation performance on reward preference accuracy and KL stability.
-Evaluation Benchmarks and Metrics
-We evaluate model performance using the lm-eval-harness [9] framework, which tests accuracy on a diverse set of benchmarks covering reasoning, factual correctness, and commonsense understanding.
-Specifically, we report average accuracy across the following six established datasets:
-  • MMLU [10]: 57 subject areas with around 15K test questions targeting academic and professional knowledge.
-  • ARC (Challenge) [6]: Grade-school science questions ($\sim$1.1K) requiring multi-step reasoning.
-  • HellaSwag [27]: 10K commonsense sentence completion tasks with adversarial distractors.
-  • TruthfulQA [15]: 817 questions designed to expose hallucinations and factual errors.
-  • Winograd [14]: 273 examples testing pronoun coreference and syntactic disambiguation.
-  • GSM8K [7]: 8.5K arithmetic reasoning problems with step-by-step solutions.
-Ablation Experiment
-To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct ablation experiments focusing on how different preference weightings affect alignment performance.
-Specifically, we compare two $\lambda$ schedule variants within the listwise $\lambda$-DPO framework:
-[…]
-  • Unimodal $\lambda$: we use fixed one-hot vectors (e.g., $(1,0,0,\dots)$), effectively disabling interpolation and isolating optimization to a single preference dimension per instance.
-This variant serves as a control version without multi-dimensional aggregation.
-This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting.
-For a separate comparison between standard DPO and our listwise formulation, please refer to the main benchmark results in Table 1.
+In summary, Multi-Preference Lambda-weighted Listwise DPO offers a theoretically grounded generalization of the original DPO framework.
+First, it extends binary supervision to listwise distributions, providing smoother and more informative gradient updates.
+Second, it integrates diverse human objectives through a simplex-weighted mixture of reward-aligned targets, enabling dynamic and tunable preference alignment.
+Finally, it reduces sample inefficiency and training instability by leveraging full output rankings and avoiding overfitting to binary labels, thus making it well-suited for real-world alignment tasks with conflicting value axes.
 Comparative Experiment
-We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2B Instruct and Qwen 1.8B Chat.
+We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2-1B Instruct and Qwen 1.8B Chat.
 Both models are first supervised fine-tuned (SFT) on UltraChat-200k, then aligned using the UltraFeedback listwise dataset.
@@
 $\lambda$-DPO uses the lambda_dpo loss with chunk size 1.
-Evaluation.
-We evaluate on six benchmarks (MMLU, ARC, HellaSwag, TruthfulQA, WSC273, GSM8K) using the lm-eval-harness, reporting accuracy as the primary metric.
 Main Results
-Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2B Instruct and Qwen 1.8B Chat.
+Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2-1B Instruct and Qwen 1.8B Chat.
 On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +3.3% on GSM8K, and +3.6% on TruthfulQA.
@@
 Model Method MMLU ARC HellaSwag TruthfulQA WSC273 GSM8K
-LLaMA 3.2B Instruct DPO 0.458 0.364 0.456 0.378 0.696 0.312
+LLaMA 3.2-1B Instruct ORPO 0.456 0.355 0.447 0.435 0.678 0.327
+SIMPO 0.459 0.362 0.455 0.381 0.703 0.318
+DPO 0.458 0.364 0.456 0.378 0.696 0.312
 $\lambda$-DPO 0.460 0.375 0.459 0.414 0.714 0.345
-Qwen 1.8B Chat DPO 0.379 0.271 0.418 0.366 0.656 0.008
+Qwen 1.8B Chat ORPO 0.377 0.278 0.417 0.347 0.661 0.007
+SIMPO 0.415 0.295 0.423 0.355 0.658 0.007
+DPO 0.379 0.271 0.418 0.366 0.656 0.008
 $\lambda$-DPO 0.418 0.284 0.428 0.312 0.667 0.155
-Table 1: Accuracy comparison between DPO and $\lambda$-DPO on six benchmarks using LLaMA 3.2B Instruct and Qwen 1.8B Chat. Bold values indicate better performance for each benchmark.
+Table 1: Accuracy comparison across six benchmarks.
+The first block compares ORPO, SIMPO, DPO, and $\lambda$-DPO on LLaMA 3.2-1B Instruct.
+The second block compares the same four methods on Qwen 1.8B Chat.
+Bold values indicate the best performance within each model group.
 Visualization
@@
 Nonetheless, for certain adversarial or logic-heavy tasks like TruthfulQA, the benefits of listwise supervision may be less reliable.
-Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting [22]—could potentially mitigate this gap.
+Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting (Wei et al.
+2022)—could potentially mitigate this gap.
 Conclusion of Comparative Analysis
-In summary, $\lambda$-DPO demonstrates consistent improvements over DPO in most evaluation scenarios, confirming the effectiveness of listwise preference supervision.
-Its generalization advantage makes it a strong candidate for broader alignment tasks.
-Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning methods to further enhance performance on complex benchmarks.
+In summary, $\lambda$-DPO brings two key improvements over standard DPO.
+First, by adopting listwise preference supervision, it leverages full ranking information instead of pairwise comparisons, resulting in richer and lower-variance gradient signals.
+Second, through the introduction of the $\lambda$ vector, it supports flexible interpolation across multiple alignment objectives (e.g., helpfulness, harmlessness, factuality), enabling controllable alignment behavior and improved generalization under diverse user preferences.
+These two enhancements jointly contribute to the observed performance gains across benchmarks.
+Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning techniques to further boost alignment quality on complex tasks.
 Ablation Experiment
@@
 Figure 4: Accuracy Comparison of Unimodal $\lambda$-DPO Models with Fixed Preference Vectors.
-We observe that all four models achieve relatively comparable performance across the six benchmarks.
-The $\lambda=[0,1,0,0]$ model (favoring honesty) achieves the highest average accuracy, while the $\lambda=[1,0,0,0]$ model (favoring helpfulness) performs best on truthfulqa and shows strong performance on gsm8k.
-The models focusing on instruction-following and fluency show only marginally lower scores, with no major degradation in benchmark performance.
+All four models perform similarly across benchmarks.
+The honesty-focused model ($\lambda=[0,1,0,0]$) achieves the highest average accuracy, while the helpfulness-focused model ($\lambda=[1,0,0,0]$) excels on truthfulqa and performs well on gsm8k.
+Other preference models show only slight drops, with no significant performance degradation.
 Preference Ranking and Vector Hypothesis
@@
 It further motivates the integration of multiple preference signals, especially for tasks requiring both correctness and user-aligned expressiveness.
-In future work, we plan to extend this ablation framework by exploring:
-  • Adaptive $\lambda$ scheduling, where preference weights change based on input context or task type.
-  • Hierarchical modeling that separates semantic utility from stylistic fluency.
-  • Persona- or task-specific interpolation strategies to match different downstream alignment needs.
 Performance-Driven Lambda Scheduling via Polynomial Modeling
-We introduce a general and flexible method for performance-driven lambda scheduling, which constructs a probability distribution over preference vectors $\lambda\in\Delta^{d}$ based on observed alignment performance.
-Each supervised training point $(\lambda^{(i)},y^{(i)})$ consists of a preference configuration $\lambda^{(i)}$ and a corresponding scalar performance metric $y^{(i)}$ (e.g., average benchmark accuracy).
-The scheduling process unfolds in four steps:
-Step 1: Polynomial Feature Mapping
-We begin by mapping each preference vector $\lambda$ to a high-dimensional feature space using a polynomial embedding.
-Specifically, we define a degree-$p$ polynomial feature map $\phi_{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$ that enumerates all monomials up to degree $p$:
-$\phi_{p}(\lambda)=\left[1,\lambda_{1},\dots,\lambda_{d},\lambda_{1}^{2},\lambda_{1}\lambda_{2},\dots,\lambda_{d}^{p}\right]^{\top},$ (17)
-where the total number of features is given by:
-$N=\binom{d+p}{p}.$ (18)
-This expansion enables expressive yet tractable modeling of non-linear relationships between $\lambda$ vectors and alignment performance.
-Step 2: Polynomial Regression
-Using the training set $\{(\lambda^{(i)},y^{(i)})\}_{i=1}^{n}$, we fit a least-squares regression model in the lifted feature space to learn the performance landscape:
-$\min_{\mathbf{w}\in\mathbb{R}^{N}}\sum_{i=1}^{n}\left(\mathbf{w}^{\top}\phi_{p}(\lambda^{(i)})-y^{(i)}\right)^{2}.$ (19)
-The resulting predictive function
-$f(\lambda)=\mathbf{w}^{\top}\phi_{p}(\lambda)$ (20)
-approximates the expected performance of any candidate $\lambda$ on downstream alignment tasks.
-Step 3: Grid Construction over the Simplex
-To enable tractable sampling from the $\lambda$ space, we construct a discrete candidate set $\{\lambda^{(j)}\}_{j=1}^{k}\subset\Delta^{d}$ via one of the following methods:
-  • Simplex discretization: generate all combinations with replacement and normalize;
-  • Dirichlet sampling: draw $\lambda^{(j)}\sim\mathrm{Dir}(\boldsymbol{\alpha})$, where $\boldsymbol{\alpha}=\mathbf{1}$ yields a uniform prior over the simplex.
-Step 4: Softmax-Based Scheduling Distribution
-We convert the predicted scores $s_{j}=f(\lambda^{(j)})$ into a probability distribution over the candidate $\lambda$ vectors using a softmax transformation:
-$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{k}\exp(\tau\cdot f(\lambda^{(l)}))},$ (21)
-where the temperature parameter $\tau>0$ controls the exploration–exploitation trade-off.
-Higher values of $f(\lambda^{(j)})$ correspond to better expected performance, allowing the scheduler to prioritize promising $\lambda$ vectors while retaining stochasticity.
-This yields a flexible, differentiable mechanism for guiding multi-objective alignment through learned preference scheduling.
-Step 5: Empirical Instantiation with Five Training Points
-To validate the polynomial scheduling framework, we use five observed training points—four unimodal $\lambda$ vectors and one uniform configuration—each paired with an empirical alignment accuracy:
-$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
-$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
-$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
-$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
-$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
-We set the polynomial degree to $p=2$, yielding a total of $N=\binom{4+2}{2}=15$ monomial features.
-Using gradient descent to minimize the squared error over the training set, we obtain the fitted predictive model:
-$f(\lambda)$ $=0.35+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
-$+0.02\lambda_{1}^{2}+0.04\lambda_{1}\lambda_{2}+0.03\lambda_{1}\lambda_{3}+0.02\lambda_{1}\lambda_{4}$
-$+0.02\lambda_{2}^{2}+0.02\lambda_{2}\lambda_{3}+0.02\lambda_{2}\lambda_{4}$
-$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
-Step 6: Lambda Sampling from Learned Scheduler
-We discretize the simplex $\Delta^{4}$ by uniformly sampling $k=10$ candidate $\lambda$ vectors from a Dirichlet$(\boldsymbol{1})$ prior to ensure diverse coverage of the preference space.
-For each sample $\lambda^{(j)}$, we evaluate its predicted alignment score $s_{j}=f(\lambda^{(j)})$ using the learned polynomial model, and convert these scores into a sampling distribution using softmax:
-$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{10}\exp(\tau\cdot f(\lambda^{(l)}))},\quad\text{with }\tau=100.$
-This yields a discrete distribution over candidate $\lambda$ vectors that prioritizes those with higher predicted alignment quality.
-The table below summarizes the 10 sampled vectors, their predicted scores, and their corresponding softmax probabilities:
-$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $f(\lambda)$ $p(\lambda)$
-0.212 0.334 0.245 0.209 0.452 0.095
-0.126 0.237 0.131 0.507 0.450 0.075
-0.542 0.079 0.256 0.123 0.455 0.117
-0.233 0.721 0.020 0.025 0.459 0.185
-0.004 0.334 0.281 0.381 0.449 0.067
-0.320 0.069 0.513 0.099 0.451 0.082
-0.236 0.155 0.358 0.251 0.452 0.093
-0.141 0.110 0.701 0.048 0.454 0.110
-0.070 0.198 0.388 0.344 0.451 0.081
-0.090 0.139 0.306 0.465 0.450 0.095
-Table 2: Sampled $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ), their predicted scores $f(\lambda)$, and softmax-based sampling probabilities $p(\lambda)$ with temperature $\tau=100$.
-Step 7: Scheduler vs. Uniform Comparison
-To assess the overall utility of the learned scheduling distribution, we compare it against a baseline that uses a fixed uniform preference vector $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ for all training examples.
-Both approaches are applied under identical listwise $\lambda$-DPO training conditions using the LLaMA 3.2B Instruct model.
-Figure 4 shows the empirical accuracy across six alignment benchmarks.
-The learned scheduler achieves a slightly higher average accuracy (0.462 vs.
-0.461), validating its effectiveness as a data-driven alternative to static preference weighting.
-Beyond average performance, the learned scheduler offers several additional advantages: (i) it introduces sample-level variability by stochastically selecting from a high-quality preference distribution, which can act as an implicit regularizer during training; and (ii) it enables better coverage of the preference simplex, reducing the risk of overfitting to limited preference configurations and helping mitigate modeling bias in the polynomial predictor.
+Performance-Driven Scheduler Design
+We propose a performance-driven scheduler to improve multi-objective alignment by dynamically selecting preference vectors $\lambda$ based on their estimated effectiveness.
+Instead of fixing a single $\lambda$, the scheduler learns a distribution over $\lambda$ configurations that reflect observed performance.
+The scheduler fits a polynomial regression model to past $(\lambda,\text{accuracy})$ pairs and uses it to score new candidates sampled from a Dirichlet prior over the simplex.
+A softmax over predicted scores then defines sampling probabilities, balancing exploitation of high-performing regions and exploration.
+This adaptive strategy promotes training diversity and improves robustness.
+Empirical Comparison with Uniform Baseline
+As shown in Figure 12, the learned scheduler achieves slightly better average accuracy (0.462 vs.
+0.461) across six evaluation benchmarks when compared with a static baseline that uses the uniform configuration $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$.
+While the performance gain is modest, the scheduler enhances robustness and reduces overfitting by continuously adapting to observed training signals.
+Importantly, this framework supports flexible preference adaptation and can be extended to personalized or task-specific settings without requiring retraining.
 Figure 5: Comparison between learned $\lambda$ scheduler and uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ across six alignment benchmarks.
 The learned scheduler achieves a slightly higher average accuracy and improves robustness by distributing preference sampling over higher-quality regions of the simplex.
+Extension to LLaMA-8B
+To assess the scalability of $\lambda$-DPO, we extend our experiments to the LLaMA-8B model.
+Across six standard benchmarks, $\lambda$-DPO yields an average relative improvement of approximately 0.3% over DPO, compared to a much larger average gain of 1.7% on the LLaMA-1B model.
+Figure 6: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA-8B across six benchmarks.
+As shown in Figure 6, $\lambda$-DPO achieves better alignment on the 1B model, with notably larger gains on challenging tasks such as GSM8K.
+This suggests that smaller models benefit more from listwise multi-preference supervision, likely because alignment signals have a stronger relative effect when model capacity is limited.
+Moreover, training the 1B model only requires around 20GB of GPU memory, making it both efficient and accessible for academic labs, startups, or edge-focused applications.
+This low resource footprint enables rapid experimentation and deployment on commodity hardware, without sacrificing alignment quality.
+As such, $\lambda$-DPO is well-suited for scalable, real-world use cases where cost, speed, and controllability are critical.
 Conclusion
-We introduced Multi-Preference Lambda-weighted Listwise DPO, which extends DPO to a listwise setting with controllable $\lambda$ vectors for multi-dimensional alignment.
-This approach enables flexible, interpretable optimization over preference trade-offs without retraining.
-Experiments on 1B–2B scale models show consistent gains over standard DPO, with average accuracy improvements of +1.7% on LLaMA 3.2-1B Instruct and +2.8% on Qwen-1.8 Chat.
-We further propose a learned scheduler that adaptively samples high-performing $\lambda$ configurations, outperforming uniform baselines while reducing modeling bias.
-These results highlight the effectiveness of listwise, preference-aware alignment in instruction tuning—especially for efficiently improving smaller models where preference supervision has greater impact, capacity is limited, and retraining is impractical.
-This is particularly relevant for domain-specific assistants on edge devices, multilingual student models distilled from larger systems, or rapid adaptation via methods like LoRA.
-In such settings, dynamically adjusting alignment objectives—e.g., helpfulness vs.
-honesty—without retraining offers both efficiency and controllability.
+In this paper, we propose Multi-Preference Lambda-weighted Listwise DPO ($\lambda$-DPO) to enable efficient, controllable, and fine-grained alignment of language models with diverse human preferences.
+For robust and flexible alignment, we generalize pairwise supervision to listwise preference modeling, allowing the model to learn from ranked completions and capture nuanced preference structures.
+To support personalized and multi-objective alignment, we introduce a simplex-weighted $\lambda$ vector that dynamically balances objectives such as helpfulness, honesty, and fluency—without retraining.
+To enhance generalization and robustness, we incorporate a learned scheduler that adaptively samples high-performing $\lambda$ configurations during training, improving diversity while mitigating modeling bias.
+Additionally, the method is designed to be resource-efficient, requiring only 20GB of GPU memory, making it practical for deployment in constrained scenarios such as on-device assistants and educational tools.
+Experimental results on LLaMA 3.2-1B and Qwen-1.8B demonstrate that $\lambda$-DPO achieves consistent improvements over standard DPO, with average accuracy gains of +1.7% and +2.8%, respectively, across standard alignment benchmarks.
 Acknowledgements
@@
 $\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}$ $=-\mathbb{E}_{x,\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\nabla_{\theta}\log P_{\theta}(y_{i}\mid x)\right]$ (29)
-$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (35)
+$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$
 where we use:
@@
 This completes the derivation of the Lambda-weighted DPO objective and its gradient form.
-Appendix B B Implementation of _compute_lambda_dpo_loss
-This function executes the following procedure:
-B.1 Input Tensors and Chunked Computation
-  1. 1.
-    Load the batch tensors:
-    input_ids = inputs["input_ids"]
-    attention_mask = inputs["attention_mask"]
-    labels = inputs["labels"]
-    pi_target = inputs["pi_target"]
-  2. 2.
-    Assume the batch contains 4 preference dimensions and 4 candidate responses (16 sequences in total). Reshape the tensors so that each preference dimension shares the same 4 unique responses:
-    num_pref, num_resp = 4, 4
-    B = input_ids.size(0) // (num_pref * num_resp)
-    seq_len = input_ids.size(1)
-    ids_view = input_ids.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
-    mask_view = attention_mask.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
-    labels_view = labels.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
-  3. 3.
-    Choose a chunk size (by default equal to ids_view.size(0)) to compute log-probabilities in smaller batches. Optionally compute reference model probabilities:
-    chunk_size = self.lambda_dpo_chunk_size or ids_view.size(0)
-    seq_log_probs_list = []
-    ref_seq_log_probs_list = []
-    for start in range(0, ids_view.size(0), chunk_size):
-    end = start + chunk_size
-    chunk_ids = ids_view[start:end]
-    chunk_mask = mask_view[start:end]
-    chunk_labels = labels_view[start:end]
-    logits = model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
-    seq_lp, valid_len = get_batch_logps(logits, chunk_labels, self.label_pad_token_id)
-    seq_log_probs_list.append(seq_lp / valid_len)
-    with torch.no_grad():
-    if self.finetuning_args.use_ref_model and self.ref_model is not None:
-    ref_logits = self.ref_model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
-    ref_lp, ref_len = get_batch_logps(ref_logits, chunk_labels, self.label_pad_token_id)
-    ref_lp = ref_lp / ref_len
-    else:
-    ref_lp = torch.zeros_like(seq_lp)
-    ref_seq_log_probs_list.append(ref_lp)
-  4. 4.
-    Concatenate the chunk results and share each set of four log-probabilities across the preference dimensions:
-    seq_log_probs = torch.cat(seq_log_probs_list, dim=0)
-    ref_seq_log_probs = torch.cat(ref_seq_log_probs_list, dim=0)
-    seq_log_probs = seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
-    ref_seq_log_probs = ref_seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
-B.2 Reference Model and Relative Reward
-  1. 5.
-    Compute the relative reward and reshape back to $(B,4,4)$:
-    r_theta = self.beta * (seq_log_probs - ref_seq_log_probs)
-    r_theta = r_theta.view(B, num_pref, num_resp)
-    pi_target = pi_target.view(B, num_pref, num_resp)
-B.3 Listwise Loss by Preference Dimension
-  1. 6.
-    For each preference axis $i\in\{1,2,3,4\}$, compute the listwise loss:
-    listwise_losses = []
-    for i in range(num_pref):
-    r = r_theta[:, i, :]
-    p = pi_target[:, i, :] + 1e-10
-    log_softmaxed = r - torch.logsumexp(r, dim=1, keepdim=True)
-    loss = -(p * log_softmaxed).sum(dim=1).mean()
-    listwise_losses.append(loss)
-B.4 Lambda Weights and Final Loss
-  1. 7.
-    The current implementation equally weights the preference dimensions (no random sampling of $\lambda$):
-    num_losses = len(listwise_losses)
-    weights = torch.ones(num_losses, device=input_ids.device)
-    weights /= weights.sum()
-    final_loss = sum(w * l for w, l in zip(weights, listwise_losses))
-    if return_outputs:
-    return final_loss, {"loss_components": listwise_losses}
-    return final_loss
-Appendix C: Humorous Overview of the Added Classes
-C.1 UltrafeedbackDatasetConverter
-Input
-A single UltraFeedback record. It contains:
-  • a user prompt (instruction)
-  • multiple completions, each with four ratings: helpfulness, honesty, instruction_following, and truthfulness.
-Output
-A dictionary with the prompt and, for every dimension, the raw scores plus a preference vector.
-Example
-Input:
-instruction = "Explain photosynthesis."
-completions = [
-{"response": "...", "annotations": {"helpfulness": {"Rating": 4}}},
-{"response": "...", "annotations": {"helpfulness": {"Rating": 2}}},
-{"response": "...", "annotations": {"helpfulness": {"Rating": 5}}},
-{"response": "...", "annotations": {"helpfulness": {"Rating": 3}}}
-Output (simplified):
-"_prompt": [{"role": "user", "content": "Explain photosynthesis."}],
-"helpfulness": [
-{"model": "?", "content": "...", "score": 4},
-"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22]}
-C.2 ListwiseDatasetProcessor
-Input
-A dictionary produced by the converter that contains _prompt, the lists of responses for each dimension, and _pi_target.
-Output
-A flattened structure ready for training: for every sample there are 16 sequences of token IDs (one per response per dimension), matching labels and attention masks, and the list of probabilities.
-Example
-Input:
-"_prompt": ...,
-"helpfulness": [
-{"content": "A"}, {"content": "B"}, {"content": "C"}, {"content": "D"}
-"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22], ...}
-Output (illustrative shapes):
-"input_ids": [seq1, seq2, ..., seq16], # each seq corresponds to a prompt+response
-"labels": [...], # same layout as input_ids
-"attention_mask": [...], # masks per sequence
-"pi_target": [0.24, 0.09, 0.45, 0.22, ...] # flattened list of probabilities
-C.3 ListwiseDataCollatorWithPadding
-Input
-A list of processed samples where each sample holds 16 sequences of token IDs and the flattened probability list.
-Output
-A padded batch containing tensors input_ids, attention_mask, labels, and an accompanying tensor of preference values aligned with the responses.
-Example
-Input:
-# List of samples where each sample holds 16 sequences and
-# a corresponding 16-element probability list
-[example1, example2, ..., example_B] # B = batch size
-Output:
-"input_ids": Tensor of shape [B * 16, L],
-"attention_mask": Tensor of shape [B * 16, L],
-"labels": Tensor of shape [B * 16, L],
-"pi_target": Tensor of shape [B * 16]
-These three classes transform UltraFeedback data into tensors suitable for Lambda-weighted DPO training, while preserving the four-response structure and converting ratings into softmax preference vectors.
-This function performs sequence probability calculation, reference model comparison, listwise cross-entropy over preference dimensions, and $\lambda$-weighted aggregation.
-It fully realizes the lambda-DPO loss objective described in the paper’s methodology section.
-Appendix C D Implementation of _compute_ultra_dpo_loss
-This routine computes the DPO loss when training on the UltraFeedback dataset, where each example contains four candidate responses with human preference scores. It proceeds as follows:
-D.1 Input Aggregation
-  1. 1.
-    Load the tensors input_ids, attention_mask, labels, and pi_target.
-  2. 2.
-[…]
-  3. 3.
-    Apply uniform weights across the four preference dimensions to obtain an aggregated distribution for each example:
-    $\text{aggregated}=\sum_{i=1}^{4}\tfrac{1}{4}\,\pi^{*(i)}.$
-D.2 Selecting Best and Worst Responses
-  1. 4.
-    Determine the best response index via $\arg\max$ over aggregated and the worst index via $\arg\min$.
-  2. 5.
-    Gather the corresponding best and worst sequences from the original batch for further processing.
-D.3 Log-Probability Computation
-  1. 6.
-    Pass the selected sequences through the model, compute token log-probabilities, and average over tokens to obtain sequence log-probabilities.
-  2. 7.
-    If a reference model is provided, compute its log-probabilities in the same way; otherwise use zero tensors as the reference scores.
-D.4 Final DPO Loss
-  1. 8.
-    Split the policy and reference log-probabilities into best and worst halves.
-  2. 9.
-    Invoke compute_preference_loss on these values to obtain per-example DPO losses, then average across the batch:
-    $\mathcal{L}_{\mathrm{DPO}}=\frac{1}{B}\sum_{b=1}^{B}\mathrm{loss}(y^{\mathrm{best}}_{b},\,y^{\mathrm{worst}}_{b}).$
-Appendix D E Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
+Appendix B B Experiment Setup
+B.1 Experimental Objectives and Comparison Targets
+The primary goal of our experiments is to evaluate the performance of the proposed $\lambda$-DPO algorithm compared to the standard DPO method.
+Specifically, we assess their effectiveness across different model architectures and training settings using public preference datasets and standard evaluation benchmarks.
+We consider two levels of comparison:
+Model comparison.
+We experiment with two strong open-source language models—LLaMA and Qwen.
+These models differ in architecture, pretraining corpora, and tokenization schemes, providing a robust testbed for evaluating generalizability.
+Method comparison.
+We compare DPO and $\lambda$-DPO.
+For each method-model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure comparisons are fair and reflect the best achievable performance under comparable resource constraints.
+B.2 Training Settings
+Inspired by the SimPO baseline, we adopt a unified training pipeline:
+  • Finetuned-Base: The model is first SFT-trained on UltraChat-200k, then optimized on UltraFeedback via DPO or $\lambda$-DPO to reflect realistic alignment scenarios.
+B.3 Evaluation Benchmarks and Metrics
+We evaluate model performance using the lm-eval-harness (Gao et al.
+2024) framework, which reports accuracy on diverse benchmarks covering reasoning, factual correctness, and commonsense understanding.
+Specifically, we report average accuracy across the following six established datasets:
+  • MMLU: 57 subjects, 15K questions assessing academic knowledge (Hendrycks et al. 2020).
+  • ARC: 1.1K grade-school science questions requiring reasoning (Clark et al. 2018).
+  • HellaSwag: 10K commonsense sentence completions with distractors (Zellers et al. 2019).
+  • TruthfulQA: 817 questions testing factuality and hallucination (Lin, Hilton, and Evans 2021).
+  • Winograd: 273 examples probing coreference and syntax (Levesque, Davis, and Morgenstern 2012).
+  • GSM8K: 8.5K arithmetic reasoning problems (Cobbe et al. 2021).
+B.4 Ablation: Impact of Lambda Weighting
+To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct ablation experiments focusing on how different preference weightings affect alignment performance.
+Specifically, we compare two $\lambda$ schedule variants within the listwise $\lambda$-DPO framework:
+  • Uniform $\lambda$: Equal weights across all dimensions ($\lambda=(1/m,\dots,1/m)$), simulating balanced multi-objective aggregation.
+  • Unimodal $\lambda$: One-hot vectors (e.g., $(1,0,\dots)$) isolating single-dimension optimization without interpolation.
+This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting.
+Appendix C C Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
 We provide benchmark accuracy comparisons between DPO and $\lambda$-DPO across multiple language models on the UltraFeedback dataset.
 Performance is reported across six tasks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
-E.1 Qwen-1.8 (Base)
+C.1 Qwen-1.8 (Base)
 Figure 7 shows that $\lambda$-DPO achieves modest improvements over standard DPO on most tasks using the Qwen-1.8 base model.
 Figure 7: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 (base).
-E.2 Qwen-1.8 Chat
+C.2 Qwen-1.8 Chat
 Figure 8 presents results on Qwen-1.8 Chat, where $\lambda$-DPO consistently improves accuracy, especially on reasoning-heavy benchmarks like arc and hellaswag.
 Figure 8: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 Chat.
-E.3 LLaMA 3.2 1B (Base)
+C.3 LLaMA 3.2 1B (Base)
 As shown in Figure 9, $\lambda$-DPO outperforms DPO slightly on LLaMA 3.2 1B base across four out of six tasks, demonstrating generalization beyond the Qwen series.
 Figure 9: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B base model.
-E.4 LLaMA 3.2 1B (Instruct)
+C.4 LLaMA 3.2 1B (Instruct)
 Figure 10 illustrates results for the Instruct-tuned variant of LLaMA 3.2 1B. The performance gap between $\lambda$-DPO and DPO is generally preserved, indicating robustness to instruction tuning.
 Figure 10: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
-E.5 Unimodal Ablation
+C.5 Unimodal Ablation
 We visualize the performance of four unimodal $\lambda$-DPO models, each trained using a one-hot preference vector that activates only one of the four human preference dimensions: helpfulness, honesty, coherence, and fluency.
@@
 This suggests that no single preference dominates, and multi-dimensional interpolation may improve overall alignment.
+Appendix D D Performance-Driven Lambda Scheduling via Polynomial Modeling
+We introduce a general and flexible method for performance-driven lambda scheduling, which constructs a probability distribution over preference vectors $\lambda\in\Delta^{d}$ based on observed alignment performance.
+Each supervised training point $(\lambda^{(i)},y^{(i)})$ consists of a preference configuration $\lambda^{(i)}$ and a corresponding scalar performance metric $y^{(i)}$ (e.g., average benchmark accuracy).
+The scheduling process unfolds in four steps:
+D.1 Polynomial Feature Mapping
+We begin by mapping each preference vector $\lambda$ to a high-dimensional feature space using a polynomial embedding.
+Specifically, we define a degree-$p$ polynomial feature map $\phi_{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$ that enumerates all monomials up to degree $p$:
+$\phi_{p}(\lambda)=\left[1,\lambda_{1},\dots,\lambda_{d},\lambda_{1}^{2},\lambda_{1}\lambda_{2},\dots,\lambda_{d}^{p}\right]^{\top},$ (31)
+where the total number of features is given by:
+$N=\binom{d+p}{p}.$ (32)
+This expansion enables expressive yet tractable modeling of non-linear relationships between $\lambda$ vectors and alignment performance.
+D.2 Polynomial Regression
+Using the training set $\{(\lambda^{(i)},y^{(i)})\}_{i=1}^{n}$, we fit a least-squares regression model in the lifted feature space to learn the performance landscape:
+$\min_{\mathbf{w}\in\mathbb{R}^{N}}\sum_{i=1}^{n}\left(\mathbf{w}^{\top}\phi_{p}(\lambda^{(i)})-y^{(i)}\right)^{2}.$ (33)
+The resulting predictive function
+$f(\lambda)=\mathbf{w}^{\top}\phi_{p}(\lambda)$ (34)
+approximates the expected performance of any candidate $\lambda$ on downstream alignment tasks.
+D.3 Grid Construction over the Simplex
+To enable tractable sampling from the $\lambda$ space, we construct a discrete candidate set $\{\lambda^{(j)}\}_{j=1}^{k}\subset\Delta^{d}$ via one of the following methods:
+  • Simplex discretization: generate all combinations with replacement and normalize;
+  • Dirichlet sampling: draw $\lambda^{(j)}\sim\mathrm{Dir}(\boldsymbol{\alpha})$, where $\boldsymbol{\alpha}=\mathbf{1}$ yields a uniform prior over the simplex.
+D.4 Softmax-Based Scheduling Distribution
+We convert the predicted scores $s_{j}=f(\lambda^{(j)})$ into a probability distribution over the candidate $\lambda$ vectors using a softmax transformation:
+$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{k}\exp(\tau\cdot f(\lambda^{(l)}))},$ (35)
+where the temperature parameter $\tau>0$ controls the exploration–exploitation trade-off.
+D.5 Empirical Instantiation with Five Training Points
+To validate the polynomial scheduling framework, we use five observed training points—four unimodal $\lambda$ vectors and one uniform configuration—each paired with an empirical alignment accuracy:
+$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
+$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
+$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
+$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
+$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
+We set $p=2$, yielding $N=\binom{4+2}{2}=15$ features. Using gradient descent to minimize squared error, we obtain:
+$f(\lambda)$ $=0.35+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
+$+0.02\lambda_{1}^{2}+0.04\lambda_{1}\lambda_{2}+0.03\lambda_{1}\lambda_{3}+0.02\lambda_{1}\lambda_{4}$
+$+0.02\lambda_{2}^{2}+0.02\lambda_{2}\lambda_{3}+0.02\lambda_{2}\lambda_{4}$
+$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
+D.6 Lambda Sampling from Learned Scheduler
+We sample $k=10$ $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ) and compute predicted scores $f(\lambda^{(j)})$, converting them via softmax ($\tau=100$):
+$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{10}\exp(\tau\cdot f(\lambda^{(l)}))}$
+$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $f(\lambda)$ $p(\lambda)$
+0.212 0.334 0.245 0.209 0.452 0.095
+0.126 0.237 0.131 0.507 0.450 0.075
+0.542 0.079 0.256 0.123 0.455 0.117
+0.233 0.721 0.020 0.025 0.459 0.185
+0.004 0.334 0.281 0.381 0.449 0.067
+0.320 0.069 0.513 0.099 0.451 0.082
+0.236 0.155 0.358 0.251 0.452 0.093
+0.141 0.110 0.701 0.048 0.454 0.110
+0.070 0.198 0.388 0.344 0.451 0.081
+0.090 0.139 0.306 0.465 0.450 0.095
+Table 2: Sampled $\lambda$ vectors, their predicted scores $f(\lambda)$, and softmax sampling probabilities $p(\lambda)$ at temperature $\tau=100$.
+D.7 Scheduler vs. Uniform Comparison
+We compare our learned scheduler with a uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ baseline using LLaMA 3.2B Instruct.
+The learned scheduler achieves slightly better average accuracy (0.462 vs.
+0.461), and offers regularization via stochastic $\lambda$ sampling, enhancing robustness.
+Figure 12: Comparison of learned scheduler vs. fixed uniform $\lambda$ across six alignment benchmarks. The learned scheduler yields improved accuracy and diversity.

### v6

@@
-Multi-Preference Lambda-weighted Listwise DPO for Small-Scale Model Alignment
+Listwise Direct Preference Optimization with Multi-Dimensional Preference Mixing
 Abstract
-Large language models (LLMs) demonstrate strong generalization across a wide range of language tasks, but often generate outputs that misalign with human preferences.
-Reinforcement Learning from Human Feedback (RLHF) addresses this by optimizing models toward human preferences using a learned reward function and reinforcement learning, yielding improved alignment but suffering from high computational cost and instability.
-Direct Preference Optimization (DPO) simplifies the process by treating alignment as a classification task over binary preference pairs, reducing training overhead while achieving competitive performance.
-However, it assumes fixed, single-dimensional preferences and only supports pairwise supervision.
-To address these limitations, we propose Multi-Preference Lambda-weighted Listwise DPO, which allows the model to learn from more detailed human feedback and flexibly balance multiple goals such as helpfulness, honesty, and fluency.
-Our method models full-ranked preference distributions rather than binary comparisons, enabling more informative learning signals.
-The $\lambda$ vector controls the relative importance of different alignment goals, allowing the model to generalize across diverse human objectives.
-During inference, $\lambda$ can be adjusted without retraining, providing controllable alignment behavior for downstream use.
-We also introduce a learned scheduler that dynamically samples performant $\lambda$ configurations to improve robustness.
-Notably, our method requires only 20GB of GPU memory for training, making it suitable for compute-constrained settings such as academic labs, educational tools, or on-device assistants.
-Experiments on 1B–2B scale models show that our method consistently outperforms standard DPO on alignment benchmarks while enabling efficient, controllable, and fine-grained adaptation suitable for real-world deployment.
-Introduction
-Large language models (LLMs) such as GPT (Brown et al.
-2020) and LLaMA (Touvron and et al.
-2023) have emerged as powerful backbones in natural language processing, demonstrating impressive capabilities across tasks such as question answering, summarization, and instruction following.
-However, despite their success, LLMs often produce outputs that do not align well with human expectations and values.
-This misalignment frequently manifests as responses that are verbose, inaccurate, or unhelpful.
-It stems from their pretraining objective: next-token prediction, which emphasizes statistical co-occurrence over communicative intent, and inherently lacks exposure to explicit human feedback (Ouyang et al.
-2022).
-To address this, a line of alignment techniques has been developed to incorporate preference supervision.
-Among them, Reinforcement Learning from Human Feedback (RLHF) has become a standard paradigm (Christiano et al.
-2017; Ouyang et al.
-2022; Bai et al.
-2022).
-By training a reward model from human comparisons and optimizing it via reinforcement learning (e.g., PPO (Schulman et al.
-2017)), RLHF enables models to better match human values.
-However, the reliance on expensive reward modeling, unstable optimization dynamics, and high computational cost make RLHF difficult to scale and deploy reliably (Ziegler et al.
-2019).
-As a lightweight alternative, Direct Preference Optimization (DPO) has recently been proposed (Rafailov et al.
-2023).
-DPO reframes alignment as a classification problem over pairwise preferences, directly optimizing model outputs to prefer preferred responses over less preferred ones.
-Without requiring an explicit reward model or reinforcement loop, DPO achieves strong alignment with reduced training overhead.
-Nevertheless, existing DPO methods typically assume a single, fixed alignment objective and binary supervision.
-This limits their applicability in practical scenarios, where alignment is inherently multi-dimensional—balancing objectives such as helpfulness, factuality, honesty, and fluency (Bai et al.
-2022; Wu et al.
-2021).
-To accommodate this diversity, several controllable generation strategies have been explored, such as prompt tuning (Liu and et al.
-2021), control tokens (Keskar et al.
-2019), and multi-reward reinforcement learning (Lee et al.
-2023).
-However, these approaches often require retraining under each new preference configuration, introduce architectural complexity, or generalize poorly in unseen alignment regimes.
-In this work, we propose Multi-Preference Lambda-weighted Listwise DPO, a unified framework that extends the standard DPO formulation along two critical dimensions: supervision granularity and preference flexibility.
-Concretely, we generalize pairwise supervision to listwise preference modeling, enabling the model to learn from multiple ranked completions per prompt.
-This richer supervision signal allows the model to reason over global ranking structures, rather than relying on isolated binary comparisons.
-As a result, it improves gradient quality and better captures nuanced distinctions between high-quality and suboptimal responses.
-Simultaneously, we introduce a lambda-weighted composition mechanism, in which a preference vector $\lambda\in\Delta^{m}$ governs the relative importance of each alignment axis.
-This design allows a single model to flexibly interpolate between multiple objectives without retraining.
-To further enhance adaptability, we develop a learned scheduling strategy that dynamically adjusts the sampling distribution over $\lambda$ vectors based on their empirical performance.
-Empirical results on two strong 1B–2B scale LLMs (LLaMA and Qwen) across six alignment benchmarks demonstrate that our approach consistently outperforms standard DPO.
-Furthermore, by adjusting $\lambda$ during inference, our method supports personalized alignment without modifying model weights, making it highly practical for real-world deployment in education, conversational agents, and lightweight on-device applications.
-Related Work
-RLHF and DPO.
-Reinforcement learning from human feedback (RLHF) (Ouyang et al.
-2022; Bai et al.
-2022) has become a widely used paradigm for aligning large language models (LLMs) with human preferences.
-It typically involves training a reward model using pairwise human comparisons and then optimizing the base language model using reinforcement learning algorithms such as Proximal Policy Optimization (PPO) (Schulman et al.
-2017).
-This two-stage approach has demonstrated strong alignment performance, notably in systems like InstructGPT (Ouyang et al.
-2022) and Claude (Bai et al.
-2022).
-However, RLHF suffers from several limitations: reward models can be noisy or biased, PPO introduces high variance and instability during training, and the overall pipeline is computationally intensive and hard to scale (Ziegler et al.
-2019; Christiano et al.
-2017).
-To reduce this complexity, direct preference optimization (DPO) (Rafailov et al.
-2023) reframes preference alignment as a supervised classification problem.
-Instead of learning a reward model, DPO directly optimizes the policy to prefer preferred responses using a KL-regularized binary cross-entropy loss.
-This improves both training stability and implementation simplicity.
-Variants such as LLaMA-DPO (Liu et al.
-2023) and Quark (Yao et al.
-2023) have extended this framework with architectural tuning and reward quantization, showing strong performance in single-objective settings.
-However, most existing work in this area assumes fixed binary preferences and lacks the flexibility to accommodate multiple alignment goals or dynamically varying human objectives.
-Listwise Preference Learning.
-While pairwise preference modeling is common, listwise preference learning has been shown to better capture global ranking information (Burges et al.
-2005; Xia et al.
-2008).
-Instead of treating comparisons independently, listwise approaches model the entire ranking over a set of candidates, providing richer supervision and reducing sample complexity.
-In the context of language model alignment, UltraFeedback (Du et al.
-2023) provides large-scale human feedback data with complete ranking annotations over multiple completions.
-This enables more expressive and fine-grained alignment signals beyond binary feedback.
-Recent works such as GLOW (Xu et al.
-2023) leverage such ranked data to train reward models using listwise objectives.
-These models are then used in RLHF-style pipelines to align policies.
-However, they still rely on intermediate reward learning and reinforcement optimization.
-In contrast, our method directly optimizes the policy using listwise cross-entropy loss, bypassing the need for explicit reward modeling.
-This not only simplifies the training pipeline but also provides more stable gradients and better captures the structure of human preferences.
-Multi-Objective and Controllable Alignment.
-Human alignment is inherently multi-faceted, involving trade-offs among objectives such as helpfulness, honesty, informativeness, and harmlessness.
-Prior work has explored multi-objective reinforcement learning frameworks (Korbak et al.
-2023; Lee et al.
-2023) that combine multiple reward signals through scalarization or decomposition.
-However, these methods typically require retraining or manual tuning of trade-off weights for each new configuration.
-Similarly, reward interpolation techniques (Wu et al.
-2021) attempt to blend preferences using convex combinations of pretrained reward models, but they still involve separate optimization for each alignment setting.
-To enable better generalization, our method integrates a weight vector $\lambda$ into the training process, allowing the model to internalize a continuous spectrum of preference trade-offs.
-This is conceptually related to universal value function approximators (Schaul et al.
-2015), which generalize value functions across goals or reward vectors.
-In the alignment context, we treat $\lambda$ as a parameter controlling the target alignment behavior, enabling downstream control without model retraining.
-Alternative strategies for controllable generation include prompt-based steering (Brown et al.
-2020; Wei et al.
-2022), control tokens (Keskar et al.
-2019), and constrained decoding techniques (Lu et al.
-2022).
-While effective in certain settings, these methods often struggle with consistency, lack semantic grounding in human-aligned dimensions, or require careful hand-engineering.
-Our approach differs in that it integrates multi-dimensional alignment directly into the learning objective, enabling fine-grained, interpretable control through $\lambda$ without modifying prompts, decoding strategies, or model architecture.
-Preliminaries
-We begin by reviewing the foundation of Direct Preference Optimization (DPO) and its connection to KL-regularized reinforcement learning from human feedback (RLHF).
-Our setting assumes the existence of a base language model $\pi_{\text{ref}}$, a supervised fine-tuned (SFT) model, and a collection of ranked or pairwise human preferences.
-We aim to train a target policy $\pi_{\theta}$ that better aligns with human preferences while remaining close to $\pi_{\text{ref}}$.
-Pairwise Preference Modeling.
-Let $x$ denote the prompt, and $y_{w}\succ y_{l}$ represent a human preference between a pair of completions: $y_{w}$ (the preferred answer) and $y_{l}$ (the less preferred one).
-These preferences are sampled from an underlying distribution $p^{*}(y_{w},y_{l}|x)$, which is typically assumed to be induced by a latent reward function $r(x,y)$.
-A commonly adopted model is the Bradley-Terry (BT) model (Bradley and Terry 1952):
-$p^{*}(y_{w}\succ y_{l}\mid x)=\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$ (1)
-Given a dataset of comparisons $\mathcal{D}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, we can learn the reward function $r(x,y)$ by minimizing the negative log-likelihood under the BT model:
-$\mathcal{L}_{\text{BT}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r(x,y_{w})-r(x,y_{l}))\right]$ (2)
-where $\sigma(z)=1/(1+\exp(-z))$ is the logistic sigmoid.
+Recent alignment methods based on Direct Preference Optimization (DPO) reformulate preference learning as supervised optimization over pairwise comparisons, offering improved efficiency and stability over reinforcement learning from human feedback (RLHF).
+However, existing DPO-style methods implicitly assume a single fixed preference objective, which limits their ability to model the structured and sometimes conflicting nature of real-world human judgments that span multiple preference dimensions.
+In this work, we propose Listwise Direct Preference Optimization ($\lambda$-DPO), a unified framework that simultaneously improves supervision granularity and preference flexibility.
+Instead of collapsing multi-dimensional preference signals into a single ranking, $\lambda$-DPO constructs a mixture of listwise preference distributions weighted by a preference vector $\lambda$ on the probability simplex, enabling a single model to internalize a continuous spectrum of preference trade-offs.
+To further improve robustness, we introduce a performance-driven stochastic $\lambda$ scheduler that adaptively samples preference weights based on empirical downstream performance, explicitly mitigating the risks of misspecification inherent to static weighting schemes.
+We evaluate our method across multiple model families and scales on six widely used benchmarks.
+Experimental results show the consistent improvement against baselines.
+Listwise Direct Preference Optimization with Multi-Dimensional Preference Mixing
+Yuhui Sun University of Alberta          Xiyao Wang University of Toronto          Zixi Li Zhejiang University
+YiTian Ding School of Computer Science McGill University          Tianyang Ling University of Alberta          Jialuo Chen University of Toronto
+Tianyi Yu Alibaba Group (Ant Group)          Zhenlong Yuan Chinese Academy of Sciences          Jinman Zhao University of Toronto jzhao@cs.toronto.edu
+1 Introduction
+Large language models (LLMs) have demonstrated strong generalization capabilities such as question instruction following Brown et al.
+(2020); Touvron et al.
+(2023); OpenAI (2024).
+Despite these advances, their behavior in real-world applications frequently remains misaligned with human expectations, manifesting in issues such as factual inaccuracies, unsafe responses, and failures to appropriately interpret user intent Bai et al.
+(2022); Ouyang et al.
+(2022); Lin et al.
+(2022).
+Unlike early formulations that treated alignment as the optimization of a single performance metric, practical alignment objectives are inherently multi-dimensional, requiring models to simultaneously balance factors such as helpfulness, factuality, safety, fluency, and task-specific constraints often with intrinsic tensions and trade-offs among them Christiano et al.
+(2017); Wu et al.
+(2021); Korbak et al.
+(2023).
+Such human preferences are difficult to compress into a single scalar reward or fixed objective function; instead, they are more naturally expressed through relative comparisons and rankings over multiple candidate responses Bradley and Terry (1952); Burges et al.
+(2005); Cui et al.
+(2023).
+Consequently, the central challenge of alignment lies not in learning a single static “optimal” answer, but in modeling and generalizing the structured manifestation of multi-dimensional human preferences over sets of candidate outputs.
+Multiple methods have gradually emerged as a dominant paradigm for aligning LLMs with human expectations.
+Among them, reinforcement learning from human feedback (RLHF) optimizes generation policies using preference signals, typically via policy gradient methods such as Proximal Policy Optimization (PPO) Schulman et al.
+(2017) has demonstrated substantial improvements Ouyang et al.
+(2022).
+However, RLHF often suffers from high engineering and computational overhead, as well as instability arising from reward modeling errors and reinforcement learning dynamics.
+Direct Preference Optimization (DPO) Rafailov et al.
+(2023) reformulates alignment as a supervised learning problem, directly optimizing the policy using pairwise preference comparisons.
+Building upon the DPO framework, a growing body of work has proposed various extensions and variants such as CPO Xu et al.
+(2024), SimPO Meng et al.
+(2024), RRHF Yuan et al.
+(2023), ORPO Hong et al.
+(2024) and IPO Azar et al.
+(2024).
+These work modify the optimization objective, regularization strategy, or preference formulation to improve robustness and efficiency under different training regimes.
+Existing preference-alignment methods typically rely on an implicit modeling assumption that human preferences induce a single, fixed ranking over candidate responses.
+Recently, several listwise extensions of DPO such as LiPO Liu et al.
+(2025), Soft Preference Optimization Sharifnassab et al.
+(2024) have been proposed to directly optimize ranking distributions over sets of candidate responses, primarily aiming to improve supervision density and training stability.
+However, these methods still fundamentally assume that, for a given prompt, candidate responses correspond to a semantically fixed ranking target.
+In practice, this assumption is typically implemented by collapsing multi-dimensional preference signals into a single ranking objective.
+Nevertheless, real-world human preference data do not always support this formulation.
+For example, in multi-dimensional preference datasets such as UltraFeedback, candidate responses to the same prompt are evaluated independently along multiple criteria which often yield conflicting judgments across dimensions Cui et al.
+(2023).
+Based on the above observations, we propose multi-preference weighted listwise DPO ($\lambda$-DPO), a unified framework that simultaneously improves supervision granularity and preference flexibility.
+Specifically, instead of simplifying preference signals into binary pairwise comparisons, our method leverages the multi-candidate ranking information available for each prompt and directly learns to match a listwise preference distribution with the model-induced distribution.
+This formulation provides richer and lower-variance training signals while capturing global ranking relationships over candidate sets.
+To account for the inherently multi-dimensional nature of real-world alignment objectives, we further introduce a preference weight vector $\lambda$ defined on the probability simplex, which combines listwise preference distributions corresponding to different preference dimensions via a weighted mixture, yielding a controllable target distribution $p_{\lambda}$.
+This design allows a single model to internalize a continuous spectrum of preference trade-offs during training and enables inference-time controllable alignment through adjusting $\lambda$ without retraining.
+In addition, we design a learned $\lambda$ scheduling strategy that dynamically adapts the sampling distribution of $\lambda$ based on empirical performance, improving training robustness and generalization while remaining compatible with resource-constrained settings.
+In summary, this paper makes the following contributions:
+  • We identify a fundamental limitation of existing DPO and its extensions, which rely on binary supervision and a single fixed alignment objective, making them insufficient to capture the structure of multi-dimensional human preferences.
+To address this, we propose a multi-preference weighted listwise DPO framework.
+  • We empirically validate the effectiveness of $\lambda$-DPO across multiple model scales and evaluation benchmarks, demonstrating consistent improvements over existing DPO settings.
+  • Through ablation and analysis experiments, we systematically study the effects of different preference weight configurations and the learned $\lambda$ scheduling strategy, shedding light on how preference weighting influences training stability and alignment performance.
+Figure 1: Overall framework.
+2 Related Work
+Alignment and RLHF.
+RLHF Ouyang et al.
+(2022) is a foundational paradigm for aligning LLMs.
+Its core approach learns a reward model from human preference comparisons and optimizes the generation policy using policy gradient methods such as REINFORCE Sutton et al.
+(1999), PPO Schulman et al.
+(2017) or REINFORCE++ Hu et al.
+(2025).
+While effective, RLHF typically incurs high computational cost and suffers from training instability.
+To address these limitations, recent work explores alternatives that bypass explicit value or reward modeling, such as GRPO Shao et al.
+(2024), or reformulate alignment as a supervised learning problem, such as DPO Rafailov et al.
+(2023).
+DPO Variation.
+DPO has inspired a growing family of alignment methods that refine its objective or optimization strategy.
+Representative variants include SimPO Meng et al.
+(2024) simplifies the DPO objective by removing explicit reference policies; ORPO Hong et al.
+(2024) unifies preference optimization with likelihood training; CTO Xu et al.
+(2024) and IPO Azar et al.
+(2024) reinterpret preference learning from contrastive or implicit policy optimization perspectives; StepDPO Lai et al.
+(2024); Xu et al.
+(2025) extends DPO to step-level supervision for multi-step reasoning; and RDPO Park et al.
+(2024) explicitly disentangles response length from quality to mitigate length bias in preference optimization.
+More recently, online DPO methods study in interactive or streaming settings, highlighting the crucial role of sampling strategies in stabilizing and improving online alignment performance Shi et al.
+(2025); Tu et al.
+(2025); Pang et al.
+(2024).
+Despite differences in formulation and optimization details, these methods largely share a common reliance on pairwise supervision and assume a single implicit alignment objective.
+Recent surveys Xiao et al.
+(2025) provide systematic overviews of DPO-style methods, highlighting both their efficiency advantages and shared modeling assumptions.
+Listwise Preference Optimization.
+Recent work has explored preference optimization from pairwise comparisons to listwise feedback, aiming to more fully exploit the relative information among multiple candidate responses for a given prompt.
+Representative approaches include LiPO Liu et al.
+(2025) and Soft Preference Optimization Sharifnassab et al.
+(2024) formulate alignment as directly matching model distributions to target ranking distributions over response lists.
+Ordinal Preference Optimization Zhao et al.
+(2025) optimizes differentiable ranking metrics, and listwise extensions of DPO that leverage soft rankings or best-of-$K$ feedback.
+Beyond language models, listwise preference optimization has also been applied in other domains, such as vision–language models Pesaran Zadeh et al.
+(2025) and diffusion-based generative models Bai et al.
+(2025).
+3 Preliminaries
+We briefly review preference modeling and DPO. These form the foundation of our approach.
+Preference Modeling.
+Let $x$ denote a prompt, and $y_{w}\succ y_{l}$ a human preference indicating that completion $y_{w}$ is preferred over $y_{l}$.
+We assume such comparisons are drawn from an underlying preference distribution $p^{*}(y_{w}\succ y_{l}\mid x)$.
+A common instantiation is the Bradley–Terry model Bradley and Terry (1952) that parameterizes preferences via a latent scoring function $r(x,y)$:
+$p^{*}(y_{w}\succ y_{l}\mid x)$
+$=$ $\frac{\exp(r(x,y_{w}))}{\exp(r(x,y_{w}))+\exp(r(x,y_{l}))}$
+$=$ $\sigma\!\left(r(x,y_{w})-r(x,y_{l})\right)$ (1)
 KL-Regularized RL Objective.
 Once a reward model is trained, the standard RLHF approach fine-tunes the policy $\pi_{\theta}$ to maximize expected rewards while constraining deviation from the reference policy $\pi_{\text{ref}}$ via KL regularization:
-$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}\left[r(x,y)-\beta\cdot\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\right]$ (3)
+$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(y|x)}$ $\big[r(x,y)$ (2)
+$-\beta\mathrm{KL}\left(\pi_{\theta}(y|x)\|\pi_{\text{ref}}(y|x)\right)\big]$
 where $\beta>0$ controls the strength of the KL constraint.
 This objective encourages the policy to produce high-reward outputs while staying close to $\pi_{\text{ref}}$.
-However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO (Schulman et al.
-2017) to optimize.
+However, in practice, it requires sampling, reward evaluation, and reinforcement learning algorithms such as PPO Schulman et al.
+(2017) to optimize.
 Reparameterization and DPO.
@@
 Using Equation 3, we reparameterize the reward difference into the log-ratio of policies:
-$\beta(r(x,y_{w})-r(x,y_{l}))=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ (5)
+$\beta(r(x,y_{w})-r(x,y_{l}))$
+$=\log\frac{\pi_{r}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{r}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ (4)
 Replacing $\pi_{r}$ with $\pi_{\theta}$ yields the DPO loss:
-$-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (6)
+$-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\bigg[\log\sigma$ $\bigg(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}$ (5)
+$-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\bigg)\bigg]$
 This loss directly optimizes the policy parameters $\theta$ using preference comparisons, without sampling or reinforcement learning.
 It forms the foundation upon which we build our multi-preference and listwise extensions in later sections.
-Figure 1: Overview of the $\lambda$-DPO architecture.
-The system integrates listwise preference modeling and a controllable $\lambda$ vector to align model outputs with multi-dimensional human preferences.
-Multi-Preference Lambda-weighted Listwise DPO
+4 Multi-Preference Lambda-weighted Listwise DPO
 Motivated by the limitations of pairwise DPO under a single fixed preference distribution, we propose two orthogonal enhancements: listwise supervision and simplex-weighted optimization.
@@
 Together, these extensions allow controllable and scalable alignment with more nuanced human feedback.
-Listwise Preference Modeling.
+The overview pipeline is illustrated in Figure 1.
+4.1 Listwise Preference Modeling
 Let $\{y_{1},y_{2},\dots,y_{N}\}$ be a set of candidate completions for a prompt $x$, and let $p^{*}(y_{i}|x)$ be a listwise human preference distribution over these candidates.
@@
 We then define the listwise DPO objective as a cross-entropy loss between human preferences $p^{*}(y_{i}|x)$ and model predictions:
-$\mathcal{L}_{\text{listwise}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (8)
-Lambda-weighted Preference Composition.
+$\mathcal{L}_{\text{listwise}}(\theta)=$ (7)
+$-\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[\sum_{i=1}^{N}p^{*}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$
+4.2 Lambda-weighted Preference Composition
 To enable multi-objective control, we assume access to $m$ distinct preference dimensions, each defined by a distribution $p^{*(k)}(y_{i}|x)$ for objective $k\in\{1,\dots,m\}$.
@@
 The full Lambda-weighted Listwise DPO objective becomes:
-$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$ (10)
+$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=$ (9)
+$-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}|x)\cdot\log P_{\theta}(y_{i}|x)\right]$
 This expectation can be approximated by sampling $\lambda\sim\text{Unif}(\Delta^{m})$ or conditioning on user-specified preferences.
-What does $\lambda$-DPO update do?
+4.3 What does $\lambda$-DPO update do?
 To understand the optimization dynamics, consider the gradient of the loss:
-$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\right]$ (11)
+$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=-\mathbb{E}_{x,\lambda}\bigg[$ $\sum_{i}\left(p^{\lambda}(y_{i}|x)-P_{\theta}(y_{i}|x)\right)$
+$\ \ \cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\bigg]$ (10)
 This gradient updates the model to increase likelihood of responses aligned with weighted human preferences, while decreasing likelihood of responses overrepresented in the reference policy.
 Importantly, when $\lambda$ is fixed (e.g., $\lambda=(1,0,0)$), the objective reduces to a single-objective listwise DPO loss.
-Lambda-DPO Pipeline.
+Due to space limitations, full derivations of the $\lambda$-DPO objective and its gradient are deferred to Appendix A.
+4.4 The Pipeline
 The training process proceeds as follows:
@@
     Optimize $\pi_{\theta}$ using the listwise cross-entropy loss between $p^{\lambda}$ and $P_{\theta}$.
-In practice, we can reuse datasets like UltraFeedback (Du et al.
-2023), where multiple preference criteria are annotated.
-During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
-Theoretical Analysis of Multi-Preference Lambda-weighted Listwise DPO
-While Direct Preference Optimization (DPO) offers a principled and scalable approach to aligning language models with human feedback, it is fundamentally limited by its reliance on pairwise supervision.
-In this section, we provide a detailed theoretical analysis of these limitations and motivate our proposed extensions—namely, listwise supervision and preference mixing via simplex weighting—to overcome them in multi-objective alignment settings.
-Pairwise DPO and Its Limitations.
-Given a prompt $x$ and a pair of sampled completions $y_{w}$ (preferred) and $y_{l}$ (dispreferred), DPO defines a binary target distribution over the pair as:
-$p^{*}(y_{i}\mid x)=\begin{cases}1&\text{if }y_{i}=y_{w}\\ 0&\text{if }y_{i}=y_{l}\end{cases}$ (12)
-This construction is simple and efficient, allowing the loss to be computed from binary human preferences.
-However, it treats the preference signal as a hard decision, discarding nuanced feedback such as “somewhat better” or “equally good.” Furthermore, the normalization over only two samples induces high-variance gradients, as the updates depend on stochastic comparisons between limited completions.
-Since only the relative preference between $y_{w}$ and $y_{l}$ is modeled, all other completions (which might be equally or more suitable) are ignored.
-This can lead to suboptimal generalization, especially when the number of plausible high-quality completions is large.
-Listwise Distribution and Multi-Preference Supervision.
-To address the expressiveness gap, we generalize DPO to operate over a list of completions $\{y_{1},\dots,y_{N}\}$ using listwise supervision.
-Specifically, given a reward model $r(x,y_{i})$ that quantifies the quality of each completion, we define a normalized target distribution over the list using softmax:
-$p^{*}(y_{i}\mid x)=\frac{\exp(r(x,y_{i}))}{\sum_{j=1}^{N}\exp(r(x,y_{j}))}$ (13)
-This formulation assigns probability mass proportionally to reward values, providing smoother gradients and improving credit assignment.
-Unlike the binary setting, listwise DPO makes use of the full ranking structure, offering a richer training signal and faster convergence.
-Moreover, human preferences in practical applications are inherently multi-dimensional—comprising objectives like helpfulness, safety, factuality, politeness, and conciseness.
-Suppose we define $m$ distinct reward models, each representing one objective.
-We can construct a family of listwise distributions $\{p^{*(k)}(y_{i}|x)\}_{k=1}^{m}$, each aligned with a single criterion.
-A convex combination of these objectives is parameterized by a preference vector $\lambda\in\Delta^{m}$, forming a new composite distribution:
-$p^{\lambda}(y_{i}\mid x)=\sum_{k=1}^{m}\lambda_{k}\cdot p^{*(k)}(y_{i}\mid x)$ (14)
-This formulation enables controllable alignment by adjusting $\lambda$, allowing practitioners to emphasize certain values over others (e.g., prioritize safety over informativeness).
-The model is trained to match this target distribution via cross-entropy:
-$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (15)
-Its gradient reveals the direction of improvement, given by $\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}=$
-$-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (16)
-which encourages the model’s policy distribution to match the aggregated preference signal from multiple human value dimensions.
-Summary.
-In summary, Multi-Preference Lambda-weighted Listwise DPO offers a theoretically grounded generalization of the original DPO framework.
-First, it extends binary supervision to listwise distributions, providing smoother and more informative gradient updates.
-Second, it integrates diverse human objectives through a simplex-weighted mixture of reward-aligned targets, enabling dynamic and tunable preference alignment.
-Finally, it reduces sample inefficiency and training instability by leveraging full output rankings and avoiding overfitting to binary labels, thus making it well-suited for real-world alignment tasks with conflicting value axes.
-Comparative Experiment
-We compare standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups on two models: LLaMA 3.2-1B Instruct and Qwen 1.8B Chat.
-Both models are first supervised fine-tuned (SFT) on UltraChat-200k, then aligned using the UltraFeedback listwise dataset.
-Training Details.
-We use LoRA (rank 8, all-layer target) with a learning rate of $5\times 10^{-6}$, batch size 1, gradient accumulation steps 8, cosine learning rate schedule, and 5 epochs.
-The KL penalty $\beta$ is set to 0.1.
-$\lambda$-DPO uses the lambda_dpo loss with chunk size 1.
-Main Results
-Table 1 reports the accuracy of DPO and $\lambda$-DPO on six representative benchmarks for both LLaMA 3.2-1B Instruct and Qwen 1.8B Chat.
-On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +3.3% on GSM8K, and +3.6% on TruthfulQA.
-The improvement on WSC273 is also evident (+1.8%).
-For Qwen 1.8B Chat, $\lambda$-DPO outperforms DPO on all benchmarks except TruthfulQA, where DPO achieves a slightly higher score.
-Notably, large gains are observed on GSM8K (+14.7%) and MMLU (+3.9%), demonstrating the generalization benefits of $\lambda$-DPO across different model families.
-Model Method MMLU ARC HellaSwag TruthfulQA WSC273 GSM8K
-LLaMA 3.2-1B Instruct ORPO 0.456 0.355 0.447 0.435 0.678 0.327
-SIMPO 0.459 0.362 0.455 0.381 0.703 0.318
-DPO 0.458 0.364 0.456 0.378 0.696 0.312
-$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714 0.345
-Qwen 1.8B Chat ORPO 0.377 0.278 0.417 0.347 0.661 0.007
-SIMPO 0.415 0.295 0.423 0.355 0.658 0.007
-DPO 0.379 0.271 0.418 0.366 0.656 0.008
-$\lambda$-DPO 0.418 0.284 0.428 0.312 0.667 0.155
-Table 1: Accuracy comparison across six benchmarks.
-The first block compares ORPO, SIMPO, DPO, and $\lambda$-DPO on LLaMA 3.2-1B Instruct.
-The second block compares the same four methods on Qwen 1.8B Chat.
+In practice, we can reuse datasets where multiple preference criteria are annotated. During inference, changing $\lambda$ enables personalized alignment without modifying model parameters.
+Model Method MMLU ARC HellaSwag TruthfulQA WSC273 MathQA Average
+Llama 3.2-1B Instruct SIMPO 47.52 39.75 46.06 26.07 71.79 30.72 43.65
+DPO 47.52 40.37 45.87 26.07 71.43 30.89 43.69
+$\lambda$-DPO 48.71 45.61 46.15 29.01 75.09 32.83 46.23
+Qwen 1.8B Chat SIMPO 44.29 32.27 43.11 15.91 60.81 24.56 36.83
+DPO 44.30 32.72 43.17 15.91 61.54 23.38 36.84
+$\lambda$-DPO 45.01 38.51 43.82 18.12 64.84 29.41 39.95
+Llama 3-8B Instruct SIMPO 63.21 59.97 77.45 45.17 77.66 31.93 59.23
+DPO 61.49 63.16 76.33 45.53 72.53 31.59 58.44
+$\lambda$-DPO 61.03 73.25 76.44 46.63 85.35 36.87 63.26
+Qwen2 7B Instruct SIMPO 68.58 68.65 62.63 45.17 76.92 39.80 60.29
+DPO 68.11 69.50 62.50 45.41 75.09 42.53 60.52
+$\lambda$-DPO 68.17 69.22 63.60 46.27 78.75 42.91 61.49
+Table 1: Accuracy (%) comparison across six widely used benchmarks for different preference optimization methods and model backbones.
+Each block corresponds to a specific model, and rows compare SIMPO, DPO, and $\lambda$-DPO.
+Columns report performance on MMLU, ARC, HellaSwag, TruthfulQA, WSC273, and MathQA, along with the averaged score across all benchmarks.
 Bold values indicate the best performance within each model group.
-Visualization
-To further illustrate these results, we provide bar charts comparing DPO and $\lambda$-DPO performance on the two models.
-Figure 2: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
-Figure 3: Accuracy comparison between DPO and $\lambda$-DPO on Qwen 1.8B Chat.
-Performance Analysis and Discussion
-As shown in both Table 1 and Figures 23, $\lambda$-DPO consistently improves performance across most tasks.
-The method yields notable improvements on ARC, GSM8K, and MMLU, while maintaining parity on Winogrande.
-TruthfulQA is the only benchmark where DPO slightly outperforms $\lambda$-DPO (in the Qwen model).
-These results suggest that $\lambda$-DPO effectively captures human preference signals across diverse task types.
-Its listwise preference modeling contributes to better alignment with human-evaluated quality metrics such as factuality, helpfulness, and fluency.
-Nonetheless, for certain adversarial or logic-heavy tasks like TruthfulQA, the benefits of listwise supervision may be less reliable.
-Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting (Wei et al.
-2022)—could potentially mitigate this gap.
-Conclusion of Comparative Analysis
-In summary, $\lambda$-DPO brings two key improvements over standard DPO.
-First, by adopting listwise preference supervision, it leverages full ranking information instead of pairwise comparisons, resulting in richer and lower-variance gradient signals.
-Second, through the introduction of the $\lambda$ vector, it supports flexible interpolation across multiple alignment objectives (e.g., helpfulness, harmlessness, factuality), enabling controllable alignment behavior and improved generalization under diverse user preferences.
-These two enhancements jointly contribute to the observed performance gains across benchmarks.
-Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning techniques to further boost alignment quality on complex tasks.
-Ablation Experiment
-To further examine the effectiveness of different preference vector settings, we conduct an ablation experiment focusing on four unimodal $\lambda$-DPO models.
-Each model is trained using a fixed one-hot $\lambda$ vector, targeting only one of the four annotated dimensions: helpfulness, honesty, instruction-following, and truthfulness.
-This setup allows us to isolate the influence of each individual preference axis and compare their respective impacts across evaluation benchmarks.
-Unimodal Comparison Across Benchmarks
-We evaluate the four unimodal $\lambda$-DPO variants across six standard benchmarks—mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
-Figure 4 shows the accuracy distribution of each model across these tasks.
-Figure 4: Accuracy Comparison of Unimodal $\lambda$-DPO Models with Fixed Preference Vectors.
-All four models perform similarly across benchmarks.
-The honesty-focused model ($\lambda=[0,1,0,0]$) achieves the highest average accuracy, while the helpfulness-focused model ($\lambda=[1,0,0,0]$) excels on truthfulqa and performs well on gsm8k.
-Other preference models show only slight drops, with no significant performance degradation.
-Preference Ranking and Vector Hypothesis
-To better assess how each individual preference dimension contributes to generalization performance, we compute the average accuracy of the four unimodal $\lambda$-DPO models across six benchmark tasks:
-  • $\lambda=[1,0,0,0]$: 0.4593
-  • $\lambda=[0,1,0,0]$: 0.4598
-  • $\lambda=[0,0,1,0]$: 0.4543
-  • $\lambda=[0,0,0,1]$: 0.4548
-  • $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$: 0.4612
-The results indicate that all four unimodal preference dimensions contribute to model alignment, with helpfulness and honesty yielding slightly better generalization.
-However, the uniform interpolation $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ achieves the highest overall accuracy (0.4612), outperforming all unimodal variants.
-This highlights the benefit of explicitly retaining the $\lambda$ vector for multi-objective supervision.
-Unlike standard DPO, which implicitly collapses preferences into a fixed form, $\lambda$-DPO provides fine-grained control over alignment trade-offs and leads to more robust generalization.
-Discussion and Ablation Analysis
-To further understand how individual preference dimensions influence alignment, we observe from Figure 4 that:
-  • The $\lambda=[1,0,0,0]$ (helpfulness) model performs best on truthfulqa and achieves solid results on gsm8k.
-  • The $\lambda=[0,1,0,0]$ (honesty) model leads slightly in overall average accuracy, especially strong on mmlu.
-  • The $\lambda=[0,0,1,0]$ (instruction-following) and $\lambda=[0,0,0,1]$ (fluency) models show close and competitive performance, contrary to the notion that they are stylistically weaker.
-These results indicate that optimizing for either content-driven or stylistic objectives in isolation yields comparable outcomes under our listwise framework.
-It further motivates the integration of multiple preference signals, especially for tasks requiring both correctness and user-aligned expressiveness.
-Performance-Driven Lambda Scheduling via Polynomial Modeling
-Performance-Driven Scheduler Design
-We propose a performance-driven scheduler to improve multi-objective alignment by dynamically selecting preference vectors $\lambda$ based on their estimated effectiveness.
-Instead of fixing a single $\lambda$, the scheduler learns a distribution over $\lambda$ configurations that reflect observed performance.
-The scheduler fits a polynomial regression model to past $(\lambda,\text{accuracy})$ pairs and uses it to score new candidates sampled from a Dirichlet prior over the simplex.
-A softmax over predicted scores then defines sampling probabilities, balancing exploitation of high-performing regions and exploration.
-This adaptive strategy promotes training diversity and improves robustness.
-Empirical Comparison with Uniform Baseline
-As shown in Figure 12, the learned scheduler achieves slightly better average accuracy (0.462 vs.
-0.461) across six evaluation benchmarks when compared with a static baseline that uses the uniform configuration $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$.
-While the performance gain is modest, the scheduler enhances robustness and reduces overfitting by continuously adapting to observed training signals.
-Importantly, this framework supports flexible preference adaptation and can be extended to personalized or task-specific settings without requiring retraining.
-Figure 5: Comparison between learned $\lambda$ scheduler and uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ across six alignment benchmarks.
-The learned scheduler achieves a slightly higher average accuracy and improves robustness by distributing preference sampling over higher-quality regions of the simplex.
-Extension to LLaMA-8B
-To assess the scalability of $\lambda$-DPO, we extend our experiments to the LLaMA-8B model.
-Across six standard benchmarks, $\lambda$-DPO yields an average relative improvement of approximately 0.3% over DPO, compared to a much larger average gain of 1.7% on the LLaMA-1B model.
-Figure 6: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA-8B across six benchmarks.
-As shown in Figure 6, $\lambda$-DPO achieves better alignment on the 1B model, with notably larger gains on challenging tasks such as GSM8K.
-This suggests that smaller models benefit more from listwise multi-preference supervision, likely because alignment signals have a stronger relative effect when model capacity is limited.
-Moreover, training the 1B model only requires around 20GB of GPU memory, making it both efficient and accessible for academic labs, startups, or edge-focused applications.
-This low resource footprint enables rapid experimentation and deployment on commodity hardware, without sacrificing alignment quality.
-As such, $\lambda$-DPO is well-suited for scalable, real-world use cases where cost, speed, and controllability are critical.
-Conclusion
-In this paper, we propose Multi-Preference Lambda-weighted Listwise DPO ($\lambda$-DPO) to enable efficient, controllable, and fine-grained alignment of language models with diverse human preferences.
-For robust and flexible alignment, we generalize pairwise supervision to listwise preference modeling, allowing the model to learn from ranked completions and capture nuanced preference structures.
-To support personalized and multi-objective alignment, we introduce a simplex-weighted $\lambda$ vector that dynamically balances objectives such as helpfulness, honesty, and fluency—without retraining.
-To enhance generalization and robustness, we incorporate a learned scheduler that adaptively samples high-performing $\lambda$ configurations during training, improving diversity while mitigating modeling bias.
-Additionally, the method is designed to be resource-efficient, requiring only 20GB of GPU memory, making it practical for deployment in constrained scenarios such as on-device assistants and educational tools.
-Experimental results on LLaMA 3.2-1B and Qwen-1.8B demonstrate that $\lambda$-DPO achieves consistent improvements over standard DPO, with average accuracy gains of +1.7% and +2.8%, respectively, across standard alignment benchmarks.
-Acknowledgements
-We thank the contributors of the original DPO framework for their foundational work, which inspired this extension.
-We also appreciate the open-source efforts of the Hugging Face team and the LLaMA-Factory project for enabling scalable preference-based fine-tuning.
-This work was made possible through collaborative effort and self-supported compute resources by the authors: Yuhui Sun, Xiyao Wang, Zixi Li, Zhenlong Yuan and Jinman Zhao.
-We further thank colleagues and early readers for helpful feedback that improved the theoretical formulation and presentation.
-Appendix A A Mathematical Derivations
+5 Theoretical Analysis
+While DPO offers a principled and scalable approach to aligning language models with human feedback, its reliance on pairwise supervision imposes intrinsic limitations in expressive power, particularly in multi-objective alignment settings.
+In this section, we provide a theoretical analysis of this limitation and motivate our proposed extensions for better capturing and aggregating multi-dimensional human preferences.
+5.1 Pairwise Limitation
+Given a prompt $x$ and a sampled pair of completions $y_{w}$ (preferred) and $y_{l}$ (dispreferred), Direct Preference Optimization (DPO) relies on binary preference supervision.
+From a listwise perspective, such supervision can be interpreted as inducing an implicit target distribution over the sampled pair:
+$p^{*}(y_{i}\mid x)=\begin{cases}1&\text{if }y_{i}=y_{w},\\ 0&\text{if }y_{i}=y_{l}.\end{cases}$ (11)
+This formulation is simple and computationally efficient, allowing policy optimization directly from binary human feedback.
+However, treating preferences as hard assignments discards finer-grained signals, such as partial preference strength or equivalence between responses.
+Moreover, because gradient updates depend on a single sampled pair, the resulting optimization can exhibit high variance across iterations.
+Since the objective only constrains the relative preference between $y_{w}$ and $y_{l}$, other plausible or high-quality completions for the same prompt are ignored.
+Consequently, the expressive capacity of pairwise DPO is limited when multiple responses are comparably suitable, a limitation that becomes particularly pronounced in multi-objective alignment settings.
+5.2 Pairwise Supervision as a Degenerate Listwise Objective
+Equation (11) can be viewed as a degenerate form of listwise preference supervision, where the target distribution assigns all probability mass to a single completion within a sampled pair.
+More generally, let $\mathcal{Y}(x)=\{y_{1},\dots,y_{K}\}$ denote a set of candidate completions for prompt $x$.
+A listwise preference objective aims to model a target distribution $p^{*}(\cdot\mid x)$ over $\mathcal{Y}(x)$, satisfying
+$\sum_{i=1}^{K}p^{*}(y_{i}\mid x)=1,\qquad p^{*}(y_{i}\mid x)\geq 0.$ (12)
+Under pairwise DPO, the effective target distribution collapses to a distribution supported on only two sampled completions, with all mass assigned to the preferred one.
+This corresponds to an extreme point of the probability simplex in Eq.
+(12), with zero entropy.
+As a result, the optimization objective is insensitive to the relative quality of other candidates in $\mathcal{Y}(x)$, even when multiple responses are comparably suitable under certain preference trade-offs.
+5.3 Incompatibility with Multi-Objective Preferences
+Consider a setting with multiple preference dimensions, such as helpfulness, factuality, and safety.
+Let $r^{(k)}(x,y)$ denote a latent scoring function associated with the $k$-th preference criterion.
+Each criterion induces its own ordering over candidate responses in $\mathcal{Y}(x)$.
+In general, there does not exist a single scalar function $r(x,y)$ such that
+$r(x,y_{i})>r(x,y_{j})$ (13)
+$\quad\Longleftrightarrow\quad r^{(k)}(x,y_{i})>r^{(k)}(x,y_{j})\;\;\forall k.$
+As a consequence, different preference dimensions may induce conflicting pairwise comparisons over the same response set.
+Reducing such multi-dimensional preference signals to a single binary comparison necessarily discards information about preference trade-offs.
+From this perspective, the limitation of pairwise DPO is structural: binary comparisons cannot represent preference distributions that depend on multiple, potentially competing objectives.
+$\lambda$ Setting MMLU ARC HellaSwag TruthfulQA WSC273 MathQA Average
+$[1,0,0,0]$ (Helpfulness) 46.36 48.91 46.22 28.03 70.70 33.53 45.63
+$[0,1,0,0]$ (Honesty) 46.48 49.04 46.25 27.91 70.33 33.63 45.61
+$[0,0,1,0]$ (Instruction-Following) 46.39 49.29 46.24 28.03 71.06 33.67 45.78
+$[0,0,0,1]$ (Fluency) 46.26 48.84 46.22 28.15 70.33 33.37 45.53
+$[0.25,0.25,0.25,0.25]$ (Uniform) 48.71 45.61 46.15 29.01 75.09 32.83 46.23
+Table 2: Ablation study of different $\lambda$-weight configurations in $\lambda$-DPO, conducted on the LLaMA 3.2-1B Instruct.
+6 Results
+6.1 Experimental Setup
+Model Selection.
+We evaluate our approach on representative open-weight LLMs from both the Llama3 Touvron et al.
+(2023) and Qwen Qwen (2025) families.
+Specifically, we consider Llama 3.2-1B Instruct, Llama 3-8B Instruct, Qwen 1.8B Chat and Qwen2 7B Instruct.
+Benchmark Selection.
+To evaluate the effectiveness of our approach across diverse reasoning and alignment-sensitive tasks, we conduct experiments on a collection of widely used benchmarks spanning knowledge-intensive reasoning, commonsense inference, factual accuracy, and mathematical problem solving.
+Specifically, we include MMLU Hendrycks et al.
+(2021), ARC Clark and others (2018), HellaSwag Zellers et al.
+(2019), TruthfulQA Lin et al.
+(2022), WSC273 Levesque and others (2012), and MathQA Amini et al.
+(2019).
+Detailed benchmarks statistics reported in Appendix B.
+Baseline Selection.
+We compare our method against representative preference-based alignment baselines that are closely related to our formulation.
+In particular, we include DPO Rafailov et al.
+(2023) and SimPO Meng et al.
+(2024).
+Training Detail.
+We set the preference scaling parameter $\beta$ to 0.1, following common practice in prior work.
+Models are trained for three epochs with a learning rate of $5\times 10^{-6}$, using a cosine learning rate scheduler with a warmup ratio of 0.1.
+We use an batch size of 32, and the maximum sequence length is set to 2048 tokens.
+Our experiments are conducted on 4xH200 GPUs.
+Dataset.
+We conduct fine-tuning using the Ultrafeedback dataset Cui et al.
+(2023), a large-scale, instruction-following corpus designed for alignment and preference optimization.
+Ultrafeedback contains diverse prompts and multiple candidate responses annotated with multi-preference signals.
+6.2 Main Results
+We report the main experimental results in Table 1 across six widely used benchmarks.
+All results are reported in terms of accuracy (%), with higher values indicating better performance.
+Overall, $\lambda$-DPO consistently achieves the strongest average performance across all evaluated models, outperforming both SIMPO and DPO in every setting.
+The gains are particularly pronounced on benchmarks that are sensitive to alignment-related properties, such as TruthfulQA and WSC273, where $\lambda$-DPO yields clear and consistent improvements.
+This suggests that moving beyond pairwise supervision and explicitly aggregating preference signals leads to more effective alignment with nuanced human judgments.
+We further observe that $\lambda$-DPO delivers robust improvements across model scales.
+For smaller models (e.g., Llama 3.2-1B Instruct and Qwen 1.8B Chat), $\lambda$-DPO provides substantial relative gains, indicating improved data efficiency under limited capacity.
+For larger models (e.g., Llama 3-8B Instruct and Qwen2-7B Instruct), $\lambda$-DPO continues to outperform pairwise baselines, demonstrating that the proposed approach scales effectively with model size.
+Taken together, these results validate the effectiveness and generality of $\lambda$-DPO across diverse benchmarks and model families.
+A closer inspection of Table 1 shows that the performance gains of $\lambda$-DPO are most pronounced on benchmarks that are sensitive to alignment-related properties, particularly TruthfulQA and WSC273.
+On WSC273, $\lambda$-DPO outperforms DPO by +3.66, +3.30, +12.82, and +3.66 points for Llama 3.2-1B, Qwen 1.8B, Llama 3-8B, and Qwen2-7B, respectively, and also surpasses SIMPO by +3.30, +4.03, +7.69, and +1.83 points on the same models.
+Similarly, on TruthfulQA, $\lambda$-DPO improves over DPO by +2.94 (Llama 3.2-1B), +2.21 (Qwen 1.8B), +1.10 (Llama 3-8B), and +0.86 (Qwen2-7B), while exceeding SIMPO by +2.94, +2.21, +1.46, and +1.10 points, indicating more truthful and human-consistent responses across model scales.
+The especially large gain on WSC273 for Llama 3-8B highlights that aggregating preference signals beyond pairwise supervision can substantially improve alignment-critical commonsense judgment.
+In contrast, performance differences on other benchmarks, including ARC, MathQA, and the broader commonsense task HellaSwag, are generally more moderate, suggesting that the primary advantage of $\lambda$-DPO lies in better capturing nuanced human judgments without sacrificing generalization.
+$\lambda$ Setting MMLU ARC HellaSwag TruthfulQA WSC273 MathQA Average
+Uniform 48.71 45.61 46.15 29.01 75.09 32.83 46.23
+Learned Scheduler 48.35 48.76 46.20 29.03 75.06 33.53 46.82
+Table 3: Uniform vs. Learned $\lambda$, conducted on the LLaMA 3.2-1B Instruct.
+6.3 Additional Analysis
+Uniform vs. Unimodal Preference Weights.
+We investigate the effect of different $\lambda$ configurations by comparing a uniform mixture of preference dimensions with unimodal (one-hot) alternatives, and the model is trained to optimize a single objective at a time.
+Table 2 reports results on LLaMA 3.2-1B Instruct.
+Across unimodal settings, performance is relatively similar, and no single preference dimension consistently dominates across all benchmarks.
+This suggests that the gains of $\lambda$-DPO are not driven by optimizing any individual objective in isolation.
+In contrast, the uniform mixture $\lambda=(0.25,0.25,0.25,0.25)$ achieves the best overall average performance, yielding improvements on alignment-sensitive tasks such as TruthfulQA and WSC273 while maintaining competitive accuracy on other benchmarks.
+Although certain unimodal configurations achieve slightly higher scores on individual tasks (e.g., instruction-following on ARC), they fail to generalize consistently across the benchmark suite.
+These results support our central hypothesis that explicitly aggregating multi-dimensional preference signals leads to more robust and balanced alignment than optimizing a single preference dimension alone.
+Uniform vs. Learned $\lambda$.
+Static uniform $\lambda$ configurations face two key limitations.
+First, different benchmarks exhibit different sensitivities to preference dimensions, whereas a uniform $\lambda$ implicitly assumes that all objectives are equally important.
+This mismatch prevents the model from adapting to benchmark-specific preference structure and can lead to suboptimal performance.
+Second, although we fit a polynomial performance model $f(\lambda)$ to approximate downstream accuracy over the simplex, this approximation is inherently imperfect: $\lambda$ values predicted as weak by $f$ may still yield strong true performance.
+Thus, selecting only the maximizer of $f$ risks collapsing to a poorly explored region of the preference space.
+Instead of committing to a single $\lambda$, we construct a stochastic scheduler that samples preference weights in a performance-aware manner.
+Specifically, we define an idealized continuous sampling distribution over the simplex
+$p(\lambda)\;\propto\;\exp(\tau f(\lambda)),\qquad\lambda\in\Delta_{4},$
+where $\tau>0$ controls the exploration–exploitation balance.
+Since this density cannot be computed in closed form, we discretize it by sampling candidate $\{\lambda^{(j)}\}_{j=1}^{k}$ from a Dirichlet prior and applying a softmax:
+$p(\lambda^{(j)})=\frac{\exp(\tau f(\lambda^{(j)}))}{\sum_{l}\exp(\tau f(\lambda^{(l)}))}.$
+This produces a learned $\lambda$ scheduler that prioritizes promising trade-offs while still exploring alternative preference allocations, explicitly guarding against misspecification of $f$.
+The result is presented in Table 3, with the full formulation and implementation details of the learned $\lambda$ scheduler provided in Appendix C.
+Empirically, this design yields clear gains.
+Compared to the uniform mixture $\lambda=(0.25,0.25,0.25,0.25)$, which achieves an average score of $46.23$, the learned scheduler improves the overall average to $46.82$.
+It further delivers meaningful benchmark-specific improvements, such as ARC ($45.61\rightarrow 48.76$) and MathQA ($32.83\rightarrow 33.53$), while maintaining strong alignment-sensitive performance on TruthfulQA and WSC273.
+These results demonstrate that performance-driven stochastic $\lambda$ adaptation is more robust than any static weighting scheme.
+Benchmark-Specific Preference Sensitivity.
+Using the results reported in Table 2, we further analyze how different benchmarks respond to individual preference dimensions.
+Rather than focusing on overall performance aggregation, this analysis adopts a benchmark-centric perspective and examines which preference objectives are most aligned with specific evaluation tasks.
+We observe that no single preference dimension dominates across all benchmarks.
+For example, ARC exhibits its strongest performance under the instruction-following–focused configuration, suggesting that precise adherence to task instructions plays a critical role in this benchmark.
+In contrast, TruthfulQA benefits more consistently from balanced or honesty-oriented preference mixtures, reflecting its sensitivity to factual correctness and avoidance of hallucination.
+Similarly, WSC273 shows improved performance under configurations that emphasize balanced reasoning and fluency, rather than any single isolated objective.
+These observations highlight that different benchmarks implicitly encode distinct preference structures
+7 Conclusions
+In this work, we revisit preference optimization for language model alignment from a multi-dimensional perspective.
+We identify that existing DPO-based methods implicitly assume a single fixed preference objective, which limits their expressive power when human judgments are structured, heterogeneous, and sometimes conflicting across dimensions.
+To address this limitation, we propose $\lambda$-weighted Listwise Direct Preference Optimization ($\lambda$-DPO), a unified framework that combines listwise supervision with flexible preference aggregation for multi-dimensional alignment.
+Extensive experiments across multiple model families and six widely used benchmarks demonstrate that $\lambda$-DPO consistently outperforms strong DPO-based baselines.
+8 Limitations
+While our method improves the flexibility and expressiveness of preference optimization, this work still has several limitations that warrant further investigation.
+Our evaluation primarily relies on accuracy over standard benchmarks as a proxy for alignment quality.
+Although these benchmarks capture important aspects such as reasoning ability, factual correctness, and instruction-following, they do not fully reflect the complexity of real-world deployment scenarios, including long-horizon interactions, dynamically changing user preferences, and safety-critical behaviors.
+Appendix A Mathematical Derivations
 A.1 Deriving the Final Training Objective
@@
 Given a human-provided listwise preference distribution $p^{*}_{i}$, we minimize the cross-entropy between $p^{*}_{i}$ and $P_{\theta}$:
-$\mathcal{L}^{\text{DPO}}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}\right)\right]$ (20)
+$\mathcal{L}^{\text{DPO}}(\theta)=$ $\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\bigg[-\sum_{i=1}^{N}p^{*}_{i}$
+$\cdot\log\left(\frac{\exp(r_{\theta}(x,y_{i}))}{\sum_{j=1}^{N}\exp(r_{\theta}(x,y_{j}))}\right)\bigg]$ (17)
 Substituting the reward expression and simplifying:
-$\mathcal{L}^{\text{DPO}}(\theta)&=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}Z(x)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}Z(x)^{\beta}}\right)\right]\\ &=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]\end{aligned}$
+$\begin{aligned} &\mathcal{L}^{\text{DPO}}(\theta)\\ =&\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}Z(x)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}Z(x)^{\beta}}\right)\right]\\ =&\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p^{*}_{i}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]\end{aligned}$
 This becomes the canonical DPO loss.
 To extend this to multiple preference dimensions, we assume there are $m$ reward components indexed by $k$, each with its own preference distribution $p_{i}^{*(k)}$:
-$\mathcal{L}^{(k)}(\theta)=\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\right]$ (21)
+$\mathcal{L}^{(k)}(\theta)=$ $\mathbb{E}_{(x,\{y_{i}\})\sim\mathcal{D}}\bigg[-\sum_{i=1}^{N}p_{i}^{*(k)}$
+$\cdot\log\left(\frac{\left(\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\text{ref}}(y_{i}|x)}\right)^{\beta}}{\sum_{j=1}^{N}\left(\frac{\pi_{\theta}(y_{j}|x)}{\pi_{\text{ref}}(y_{j}|x)}\right)^{\beta}}\right)\bigg]$ (18)
 We then define a user-controllable or sampled weight vector $\lambda\in\Delta^{m}$ (the $m$-dimensional probability simplex: $\lambda_{k}\geq 0$, $\sum_{k}\lambda_{k}=1$), and define the aggregated loss:
@@
 To enable generalization over arbitrary preferences, we optimize the expected loss over $\lambda\sim\text{Unif}(\Delta^{m})$:
-$\mathcal{L}_{\text{final}}(\theta)=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\mathcal{L}^{\lambda}(\theta)\right]=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)\right]$ (23)
+$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\mathcal{L}^{\lambda}(\theta)\right]$ (20)
+$=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathcal{L}^{(k)}(\theta)\right]$ (21)
 A.2 From Final Objective to Lambda-DPO Form
@@
 So the aggregated expected loss becomes:
-$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\left[\sum_{k=1}^{m}\lambda_{k}\cdot\mathbb{E}_{(x,\{y_{i}\})}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]\right]$ (25)
+$\mathcal{L}_{\text{final}}(\theta)$ $=\mathbb{E}_{\lambda\sim\text{Unif}(\Delta^{m})}\Bigg\{\sum_{k=1}^{m}\lambda_{k}$
+$\cdot\mathbb{E}_{(x,\{y_{i}\})}\left[-\sum_{i=1}^{N}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]\Bigg\}$ (23)
 $=\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[-\sum_{i=1}^{N}\sum_{k=1}^{m}\lambda_{k}p_{i}^{*(k)}\log P_{\theta}(y_{i}|x)\right]$ (24)
@@
 So the objective becomes:
-$\mathcal{L}_{\lambda\text{-DPO}}(\theta)=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (28)
+$\mathcal{L}_{\lambda\text{-DPO}}(\theta)$
+$=-\mathbb{E}_{(x,\{y_{i}\}),\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\log P_{\theta}(y_{i}\mid x)\right]$ (26)
 Thus, by switching the summation order (i.e., aggregating over preference weights before computing loss), we arrive at an equivalent but more compact objective that is easier to optimize and sample in practice.
-A.3 Gradient of Lambda-DPO Loss
+A.3 A.3 Gradient of Lambda-DPO Loss
 We compute the gradient of $\mathcal{L}_{\lambda\text{-DPO}}$:
-$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}$ $=-\mathbb{E}_{x,\lambda}\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\nabla_{\theta}\log P_{\theta}(y_{i}\mid x)\right]$ (29)
-$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$
+$\nabla_{\theta}\mathcal{L}_{\lambda\text{-DPO}}$
+$=-\mathbb{E}_{x,\lambda}$ $\left[\sum_{i=1}^{N}p^{\lambda}(y_{i}\mid x)\cdot\nabla_{\theta}\log P_{\theta}(y_{i}\mid x)\right]$ (27)
+$=-\mathbb{E}_{x,\lambda}$ $\bigg[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)$
+$\qquad\quad\cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\bigg]$ (28)
 where we use:
@@
 This completes the derivation of the Lambda-weighted DPO objective and its gradient form.
-Appendix B B Experiment Setup
-B.1 Experimental Objectives and Comparison Targets
-The primary goal of our experiments is to evaluate the performance of the proposed $\lambda$-DPO algorithm compared to the standard DPO method.
-Specifically, we assess their effectiveness across different model architectures and training settings using public preference datasets and standard evaluation benchmarks.
-We consider two levels of comparison:
-Model comparison.
-We experiment with two strong open-source language models—LLaMA and Qwen.
-These models differ in architecture, pretraining corpora, and tokenization schemes, providing a robust testbed for evaluating generalizability.
-Method comparison.
-We compare DPO and $\lambda$-DPO.
-For each method-model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure comparisons are fair and reflect the best achievable performance under comparable resource constraints.
-B.2 Training Settings
-Inspired by the SimPO baseline, we adopt a unified training pipeline:
-  • Finetuned-Base: The model is first SFT-trained on UltraChat-200k, then optimized on UltraFeedback via DPO or $\lambda$-DPO to reflect realistic alignment scenarios.
-B.3 Evaluation Benchmarks and Metrics
-We evaluate model performance using the lm-eval-harness (Gao et al.
-2024) framework, which reports accuracy on diverse benchmarks covering reasoning, factual correctness, and commonsense understanding.
-Specifically, we report average accuracy across the following six established datasets:
-  • MMLU: 57 subjects, 15K questions assessing academic knowledge (Hendrycks et al. 2020).
-  • ARC: 1.1K grade-school science questions requiring reasoning (Clark et al. 2018).
-  • HellaSwag: 10K commonsense sentence completions with distractors (Zellers et al. 2019).
-  • TruthfulQA: 817 questions testing factuality and hallucination (Lin, Hilton, and Evans 2021).
-  • Winograd: 273 examples probing coreference and syntax (Levesque, Davis, and Morgenstern 2012).
-  • GSM8K: 8.5K arithmetic reasoning problems (Cobbe et al. 2021).
-B.4 Ablation: Impact of Lambda Weighting
-To better understand the impact of the $\lambda$ vector in $\lambda$-DPO, we conduct ablation experiments focusing on how different preference weightings affect alignment performance.
-Specifically, we compare two $\lambda$ schedule variants within the listwise $\lambda$-DPO framework:
-  • Uniform $\lambda$: Equal weights across all dimensions ($\lambda=(1/m,\dots,1/m)$), simulating balanced multi-objective aggregation.
-  • Unimodal $\lambda$: One-hot vectors (e.g., $(1,0,\dots)$) isolating single-dimension optimization without interpolation.
-This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting.
-Appendix C C Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
-We provide benchmark accuracy comparisons between DPO and $\lambda$-DPO across multiple language models on the UltraFeedback dataset.
-Performance is reported across six tasks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
-C.1 Qwen-1.8 (Base)
-Figure 7 shows that $\lambda$-DPO achieves modest improvements over standard DPO on most tasks using the Qwen-1.8 base model.
-Figure 7: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 (base).
-C.2 Qwen-1.8 Chat
-Figure 8 presents results on Qwen-1.8 Chat, where $\lambda$-DPO consistently improves accuracy, especially on reasoning-heavy benchmarks like arc and hellaswag.
-Figure 8: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 Chat.
-C.3 LLaMA 3.2 1B (Base)
-As shown in Figure 9, $\lambda$-DPO outperforms DPO slightly on LLaMA 3.2 1B base across four out of six tasks, demonstrating generalization beyond the Qwen series.
-Figure 9: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B base model.
-C.4 LLaMA 3.2 1B (Instruct)
-Figure 10 illustrates results for the Instruct-tuned variant of LLaMA 3.2 1B. The performance gap between $\lambda$-DPO and DPO is generally preserved, indicating robustness to instruction tuning.
-Figure 10: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
-C.5 Unimodal Ablation
-We visualize the performance of four unimodal $\lambda$-DPO models, each trained using a one-hot preference vector that activates only one of the four human preference dimensions: helpfulness, honesty, coherence, and fluency.
-Accuracy is reported across six standard benchmarks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
-Figure 11: Accuracy comparison of four unimodal $\lambda$-DPO models, each trained with a single active preference dimension.
-All models perform similarly across benchmarks, with slight advantages for helpfulness and honesty on average.
-This suggests that no single preference dominates, and multi-dimensional interpolation may improve overall alignment.
-Appendix D D Performance-Driven Lambda Scheduling via Polynomial Modeling
+Appendix B Benchmark Statistics
+We report average accuracy across the following six established datasets:
+  • MMLU: 57 subjects, 15K questions assessing academic knowledge Hendrycks et al. (2021).
+  • ARC: 1.1K grade-school science questions requiring reasoning Clark and others (2018).
+  • HellaSwag: 10K commonsense sentence completions with distractors Zellers et al. (2019).
+  • TruthfulQA: 817 questions testing factuality and hallucination Lin et al. (2022).
+  • Winograd: 273 examples probing coreference and syntax Levesque and others (2012).
+  • MathQA: 37k Math reasoning problems Amini et al. (2019).
+Appendix C Performance-Driven Lambda Scheduling via Polynomial Modeling
 We introduce a general and flexible method for performance-driven lambda scheduling, which constructs a probability distribution over preference vectors $\lambda\in\Delta^{d}$ based on observed alignment performance.
@@
 The scheduling process unfolds in four steps:
-D.1 Polynomial Feature Mapping
+C.1 Polynomial Feature Mapping
 We begin by mapping each preference vector $\lambda$ to a high-dimensional feature space using a polynomial embedding.
@@
 This expansion enables expressive yet tractable modeling of non-linear relationships between $\lambda$ vectors and alignment performance.
-D.2 Polynomial Regression
+C.2 Polynomial Regression
 Using the training set $\{(\lambda^{(i)},y^{(i)})\}_{i=1}^{n}$, we fit a least-squares regression model in the lifted feature space to learn the performance landscape:
@@
 approximates the expected performance of any candidate $\lambda$ on downstream alignment tasks.
-D.3 Grid Construction over the Simplex
+C.3 Grid Construction over the Simplex
 To enable tractable sampling from the $\lambda$ space, we construct a discrete candidate set $\{\lambda^{(j)}\}_{j=1}^{k}\subset\Delta^{d}$ via one of the following methods:
@@
   • Dirichlet sampling: draw $\lambda^{(j)}\sim\mathrm{Dir}(\boldsymbol{\alpha})$, where $\boldsymbol{\alpha}=\mathbf{1}$ yields a uniform prior over the simplex.
-D.4 Softmax-Based Scheduling Distribution
+C.4 Softmax-Based Scheduling Distribution
 We convert the predicted scores $s_{j}=f(\lambda^{(j)})$ into a probability distribution over the candidate $\lambda$ vectors using a softmax transformation:
@@
 where the temperature parameter $\tau>0$ controls the exploration–exploitation trade-off.
-D.5 Empirical Instantiation with Five Training Points
+C.5 Empirical Instantiation with Five Training Points
 To validate the polynomial scheduling framework, we use five observed training points—four unimodal $\lambda$ vectors and one uniform configuration—each paired with an empirical alignment accuracy:
-$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
-$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
-$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
-$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
-$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
+$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4563$
+$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4561$
+$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4578$
+$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4553$
+$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4623$
 We set $p=2$, yielding $N=\binom{4+2}{2}=15$ features. Using gradient descent to minimize squared error, we obtain:
-$f(\lambda)$ $=0.35+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
-$+0.02\lambda_{1}^{2}+0.04\lambda_{1}\lambda_{2}+0.03\lambda_{1}\lambda_{3}+0.02\lambda_{1}\lambda_{4}$
-$+0.02\lambda_{2}^{2}+0.02\lambda_{2}\lambda_{3}+0.02\lambda_{2}\lambda_{4}$
-$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
-D.6 Lambda Sampling from Learned Scheduler
-We sample $k=10$ $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ) and compute predicted scores $f(\lambda^{(j)})$, converting them via softmax ($\tau=100$):
+$f(\lambda)$ $=0.36+0.09\lambda_{1}+0.09\lambda_{2}+0.09\lambda_{3}+0.09\lambda_{4}$
+$+0.0061\lambda_{1}^{2}+0.0280\lambda_{1}\lambda_{2}+0.0280\lambda_{1}\lambda_{3}$
+$+0.0280\lambda_{1}\lambda_{4}+0.0060\lambda_{2}^{2}+0.0280\lambda_{2}\lambda_{3}$
+$+0.0280\lambda_{2}\lambda_{4}+0.0068\lambda_{3}^{2}$
+$+0.0280\lambda_{3}\lambda_{4}+0.0056\lambda_{4}^{2}.$
+C.6 Lambda Sampling from Learned Scheduler
+We sample $k=10$ $\lambda$ vectors from Dirichlet(1,1,1,1)(1,1,1,1) and compute predicted scores $f(\lambda^{(j)})$, converting them via softmax ($\tau=100$):
 $p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{10}\exp(\tau\cdot f(\lambda^{(l)}))}$
 $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $f(\lambda)$ $p(\lambda)$
-0.212 0.334 0.245 0.209 0.452 0.095
-0.126 0.237 0.131 0.507 0.450 0.075
-0.542 0.079 0.256 0.123 0.455 0.117
-0.233 0.721 0.020 0.025 0.459 0.185
-0.004 0.334 0.281 0.381 0.449 0.067
-0.320 0.069 0.513 0.099 0.451 0.082
-0.236 0.155 0.358 0.251 0.452 0.093
-0.141 0.110 0.701 0.048 0.454 0.110
-0.070 0.198 0.388 0.344 0.451 0.081
-0.090 0.139 0.306 0.465 0.450 0.095
+0.212 0.334 0.245 0.209 0.462 0.108
+0.126 0.237 0.131 0.507 0.461 0.098
+0.542 0.079 0.256 0.123 0.461 0.099
+0.233 0.721 0.020 0.025 0.459 0.082
+0.004 0.334 0.281 0.381 0.462 0.101
+0.320 0.069 0.513 0.099 0.462 0.104
+0.236 0.155 0.358 0.251 0.462 0.108
+0.141 0.110 0.701 0.048 0.461 0.095
+0.070 0.198 0.388 0.344 0.462 0.104
+0.090 0.139 0.306 0.465 0.462 0.100
 Table 4: Sampled $\lambda$ vectors, their predicted scores $f(\lambda)$, and softmax sampling probabilities $p(\lambda)$ at temperature $\tau=100$.
-D.7 Scheduler vs. Uniform Comparison
-We compare our learned scheduler with a uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ baseline using LLaMA 3.2B Instruct.
-The learned scheduler achieves slightly better average accuracy (0.462 vs.
-0.461), and offers regularization via stochastic $\lambda$ sampling, enhancing robustness.
-Figure 12: Comparison of learned scheduler vs. fixed uniform $\lambda$ across six alignment benchmarks. The learned scheduler yields improved accuracy and diversity.
+C.7 Scheduler vs. Uniform Comparison
+We compare our learned scheduler with a uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ baseline using Llama 3.2-1B Instruct.
+The learned scheduler achieves slightly better average accuracy (%) (46.82 vs.
+46.23), while additionally introducing stochastic $\lambda$ sampling as a regularization mechanism, improving robustness.